A Complete Bibliography of Publications in IEEE Software: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

05 April 2024 Version 2.06

Title word cross-reference

19 [BZA21, GCF22, HHJ^{+}21, MCPdS21, MGD^{+}22, vdBLvdHvG20].

2021 [Ano21a]. 2022 [Ano22a]. 2023 [Ano23a]. 21st [Ano23-82]. 2430 [Abr21]. 25-Year [VORP24]. 26262 [AS21]. 26262-6 [AS21].

40-Year [Eld24a].

50-Year [Zac21]. 500th [BP22].

6 [AS21].

Abandonment [LMPT22]. Abran [Ano22-63]. Abstractions [Fai23b]. Academia [CJ23, GSH20]. Academic [GR21]. Academy [Rai20]. Accelerator [MS21]. Access [Ano20i, SGLSA^{+}21]. Accessibility [JM24c]. Accidental [Ant20]. ACES [EKv20]. Achieve [GMZ21, Ozk21a]. Across [HGS^{+}21]. Act [WBR24]. Action [SBHK21]. Actionable [TJG21]. Actions [Hen24]. Adam [Bey20]. Adapting [MDB22, WDT21]. Adaptive [Mag23, MJGP21, dSTdAdS22]. Address [WCG22]. Adoption [PZZ^{+}20]. Advance [RLTD23]. Advanced [SZ22]. Advancing [ZHX22]. Advantage [ZMR21]. Advice [HLNP23]. Affirmative [SBHK21]. Afford [JM24a]. After [MSK^{+}21]. Against [BK24, Syk20]. Age [PZC^{+}22, Syk20, TK24]. Agile [Ano21b, Ano22b, AS21, BHHvH20, CHAVMD21, Doo20, EKB21, FMP21, FRP^{+}21, Gre20, GGJ20, HDT^{+}23, Hol21a, Kee22a, Poo20, PLC20, VCMR22, Ano21a]. Agile4MLS [VCMR22]. Agility [DS24, PLC20]. Aging [LLA^{+}21]. Agnostic [AHW^{+}24, LL23]. Agreement [SL20]. AI [LZXW23, WBR24, AGC^{+}24, ASS^{+}23b, AMCM24, Ano20a, BKL^{+}21, BGH^{+}22, Bor23a, BOBC22, CHM^{+}20, CHL^{+}20, EL23, Eld24b, EB22, HDT^{+}23, HM21, MAE24, Men20, Men21, Men23, Nej21, NCL20, Ozk20e, Ozk22a, Ozk23b, Ozk23e, PSAH23, RWS^{+}24, SSvG22, SBVO22, SHG^{+}22, SSC^{+}22, SZ22, Spi23, Spi24b, SAGS24, TCHC23, TK24, ZB23]. AI-Assisted [HDT^{+}23]. AI-Augmented [Ozk23e]. AI-Based [AGC^{+}24]. AI-Driven [EB22]. AI-Embracing [RWS^{+}24]. AI-Enabled [Ozk20e, SSvG22, SHG^{+}22]. AI-Intensive [MAE24]. Alain [Ano22-63]. Alexander [Men23]. Alliance [Ano21b, Ano22b]. Almost [GMZ21]. Already [Bos24]. Amazon [Win22b]. Ambition [HvG22]. Among [BCF^{+}20]. Analysis [AS21, DC20, ER23, GR21, GCF22, Hol21a, HGS^{+}21, HP22, NI21, SAGS24]. Analytics [ALB^{+}22, GBDC22, Kim20, LL23, TJG21]. Analyzing [CZ22, SRNK22]. Anaya [Rai20]. Anderson [Rag23]. Anger [NCL20]. Angular [VSS^{+}22]. Annals [Ano20c, Ano20d, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano23e, Ano23f, Ano23-31, Ano24b]. Antipatterns [Amm23, FMP21]. API [NFH22]. APIs [Arc21, Doo23a, ZWL20]. App [Hen21, HGS^{+}21]. Application [CPSS21, LHHK20, MND^{+}21, Ozk23a, VSS^{+}22, WZ20, ZLG^{+}20, ZHX22]. Applications [Ano23-43, ATS20, Ben20, DVE^{+}23, ESvE^{+}21, RPWF21, dSTdAdS22, Ano20b, Ano20f, Ano20g, Ano20e, Ano21n, Ano21k, Ano21l, Ano21m, Ano24d, Ano24c]. Apply [Fai21]. Applying [MHPV21]. Approach [Gre20, Hen22, JG23d, TWRG23, TVH21]. Approaches [ALB^{+}22, JPvdHM23]. Apps [BZA21, GCF22, MGD^{+}22]. Archetype [dSTdAdS22]. Archetype-Based [dSTdAdS22]. Archetypes [ATS20]. Architecting [ASIM24]. Architects [PBB23]. Architectural [Ozk22b, SPAF21]. Architecture [Fai21, Fai23b, HP22, Jun23, Kee22a, Kee22b, MS21, MBMR21, Ozk23b, Ozk23c, dSTdAdS22]. Archiving [WLO21]. Arrangements [SCTR23]. Article [BC21]. Articles [Ano20x, Ano20j, Ano20y, Ano20o, Ano20z, Ano21t, Ano21u, Ano21-73, Ano21-41, Ano21-27, Ano21-28, Ano21-29, Ano22i, Ano22j, Ano22k, Ano23n, Ano23o, Ano23-41, Ano23s, Ano23t, Ano23u, Ano23-32, Ano24o]. Artificial [AK20, Ano23a, BK24, DN20, EB23, LZXW23, NDY22, RMLV^{+}22, VKKA20, WYW20]. Arun [JM24c]. Aspects [AGC^{+}24, LRZB22]. Assess [GFVS20]. Assessing [MBPB23]. Assessment [LMPT22, olTY23]. Assessments [CG23]. Assistants [BK24]. Assisted [HDT^{+}23]. Assumptions [CMSG23, GBDC22]. Assurance [GMdOB22, HP22, JHK22, PAS^{+}22]. Attention [MAM^{+}20]. Attractive [MSmM23]. Augmented [HDT^{+}23, Ozk23e]. authoring [CCSS21]. Automated [Arc21, CCSS21, DLIR^{+}24, GMZ21, Mag24b, ZLG^{+}20]. Automatic [CCPLS21, ERL^{+}21, KWM^{+}21, LPRC21]. Automatically [BOBZ21]. Automating [Spi24a]. Automation [AGS^{+}23, Mag24a]. Automotive [AKG^{+}20, Ant23, WPSK21, dSWLT23]. Autonomic [RJV^{+}22]. Avoiding [AM23]. Award [Ano23-39, Ano23-40]. Awards [Ano21-44, Ano21-43, Ano23-65, Ano23-66, Ano24-32, Ano24-36, Ano24-37, Ano20-57]. Aware [CCDM22, RPWF21, SHG^{+}24]. Awareness [NH20]. Away [HB22, JM21, JM24a]. Axispot [IINN22].

Babel [NHF24]. Back [MBMR21, Ozk23d]. Backgrounds [ÜiYÖTD23]. Bang [MBL^{+}20]. Banking [MS21]. Banks [MS21]. Barriers [ÜiYÖTD23]. Barry [Ozk22c]. Based [AGC^{+}24, ASB^{+}24, ALB^{+}22, BCE^{+}21, FVK21, GFVS20, KHK20, LMFQ24, LLP^{+}20, NRPN21, Ozk21f, SL20, TVH21, ÜiYÖTD23, ZLG^{+}20, dSTdAdS22]. Be [TEB^{+}22]. Become [EF23]. Behavior [JJE^{+}20, MSW24]. Behavioral [CMP^{+}21, Ozk20a, PBC^{+}20]. Being [GMA^{+}24, LLA^{+}21]. Beliefs [dSSCMS23]. Benchmarking [EUG^{+}21]. Benefits [Ano20-45, Ano20-56, Ano21-67, Ano21-68, Ano21-69, Ano21-70, Ano21-71, Ano21-72, Ano23-63, Ano24q, Ano24-30, SZ22, ÜiYÖTD23]. Benifits [Ano23-64]. Bert [Hen20a]. Best [MH23]. Better [BGH^{+}22, KNP^{+}24, Men21, SBVO22, TK21]. Between [BOBZ21, GGT^{+}20, Poo20, RFP^{+}21, dSSCMS23]. Beyond [MS22, RLTD23]. Bias [NRPN21]. Biased [Syk20]. Big [Ano20p, Ano20q, Ano20r, Ano20-28, Ano20-29, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano22t, Ano22p, Ano22q, Ano23y, Ano23z, Ano24j, MBL^{+}20]. Black [Arc21, AHW^{+}24]. Black- [Arc21]. Block [PZZ^{+}20]. Blockchain [BMBDMT22, CS20, LLP^{+}20, PZZ^{+}20, SL20, ELFCFL20, LP20]. Blockchain-Based [LLP^{+}20, SL20]. Blockchains [CCDM22]. Blocks [Ozk20b]. Bloomberg [KWM^{+}21]. Board [Ano20-55, Ano21-75, Ano22-42, Ano22-44, Ano23-70]. Boards [Ano22-55, Ano20-53, Ano20-54, Ano20-58, Ano21-66, Ano21-74, Ano22-45, Ano23-69]. Boehm [Ozk22c]. Booting [CJ23]. Boris [Bla20]. Bot [BGBM24]. Bots [WZGS23]. Bottle [Cla23]. Bounty [MS20]. Box [Arc21]. Boxes [AHW^{+}24]. BPMN [Fun23]. Breaking [SMS20]. Breakthroughs [SD21]. Bridging [lÜlzD24]. Bring [CABD20].

Bringing [vV20]. Brought [Kee22a]. Browser [Shr23]. Bug [MS20, ZMR21]. Bugs [Rag21]. Build [JG23c]. Building [JM21, Ozk20b, Ozk22f, vdBLvdHvG20]. Builds [LZ22]. Bullets [Ozk23f]. Bumps [LSPN21]. Business [EVG22, Fun23, LHHK20, PLC20]. Butcher [Win22a].

C [Tag24a]. Caching [MND^{+}21]. CAI [Ano23a]. Call [Ano20-46, Ano20-47, Ano20x, Ano20j, Ano20y, Ano20o, Ano20s, Ano20t, Ano20z, Ano20-30, Ano20-31, Ano21-44, Ano21-43, Ano21-54, Ano21-55, Ano21-56, Ano21-57, Ano21-58, Ano21t, Ano21u, Ano21-73, Ano21-41, Ano21-27, Ano21-28, Ano21-29, Ano21-42, Ano22d, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22i, Ano22j, Ano22k, Ano22u, Ano22v, Ano23-48, Ano23-49, Ano23-50, Ano23-46, Ano23n, Ano23o, Ano23-41, Ano23s, Ano23t, Ano23u, Ano23-32, Ano23-27, Ano24r, Ano24s, Ano24p, Ano24o, Ano24-35, Ano24i, GR21]. Calm [GSM22]. Can [Gil21, HM21, NDY22, Ozk21a, Ozk23b, PZC^{+}22, SD21, SBVO22, Sel20, SZ22, TEB^{+}22]. Canellas [JM24b]. Cannabis [PSSH23]. Can’t [JM24a]. Capabilities [LPT^{+}22]. Capstone [Mah23]. Care [dSTdAdS22]. Career [Ano23-42, Ano23-51, Ano23-52, Ano24t, Ano24u, Ano24v, HLNP23, Hot23, dSSCMS23, dSSSVM23]. Careers [CPGR^{+}23, FSF^{+}23]. Case [JAP23, KTH^{+}22, MGD^{+}22, NHF24, NCL20, SPS23, WFSH21]. Categories [HGS^{+}21]. Center [Ano23-42, Ano23-51, Ano23-52, Ano24t, Ano24u, Ano24v]. Centered [GMdOB22, HMU^{+}20, HDT^{+}23, JG23d]. Centric [ANR22, CKM^{+}21, Hen22]. Century [Ano23-82]. Certification [BVG^{+}23, VORP24]. Certifications [GvV22]. cfp [Ano22r, Ano22-36]. Chains [DLIR^{+}24, LZ22]. Chair [BC21]. Challenge [EH23a, GBDC22, TSJ21]. Challenges [AaPC23, BGL^{+}21, BCR21, DLJ^{+}21, DM20, KTH^{+}22, MAWH22, MJGP21, TZ22, TCHC23, WCG22, ZWL20]. Challenging [CMSG23]. Change [CHAVMD21, KPM^{+}22]. Changed [WPC22]. Changes [WDT21]. Changing [Eld24b, MDB22, WDT21]. Characterizing [SCTR23]. Chasing [ML21]. Chatbot [PSJPGdL21]. Chatbots [MNDT22]. Chats [SYW22]. Checking [Ozk23c]. Cherny [Bla20]. Chief [Ano23g, Ano23-47, Ano24-31, Eld24a, Ozk23d]. Children [PZC^{+}22]. China [ZCG^{+}21]. Choosing [Pan22, PSJPGdL21]. Chris [Con20, MH23]. Civic [KPM^{+}22]. Claims [JM21]. Claims-Processing [JM21]. Clarifying [FMP21]. Clarity [Hol23]. Class [Mah23]. Client [DVE^{+}23]. Closing [GGT^{+}20]. Cloud [AGS^{+}23, SL20]. Clueless [HvG22]. Co [WCF^{+}23, olTY23]. Co-Designing [WCF^{+}23]. Co-op [olTY23]. Coaching [MHPV21]. Coanalysis [GMdOB22]. Code [Ant20, AGS^{+}23, BGBM24, CC23, Doo22, ER23, Fai20a, Fun23, Hol21b, KPM^{+}22, LaT23, LSPN21, Ozk23c, QT23, SPS23, SWS23, SAPH23, SRR^{+}22, Was23b, Wen23, WZ20, Zac21, dMOU^{+}23]. Codifying [TEB^{+}22]. Cognitive [Mag23, Mag24b]. Collaboration [CS22, DLIR^{+}24, HHJ^{+}21, JvdHPE22, PSAH23]. Collaborations [Ant23, CJ23, Fre23]. Collaborative [GSH20, LRZB22, SSQ^{+}20]. Collect [KHF21]. Collection [LZXW23, BBM^{+}21]. Combined [PREL21]. Combining [ALB^{+}22]. Commands [Spi23]. Commercial [SBVO22]. CommunesPlone [VJ23]. Communication [DVE^{+}23]. Communities [ASS^{+}23a, TZ22]. Community [CPT^{+}20, Ozk21e]. Companies [MDB22, ZCG^{+}21]. Company [DM20, EF23, KHK20]. Comparison [ATA^{+}21, Dau23, dSSCMS23, ZMR21].

Competence [EH23a]. Compiler

[BBM^{+}21]. Complex [ASIM24, Mag23]. Complexity [Ant20, AK22]. Compliance [AS21, GSM22]. Components [LMPT22]. Composition [AKK22]. Computer [Ano20f, Ano20g, Ano20e, Ano20-44, Ano20-48, Ano20-45, Ano20-53, Ano20-49, Ano20-57, Ano20-50, Ano20-46, Ano20-54, Ano20-56, Ano20-51, Ano20-47, Ano20-55, Ano20-52, Ano20i, Ano21n, Ano21k, Ano21l, Ano21m, Ano21-54, Ano21-67, Ano21-60, Ano21-55, Ano21-61, Ano21-68, Ano21-62, Ano21-69, Ano21-56, Ano21-63, Ano21-70, Ano21-59, Ano21-57, Ano21-66, Ano21-64, Ano21-71, Ano21-58, Ano21-72, Ano21-65, Ano21v, Ano21w, Ano21x, Ano21y, Ano22-32, Ano22-37, Ano22-38, Ano22-33, Ano22-31, Ano22-36, Ano22-39, Ano22-34, Ano22-40, Ano22-35, Ano22-41, Ano22-47, Ano23-43, Ano23-48, Ano23-58, Ano23-64, Ano23-49, Ano23-59, Ano23-60, Ano23-50, Ano23-51, Ano23-56, Ano23-65, Ano23-53, Ano23-61, Ano23-44, Ano23-63, Ano23-55, Ano23-62, Ano23-57, Ano23-47, Ano23-46, Ano23-45, Ano23-66, Ano23-54, Ano23-52, Ano23p, Ano23q, Ano23r]. Computer [Ano24c, Ano24-31, Ano24r, Ano24q, Ano24-27, Ano24w, Ano24t, Ano24s, Ano24-30, Ano24-28, Ano24x, Ano24u, Ano24p, Ano24z, Ano24-29, Ano24-32, Ano24y, Ano24v, SSvG22, SBHK21, Ano20-32, Ano20b, Ano22d, Ano23g]. Computers [Ano24d, Ano20s, Ano20t, Ano20-30, Ano20-31, Ano21-42, Ano22u, Ano22r, Ano22v, Ano23-27]. Computing [Ano20h, Ano21o, Ano21p, Ano21q, Ano21s, Ano21r, Ano22y, Ano22e, Ano22g, Ano22f, Ano22h, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23-55, Ano23m, Ano24e, Ano24f, Ano24g, HPEP21, LDM^{+}21, SGLSA^{+}21, TSW21, olTY23, Ano20c, Ano20d, Ano20j, Ano20u, Ano20v, Ano20w, Ano20z, Ano21c, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21d, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano22k, Ano22l, Ano22w, Ano22x, Ano22s, Ano23b, Ano23c, Ano23e, Ano23f, Ano23s, Ano23t, Ano23u, Ano23-31, Ano23-28, Ano23-29, Ano23-30, Ano24a, Ano24b]. Computing-Where [TSW21]. ComputingEdge [Ano21e, Ano22c, Ano23d]. Concepts [AaPC23]. Conceptualization [LDM^{+}21]. Conceptualizing [EBC^{+}22]. Concerns [CLSS22, LHHK20, Ozk22b]. Conference [Ano23a, CPM^{+}20]. Conformance [Ozk23c]. Congress [Ano22a]. Connected [Ano20-73]. Connecting [LaT23, VORP24]. Connectivity [NVV^{+}24]. Consent [PZC^{+}22]. Considerations [CKT24]. Considering [AGC^{+}24]. Constructing [PA21]. Contact [BZA21, GCF22]. Contact-Tracing [BZA21, GCF22]. Contents [Ano20-75, Ano20-76, Ano20-77, Ano20-78, Ano20-79, Ano21-88, Ano21-89, Ano21-90, Ano21-91, Ano21-92, Ano21-93, Ano22-69, Ano22-70, Ano22-71, Ano22-72, Ano22-73, Ano23-91, Ano23-92, Ano23-93, Ano23-94, Ano23-95, Ano23-96, Ano24-44, Ano24-45, Ano24-46, Ano20-74]. Contexts [LBF^{+}22]. Continual [BCE^{+}21]. Continuous [DM20, GMZ21, GRMR21, KGW23, MBL^{+}20, ZPKS24, dSWLT23]. Contract [CS20, JS20, LP20]. Contracts [FLB^{+}20, HP22, WLL20, WZ20]. Contrasting [MBL^{+}20]. Contributions [FCTP21, SRNK22, Zac21]. Contributors [AVvBH22, WCG22]. Control [Fai20c, Sel20]. Controlled [FDHB20]. Controllers [Sel20]. Controlling [CAP21, Sel20]. Conversations [Spi23]. Cope [FWBW23]. Correction [Ano20-33, Ano21-45]. Corrections [Ozk23f]. Counterproductive [TEB^{+}22]. Countries [FCTP21]. Course [CG23, LMFQ24, Ozk23f, Spi24a]. Courses [BGBM24, CJ23, LPBG^{+}23]. Cover [Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano21-47, Ano21-48, Ano21-49, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano24l, Ano24m, Ano24n, Ano20-36, Ano21-50, Ano21-51, Ano21-52, Ano22z]. Covered [Ano23-56, Ano23-57, Ano24z]. COVID [BZA21, GCF22, HHJ^{+}21, MCPdS21, MGD^{+}22, vdBLvdHvG20]. COVID-19 [BZA21, GCF22, HHJ^{+}21, MCPdS21, MGD^{+}22, vdBLvdHvG20]. Creating [Fun23, LSN21]. Creativity [IAV24, JPvdHM23]. Crisis [AK20, ABD^{+}21, MCPdS21]. Critical [CHAVMD21, DMBA20]. Crnković [Ano22-54]. Cross [JJE^{+}20, Hen24]. Cross-Tool [JJE^{+}20]. Crossing [Ozk22d]. Crowdsourced [AVvBH22]. Crowdsourcing [WYW20]. CS [Ano22-42, Ano22-48, Ano23-67]. CSCFP [Ano23-68]. Current [CMSG23, VKKA20]. Curriculum [RWS^{+}24, TKHB23]. Cyber [SHG^{+}24, dSWLT23]. Cyber-Physical [SHG^{+}24]. Cyberattacks [dNRR23]. Cyberphysical [MSW24, WDT21]. Cybersecurity [EB23, MS20]. Cycles [Fai23a].

D&I [Ano23-53]. DAF [RPWF21]. Daily [SMS20]. Danish [Fre23]. Dare [Bor23b]. Data [ALB^{+}22, AVvBH22, ANR22, BATM24, CS22, CKM^{+}21, HHJ^{+}21, IINN22, Kim20, LRZB22, MH23, PA21, RMLV^{+}22, RA22, SHG^{+}22, TSJ21, WLO21, WPC22, rTRN^{+}22, vV20, Ano20p, Ano20q, Ano20r, Ano20-28, Ano20-29, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano22t, Ano22p, Ano22q, Ano23y, Ano23z, Ano24j]. Data-Centric [ANR22, CKM^{+}21]. Data-Driven [PA21, TSJ21]. Data-Intensive [BATM24]. Database [Con21]. Databases [Win22b]. Dave [Hen24]. Days [DN20]. Debt [ATA^{+}21, Bor23b, BZS21, CLSS22, CLM21, Fai20d, Fai23a, FRP^{+}21, FB21, JG23a, KHF21, LDM^{+}21, MTR^{+}21, MBPB23, NI21, Ozk21d, RFP^{+}21, TSJ21]. Debugging [SAGS24]. Decentralization [TZ22]. Decentralizing [SWS23]. Decision [Kee22a, Kee22b, LHLLL24, PZZ^{+}20, TK21, CCSS21]. Decision-Making [TK21]. Decisions [TK21]. Deep [EUG^{+}21, ZLG^{+}20]. Defect [MBL^{+}20]. Defending [BK24]. Deficiencies [TEB^{+}22]. Deficits [TKHB23]. Defining [JG23a]. Degrees [Dau23]. Delivery [GRMR21, MNDT22, ZPKS24]. Demands [TKHB23]. Dependability [GMdOB22]. Dependability-Centered [GMdOB22]. Dependable [PA21]. Dependency [RPWF21, Tag24b]. Dependency-Aware [RPWF21]. Deploying [NDY22, vdBLvdHvG20]. Deployment [DM20, Ozk20c]. Deprecation [NFH22]. Design [ANR22, CEL20, Hal21, HMU^{+}20, HP22, JHK22, KHK20, LaT23, LLP^{+}20, LTHY21, MAM^{+}20, Ozk20b, Pen20, PAS^{+}22, RSN20, TK21, WZ20, LZXW23]. Design-Led [KHK20]. Designers [PREL21]. Designing [LZXW23, Mag24b, WCF^{+}23]. Developer [BCHA23, GJH^{+}23, GJHL24, Her21, IAV24, JG23b, JG23c, JG23d, NH20, Ozk22e, SYW22]. Developers [ASS^{+}23b, AMCM24, BPS20, BCHA23, CPGR^{+}23, JvdHPE22, SD21, ZCG^{+}21, dMOU^{+}23, CCSS21]. Developing [HLNP23, MJGP21, VCMR22, WZGS23, WPP20]. Development [ABD^{+}21, AS21, BCF^{+}20, BK24, CC23, CMSG23, EB22, FWBW23, FCTP21, GGJ20, HP22, JPvdHM23, JJE^{+}20, KMFC21, LSPN21, MHPV21, OSMB20, Ozk23b, Ozk23e, PSJPGdL21, RMLV^{+}22, SRP20, SRR^{+}22, WCF^{+}23, ZPKS24]. DevOps [BH22, EH23b, HFLR20, ML21, DCD^{+}23, SPS23, ZWL20]. DevSecOps [AaPC23]. Differences [Zac21]. Different [FCTP21, LL23, Ozk20e]. Difficulties [BBM^{+}21]. Digest [CCSS21]. Digital [ALB^{+}22, AK22, CSC^{+}22, CKWB22, DSE20, EBC^{+}22, IINN22, KHK20, LPT^{+}22, MSW24, Ozk22b, PZC^{+}22, RJV^{+}22, WDT21, dSWLT23, MS21]. Dimensionality [CWGS24]. Direction [TKHB23]. Directions [DC20, MJGP21, TCHC23]. Disabilities [CPGR^{+}23]. Disaster [GRMR21, MGD^{+}22]. Disciplined [Ozk23f]. Distributed [IINN22, VCPC20]. Diverse [PSSH23]. Diversity [AK20, ABD^{+}21, Ano22-37, Ano23-55, Ano23-54, Ano24w, Ano24x, Ano24y, CMP^{+}21, CPT^{+}20]. Divide [Ozk22d]. DMN [Fun23]. DNA [OSMB20]. DNA-Inspired [OSMB20]. Do [AMCM24, JM23, KMFC21, Ozk21d, dSSSVM23, SRNK22, TJG21, WCG22]. Does [Gre22, Mag24a, Pen20]. Doesn’t [Mag24a]. Domain [BGL^{+}21, SBL^{+}24, WZ20, Hen20a]. Domain-Specific [WZ20]. Don’t [AM23, Hoc23b, MJRA23]. Dramatically [SD21]. Drive [Ano23-55]. Driven [CKM^{+}21, ER21, EB22, Fav23, FWBW23, NVV^{+}24, PA21, TSJ21]. Drivers [PZZ^{+}20]. Drives [BZA21]. DRT [WLO21]. Dual [SRP20]. Dual-Track [SRP20]. During [LSPN21, MCPdS21].

Early [TWRG23]. Early-Exposure [TWRG23]. eBay [Jun23]. Ecosystem [AKG^{+}20, ZHX22]. Ecosystems [DLJ^{+}21, NVV^{+}24, RA22]. Edge [Ano22y, Ano20h, Ano21o, Ano21p, Ano21q, Ano22e, Ano22g, Ano22f, Ano23b, Ano23h, Ano23i, Ano23j, Ano23k, Ano24e, Ano24f, Ano24g]. Editor [Ano23g, Ano24-31, Eld24a, Ozk23d, Ozk23f]. Editors [Ano23-47]. Education [Ano21-44, BK24, Dau23, GGT^{+}20, TKHB23, TK24, olTY23]. Educational [BK24, ÜiYÖTD23]. Effect [CHM^{+}20]. Effective [LBF^{+}22]. Effects [GGJ20, SSQ^{+}20]. Eight [GCF22, JPvdHM23]. eknows [GMPS24]. Election [Ano21-59]. Elephant [SvC20]. Else [WPC22]. Embedded [LMPT22, WDT21]. Embedding [SBL^{+}24]. Embracing [AaPC23, RWS^{+}24]. Emerging [KTH^{+}22, LHHK20, ZWL20]. Emotional [MHPV21]. Empirical [RWS^{+}24, GMA^{+}24]. Employability [BPS20]. Employee [JAP23]. Emulating [Was23b]. Enabled [CKWB22, Ozk20e, SSvG22, SHG^{+}22, VCMR22]. Enablers [AKG^{+}20]. Enabling [JJE^{+}20]. End [MS22]. Energy [PSAH23]. Engagement [GSH20, WYW20]. Engineer [Ozk20f, Ozk22a]. Engineering [ASB^{+}24, AK20, BKL^{+}21, BGL^{+}21, BGH^{+}22, BP22, Bor24, BOBC22, BHHvH20, CJ23, CS20, CP20, CMP^{+}21, CPSS21, CKWB22, CKB^{+}21, CKM^{+}21, DN20, Dau23, Doo23b, DMBA20, EKB21, ER21, FVK21, FSF^{+}23, GGT^{+}20, GSH20, GR21, GMA^{+}24, Gre20, Gre21, GPS23, HB22, HMU^{+}20, HLNP23, IAV24, KP20, Kim20, KGW23, KTH^{+}22, LP20, LLA^{+}21, LRZB22, LPBG^{+}23, Mag23, Mah23, MGD^{+}22, MAE24, MJRA23, Ozk20a, Ozk20e, Ozk21b, Ozk21e, Ozk21f, Ozk22d, Ozk23a, PS22, PBC^{+}20, PLT23, RWS^{+}24, SSvG22, SSC^{+}22, SZ22, SBHK21, SL20, TKHB23, TK24, TWRG23, TSJ21, TEB^{+}22, WDT21, WFSH21, ZB23, dGEH^{+}21, olTY23, Ano21c, Ano21s, Ano21r, Ano21d, Ano22h, Ano23c, Ano23l, Ano23m, Ano24a]. Engineers [BCF^{+}20, CG23, DC20, Gre22, JM24a, KPM^{+}22, LMFQ24, MCPdS21, MSmM23, Was23a]. Enhancing [JAP23]. Enrollment [SBHK21]. Enterprise [Amm23]. Entur [RLTD23]. Environments [EUG^{+}21]. Episode [BP22]. Epistemological [GR21]. Era [ASS^{+}23b, JAP23]. Errata [Ano20-34, Ano20-35]. Erratum [Ano21-46]. Error [LSPN21]. Errors [JS20, NH20]. Essential [Ant20]. Ethereum [JS20]. Ethical [CKT24, MJRA23]. Ethics [JM24a, Men23, VKKA20]. European [GCF22, WBR24]. Evaluating [Ant20, KLRT23]. Evaluation [VORP24, dGEH^{+}21]. Even [Hoc23b]. Every [Her21]. Everything [SRR^{+}22, Was23b]. Everywhere [JM23]. Evidence [ASB^{+}24, GSH20, GMA^{+}24, NRPN21]. Evidence-Based [ASB^{+}24]. Evolution [BH22, ZPKS24]. Evolve [WPSK21]. Evolving [Jun23]. EvoMaster [Arc21]. Execution [Fun23, KNP^{+}24]. Exhaustion [Hoc23a]. Experience [DM20, JAP23, Mah23, DCD^{+}23, dSSCMS23, VCMR22, WCF^{+}23, WAHJ24, olTY23]. Experienced [MTR^{+}21]. Experiences [LBF^{+}22, MCPdS21, ZWL20]. Experiential [NI21]. Experiment [FDHB20, VSS^{+}22]. Experimental [CG23]. Expert [CHL^{+}20, CCTH23, KNP^{+}24, LP21]. Explainability [CCTH23, DLIR^{+}24, LHLLL24]. Explainable [TCHC23]. Explaining [AHW^{+}24, CWGS24, MSW24]. Explanations [AM23, AHW^{+}24, LL23]. Exploratory [XJA21]. Exploring [PSSH23]. Exposure [TWRG23]. Extension [Wen23]. Extensions [Shr23]. Extracting [CAP21, CCSS21]. Eyecare [VS20].

FaaSifier [RPWF21]. Facebook [BKL^{+}21]. Factors [PZZ^{+}20, WCG22]. Factory [MAM^{+}20]. Facts [dSSCMS23]. Failures [Hoc23b, WLL20]. Fake [MH23]. Fallacy [ML21]. Fast [GPS23]. Fatigue [Hol20b]. Fault [CPM^{+}20]. Fear [NCL20]. Feature [AKK22, PREL21]. Features [NHF24]. Feedback [BGBM24, LPBG^{+}23, RA22, SSQ^{+}20]. Female [PREL21, SBHK21]. Fighting [JM24b]. filler [Ano22g, Ano22-36, Ano22-43, Ano22-44, Ano22-47, Ano22l, Ano22-49, Ano22-50, Ano22-52, Ano22p, Ano22q, Ano22r, Ano22-53, Ano22s]. First [SRNK22]. Fit [Gil21]. Five [Men20]. Fix [Fai23a]. Flaky [SAHH23]. Flow [BCHA23]. Focus [BCHA23, HGS^{+}21]. Focusing [ASS^{+}23b, CWGS24]. Fog [Hal21]. Ford [Tag24b]. Formal [CG23]. Formative [CG23]. Forward [Ozk23d]. Fostering [TWRG23]. Foundational [SPS23]. Four [CABD20, CKB^{+}21]. Fractal [RMLV^{+}22]. Framework [AaPC23, FVK21, LHHK20]. Free [Fav23]. Freshness [Tag24b]. Friction [BCHA23]. Frisbie [Shr23]. Front [Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano21-47, Ano21-48, Ano21-49, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano24l, Ano24m, Ano24n, Ano20-36, Ano21-50, Ano21-51, Ano21-52, Ano22z]. Frontier [Ozk23e]. Frustrations [HB22]. Function [LTHY21]. Functional [Abr21, BC21]. Functionality [Con20]. Functions [LSN21, NVV^{+}24]. Fund [Ano23-53, Ano24y, Ano22-37]. Future [Bos24, BCL^{+}21, CLM21, CMSG23, Gre22, LMFQ24, LPT^{+}22, Ozk21b, Rai20, SCTR23, TCHC23, TK24]. Fuzz [Win22a]. Fuzzing [BCR21, KNP^{+}24].

Galante [Doo23b]. Game [BZS21]. Games [DC20]. Gamification [EVG22]. Gap [BOBZ21, GGT^{+}20, NI21]. Garbage [KHF21]. Gender [CPT^{+}20, GLBS20, Zac21]. Gendered [MCPdS21]. General [Ano22-43, Ano22-49]. Generating [ATS20]. Generation [AGS^{+}23, Nej21]. Generative [BK24, EL23, Eld24b, Ozk23b, Spi24b]. Generic [Ano22r, Ano23-67]. Genie [Cla23]. Getting [Fai21]. Girls [Ozk21c]. Git [BBM^{+}21]. GitHub [Hen24, NRPN21, SRNK22, WZGS23]. GitOps [BH22]. GNU [BBM^{+}21]. Go [DCD^{+}23, NI21]. Goes [LSN21]. Going [MS22, TmAM24]. Gold [DJF22]. Google [MGD^{+}22]. Governance [EVM20]. Grall [Man21]. Graph [vV20]. Graphics [Ano20b, Ano20f, Ano20g, Ano20e, Ano21n, Ano21k, Ano21l, Ano21m, Ano23-43, Ano24d, Ano24c]. GraphQL [Doo23a]. Grassroots [KPM^{+}22]. Gray [GR21]. Great [Ozk22d]. Green [VLEd21]. Ground [TEB^{+}22]. Ground-Truth [TEB^{+}22]. Group [GGJ20, KLRT23]. Group-Development [GGJ20]. GUI [ZLG^{+}20]. Guidance [KNP^{+}24]. Guide [CKM^{+}21, EAM24, Ozk23b, Ozk23f]. Guidelines [WZGS23]. Guinean [KTH^{+}22]. GWT [VSS^{+}22].

Hackathons [PMGP21]. Hadoop [CZ22]. Handling [LSPN21]. Happens [TmAM24]. Harbor [CABD20]. Harder [Fai21]. Hardware [ZHX22]. Harness [SZ22]. Hazard [MH23]. Heading [TSW21]. Headwinds [Kee21]. Health [AVvBH22, ATS20, LLA^{+}21, Ozk21e, dSTdAdS22]. Health-Care [dSTdAdS22]. Help [CABD20, Pen20, SD21, SBVO22]. Here [Bos24, EB22]. Highlights [ASB^{+}24, CPM^{+}20]. History [Ano20c, Ano20d, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano23e, Ano23f, Ano24b]. Hitchhiker [CKM^{+}21]. Hoban [Doo22]. Hold [Gre22]. HOST [Ano20-42]. Hubert [Hen20a]. Human [BAEM23, FCTP21, HMU^{+}20, HDT^{+}23, JG23d, NH20, Ozk20a, WFSH21]. Human-Centered [HMU^{+}20, HDT^{+}23, JG23d]. Humans [BCHA23, GJH^{+}23, GJHL24, IAV24, JG23b, JG23c]. Humphrey [Ano22-74, Ano23-39, Ano23-40]. Hybrid [ALB^{+}22, CMSG23, JPvdHM23, JG23b, JAP23, SCTR23, SSH23, WCF^{+}23]. Hype [FWBW23, PZZ^{+}20].

IaC [AaPC23]. ICCQ

[Ano20-43, Ano21-53]. Idea [EAM24]. Ideas [Eld24c, Mag24b]. Identification [ZLG^{+}20]. Identifying [OSMB20, dMOU^{+}23]. Identity [LLP^{+}20]. IEEE [Ano20f, Ano20k, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano21-32, Ano21-33, Ano22a, Ano23a, Abr21, Ano20-44, Ano20-48, Ano20-45, Ano20-53, Ano20-49, Ano20-57, Ano20-50, Ano20-46, Ano20-54, Ano20-56, Ano20-51, Ano20-47, Ano20-55, Ano20-52, Ano20-58, Ano20-59, Ano20-60, Ano21-54, Ano21-67, Ano21-60, Ano21-55, Ano21-61, Ano21-68, Ano21-62, Ano21-69, Ano21-56, Ano21-63, Ano21-70, Ano21-59, Ano21-57, Ano21-66, Ano21-64, Ano21-71, Ano21-58, Ano21-72, Ano21-65, Ano21-73, Ano21-75, Ano21-74, Ano22-42, Ano22-32, Ano22-37, Ano22-38, Ano22-33, Ano22-31, Ano22-36, Ano22-39, Ano22-34, Ano22-40, Ano22-35, Ano22-41, Ano22-43, Ano22-46, Ano22-45, Ano22-44, Ano22-47, Ano22-48, Ano22-49, Ano22-50, Ano22-52, Ano22-51, Ano22-53, Ano23-68, Ano23-42, Ano23-67, Ano23-43, Ano23-48, Ano23-58, Ano23-64, Ano23-49, Ano23-59, Ano23-60, Ano23-50]. IEEE [Ano23-51, Ano23-56, Ano23-65, Ano23-53, Ano23-61, Ano23-44, Ano23-63, Ano23-55, Ano23-62, Ano23-57, Ano23-47, Ano23-46, Ano23-45, Ano23-66, Ano23-54, Ano23-52, Ano23-41, Ano23-69, Ano23-70, Ano23-73, Ano23-71, Ano23-72, Ano23-74, Ano24-31, Ano24r, Ano24q, Ano24-27, Ano24w, Ano24t, Ano24s, Ano24-30, Ano24-28, Ano24x, Ano24u, Ano24p, Ano24z, Ano24-29, Ano24-32, Ano24y, Ano24v, Ano24o, Ano24-33, Ano24-34, Ano24-35, Ano24-36, Ano24-37, BC21, Ano20c, Ano20d, Ano20h, Ano20g, Ano20e, Ano20i, Ano20j, Ano20o, Ano20n, Ano20l, Ano20m, Ano20u, Ano20v, Ano20w, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21o, Ano21p, Ano21q, Ano21n, Ano21k, Ano21l, Ano21m, Ano21s, Ano21r, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28].

IEEE [Ano21-29, Ano21-31, Ano21-30, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano22d, Ano22e, Ano22g, Ano22f, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22t, Ano22p, Ano22q, Ano22u, Ano22r, Ano22v, Ano22w, Ano22x, Ano22s, Ano23e, Ano23f, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23g, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23x, Ano23w, Ano23-31, Ano23y, Ano23z, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano24b, Ano24e, Ano24f, Ano24g, Ano24d, Ano24c, Ano24i, Ano24h, Ano24j, Ano24k]. Impact [BVG^{+}23, CG23, EAM24, Eld24a, Mag24b, SBHK21, VS20, dNRR23]. Impediments [AKG^{+}20]. Implications [Ozk23a]. Importance [Hen22]. Impossible [GMZ21]. Improve [EVGA24]. Improving [BBM^{+}21, CPM^{+}20, NH20]. Incentivization [DJF22]. Inclusion [Ano22-37, Ano23-55, Ano23-54, Ano24w, Ano24x, Ano24y, KTH^{+}22]. Inclusive [PMGP21]. Inclusivity [GLBS20]. Increases [PREL21]. Increasing [LZ22, MND^{+}21, SPS23]. India [VS20]. Individualism [BZA21]. Industrial [GGT^{+}20, HM21, SPAF21, SBVO22, SHG^{+}22, VCMR22, VKKA20]. Industry [BZS21, CJ23, EM22, GSH20, GMPS24, KGW23, RSN20, dSSSVM23, TKHB23, WDT21]. Infancy [LRBM23]. Inference [EUG^{+}21]. Information [Ano20-32, Ano20-48, Ano20-49, Ano20-50, Ano20-51, Ano20-52, Ano21-60, Ano21-61, Ano21-62, Ano21-63, Ano21-64, Ano21-65, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano23-58, Ano23-59, Ano23-60, Ano23-61, Ano23-62, Ano24-27, Ano24-28, Ano24-29, CAP21, LaT20, MSK^{+}21]. Infrastructure [AGS^{+}23, Bor23a, Doo22, Ozk23c, QT23, SPS23, SWS23, SAPH23]. InnerSource [DJF22]. Innovating [GMPS24].

Innovation [BATM24, CPGR^{+}23]. Innovative [HFLR20]. Insights [Bor24, CPSS21, CPT^{+}20, NRPN21, ÜiYÖTD23, WPSK21]. Inspired [OSMB20]. Instinctively [BK24]. Integrating [BK24, HMU^{+}20]. Integration [DM20, FLB^{+}20, GvV22, GMZ21, LL23, MBL^{+}20, ZPKS24]. Integrity [LZ22]. Intel [GBDC22]. Intellectual [Fai20c]. Intelligence [AK20, Ano23a, BK24, DN20, EB23, LZXW23, NDY22, RMLV^{+}22, VKKA20, WYW20]. Intelligence-Powered [WYW20]. Intelligent [Ben20]. Intensive [BATM24, CWGS24, MAE24, RJV^{+}22, RMLV^{+}22]. Interaction [LTB^{+}20]. Interactions [AKK22]. Interactive [FVK21]. Interdisciplinary [rTRN^{+}22]. Interface [LHHK20]. International [GSH20]. Internet [SHG^{+}22]. INTERO [SPE20]. Interoperability [SPE20]. Interplay [KGSA20]. Interpretability [AHW^{+}24]. Intersection [CHM^{+}20]. Interview [JM23, JM24b, JM24c, MH23, Men23]. Introducing [KGW23]. Introduction [KWM^{+}21]. Intuition [PREL21]. Invoking [LSN21]. ISO [AS21]. Isomorphic [ZLG^{+}20]. iSPY [SYW22]. Issue [Ano22d, Ano24i, ERL^{+}21, SYW22]. Issues [MS20]. Istio [MBMR21]. Italian [Fav23]. Iteration [Fai22]. Iterations [Fai20b]. Ivica [Ano22-54].

James [Rag21]. Jason [Tag24a]. Java [Her21]. JavaFX [RK20]. JavaScript [NFH22, NHF24]. Jeanna [JM23]. Jeremy [Doo20]. JIT [CPM^{+}20]. Job [Ano20-53, Ano20-54, Ano20-58, Ano21-66, Ano21-74, Ano22-55, Ano23-69]. Jobs [Ano20-55, Ano21-75, Ano22-42, Ano22-45, Ano22-44, Ano23-70]. Jobsboard [Ano22-46]. John [Rag23]. Joint [Mag24b]. Jon [Amm23]. Journal [Ano20i, Ano21v, Ano21w, Ano21x, Ano21y, Ano22-47, Ano22-48, Ano23p, Ano23q, Ano23r]. Journey [Fav23, Rai20, RLTD23, VORP24]. Joy [NCL20]. Jump [FSF^{+}23]. Jump-Starting [FSF^{+}23]. Just [MJRA23, MS22]. Justin [Hen20b].

Keeping [Hol21a]. Kevin [JM24c]. Keyless [HvG22]. KEYS [Men21]. Kinds [Fai22, NCL20]. Kit [MS21]. Knights [DJF22]. Know [Her21, KMFC21, Ozk20f, WPC22]. Knowledge [HFLR20, Ozk23b, SBL^{+}24, vV20]. Krishnavajjala [JM24c].

Labels [ERL^{+}21]. Laboratory [VORP24]. Lacey [Hen21]. Land [DS24]. Landscape [NFH22]. Language [Bor24, HP22, Ozk23a, SD21, WZ20]. Languages [NI21]. Large [Bor24, HFLR20, Ozk23a, SPE20]. Large-Scale [HFLR20]. Latency [JG23c]. Laws [Men20]. Lead [Fre23]. Lean [KHK20]. Learn [Eld24b, Sel20]. Learned [Ant23, BBM^{+}21, PBB23, RK20, ZPKS24]. Learning [BCE^{+}21, Ben20, CG23, KGSA20, LMFQ24, LTB^{+}20, PA21, PBB23, VCMR22, XJA21, ZLG^{+}20]. Learning-Based [ZLG^{+}20]. Learning-Enabled [VCMR22]. Led [KHK20]. Ledger [VCPC20]. Legacies [OSMB20]. Legacy [VSS^{+}22]. LegacyPro [OSMB20]. Legal [rTRN^{+}22]. Lessons [Ant23, BBM^{+}21, CPGR^{+}23, KPM^{+}22, LaT20, PBB23, RK20, SSvG22, ZPKS24]. Letter [Ano22-63]. Letters [Ano22-56]. Level [MND^{+}21, SPS23, SL20]. Levels [FCTP21]. Leveraging [VCMR22]. Licenses [Hol23]. Lie [AM23]. Life [PLT23]. Lifecycle [AGS^{+}23, FDHB20]. Lightweight [LTHY21]. Like [RWS^{+}24]. Listening [BCF^{+}20]. Listing [Ano20-67, Ano20-68, Ano20-69, Ano20-70, Ano20-71, Ano20-72, Ano21-82, Ano21-83, Ano21-84, Ano21-85, Ano21-86, Ano21-87, Ano22-64, Ano22-65, Ano22-66, Ano22-67, Ano22-68, Ano23-85, Ano23-86, Ano23-87, Ano23-88, Ano23-89, Ano23-90, Ano24-41, Ano24-42, Ano24-43]. Literature [GR21]. Live [PLT23, SYW22]. LiveView [Con20]. Living [BZS21]. Local [AHW^{+}24]. Log [SAGS24]. Logs [JJE^{+}20]. Long [DLIR^{+}24]. Longevity [Ozk21e]. Look [JM24a, RWS^{+}24]. Looking [MSmM23, Ozk23d]. Loops [RA22]. Loss [Fai20c]. Love [Kee22a, NCL20]. Low [CC23, Fun23]. Low-Code [Fun23]. Luca [Doo23b]. Luke [Doo22].

M. [Tag24b]. Machine [Ben20, KGSA20, LMFQ24, Ozk20a, PA21, VCMR22, WDT21, XJA21]. Magazine [Ano20e, Ano20o, Ano20n, Ano22m]. Magic [EM22]. Major [Ano21-43]. Making [Eld24c, LHLLL24, Shr21, TK21]. Malicious [AM23]. Managed [Ozk21d]. Management [BVG^{+}23, FMP21, FRP^{+}21, GRMR21, HDT^{+}23, IINN22, MGD^{+}22, MBPB23, Pan22, RFP^{+}21, TSJ21, Was23a]. Manager [CABD20]. Managing [DLJ^{+}21, FB21, JG23a, MSK^{+}21, Tag24b, HvG22]. Mangler [Wen23]. Man’s [DS24]. Manufacturing [SSvG22]. MapReduce [CZ22]. Marc [JM24b]. March [vdBLvdHvG20]. Maritime [PS23]. Massive [LPBG^{+}23, Spi24a]. Masthead [Ano20-61, Ano20-62, Ano20-63, Ano20-64, Ano20-65, Ano20-66, Ano21-76, Ano21-77, Ano21-78, Ano21-79, Ano21-80, Ano21-81, Ano22-57, Ano22-58, Ano22-59, Ano22-60, Ano22-61, Ano23-75, Ano23-76, Ano23-77, Ano23-78, Ano23-79, Ano23-80, Ano24-38, Ano24-39, Ano24-40]. Matt [Hen21, Shr23]. Matters [CWGS24]. Matthews [JM23]. Maturity [GvV22, Gre20, ML21, SPS23]. McCord [Con20]. McDonald [Tag24a]. Me [AM23, TJG21]. Measure [EVGA24]. Measured [BOBZ21]. Measurement [ATA^{+}21, lÜlzD24]. Measurements [Abr21, BC21]. Measuring [BCHA23, JG23a]. Media [BPS20]. Medical [GSM22, JM21, ME22]. Meet [Bor23a, DS24, Eld24a]. Meetings [SSH23, SMS20]. Member [Ano20-56, Ano21-67, Ano21-68, Ano21-69, Ano21-70, Ano21-71, Ano21-72, Ano23-64, Ano23-63, Ano24-30]. Metamorphic [LSN21, STZC20]. Method [BHHvH20, OSMB20]. Methods [CG23, PS22]. Microservice [MJGP21, ZWL20]. Microservices [MBMR21]. Microsoft [JAP23]. Middle [LTB^{+}20]. Migrated [MBMR21]. Migrating [DVE^{+}23, MAM^{+}20, RK20, VSS^{+}22]. Milestone [BP22]. Miller [Doo20]. Mind [BOBZ21, MAM^{+}20]. Mind-Sets [MAM^{+}20]. Mindset [MJRA23]. Mining [NFH22, NCL20]. Mislead [Ozk21d]. Missing [ERL^{+}21, ZB23]. Mitigation [KLRT23]. ML [ASS^{+}23b, AKK22]. MLOps [LMFQ24]. MLPerf [EUG^{+}21]. Mobile [BGL^{+}21, Hen21, MGD^{+}22, ZLG^{+}20]. Model [AHW^{+}24, CKM^{+}21, GvV22, KGW23, LL23, SPE20, VJ23, XJA21, olTY23]. Model-Agnostic [AHW^{+}24, LL23]. Model-Driven [CKM^{+}21]. Modeling [ASIM24, Bey20, BCL^{+}21, CPM^{+}20, CAP21]. Models [AKK22, Bor24, MSK^{+}21, Ozk23a, RFP^{+}21]. Modern [SAGS24]. Mom [Ozk21c]. Moment [Ozk21f]. Monitoring [vdBLvdHvG20]. Monolith [MBMR21]. Monolithic [MBMR21, RPWF21]. Moran [JM24c]. Motivated [CPM^{+}20]. Moving [IINN22]. Multicloud [SGLSA^{+}21]. Multiconcern [GMdOB22, HP22, JHK22, PAS^{+}22, SRR^{+}22]. Multidomain [NVV^{+}24]. Music [BCF^{+}20]. Must [BAEM23]. My [Hot23]. Myths [dSSCMS23].

Name [Hen20a]. Naming [WPP20]. Nation [Ozk22e]. Natural [SD21]. Naturally [SD21]. Nautilus [FVK21]. Navigated [RLTD23]. Navigating [CABD20, WBR24]. Need [WLL20, WPC22]. Needs [Dau23, GGT^{+}20, GSM22, LaT20, NCL20]. Network [EUG^{+}21, NVV^{+}24]. Neural [EUG^{+}21, VCPC20]. Never [WAHJ24]. Newcomers [SRNK22]. News [CP20]. Next [AGS^{+}23, Nej21, Ozk23e]. Next-Generation [Nej21]. No [DS24, Ozk23f]. No-Man’s-Land [DS24]. Node.js [RPWF21]. Nominations [Ano21-44, Ano21-43]. Non [Abr21, BC21]. Non-Functional [Abr21, BC21]. Nonmerged [NRPN21]. Nontechnical [MTR^{+}21]. Normal [ET21, EM22, Gre21]. Novel [TVH21, WLO21]. Novelty [PREL21]. Nudge [BGBM24]. Numbs [Fai20c]. Numerical [Abr21, BC21].

OAuth [Hen20b]. Objects [IINN22]. Observability [GFJL24, LHLLL24]. Observation [DLIR^{+}24]. Obsolescence [Rag23]. Office [Cla23, GMA^{+}24]. Offshoring [SvC20]. Older [BPS20]. Onboarding [GJH^{+}23, CCSS21]. One [Gil21]. Online [FDHB20, LPBG^{+}23, LTB^{+}20, Spi24a]. Only [Ozk23f]. op [olTY23]. Open [ASS^{+}23a, Ano20i, Ano21v, Ano21w, Ano21x, Ano21y, Ano22-47, Ano22-48, Ano23p, Ano23q, Ano23r, CS22, CZ22, Fav23, Fre23, FCTP21, HHJ^{+}21, Hol23, LMPT22, LRZB22, LRBM23, LPBG^{+}23, LBF^{+}22, RA22, RLTD23, Spi24a, TZ22, VJ23, WCG22, ZPKS24, rTRN^{+}22, WPC22]. Open-Source [Fre23]. openEuler [ZHX22]. Operating [Mag23]. Opportunities [Ozk23a]. Opportunity [DS24]. Orchestration [NVV^{+}24]. Organization [Jun23]. Organizational [LBF^{+}22, ML21]. Organizations [OSMB20]. Orientation [FLB^{+}20]. Oriented [MS21]. Original [VJ23, Wen23]. OSSARA [LMPT22]. Our [Fai20c]. Outback [Poo20]. Overcoming [DM20]. Overview [ATA^{+}21, LLA^{+}21].

Pain [WPP20]. Pair [Spi24b]. Pairing [SYW22]. Pandemic [Gre21]. Panel [Bor24]. Papers [Ano20-46, Ano20-47, Ano20s, Ano20t, Ano20-30, Ano20-31, Ano21-54, Ano21-55, Ano21-56, Ano21-57, Ano21-58, Ano21-42, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22u, Ano22v, Ano23-48, Ano23-49, Ano23-50, Ano23-46, Ano23-27, Ano24r, Ano24s, Ano24p, PSSH23]. Papua [KTH^{+}22]. Parallelism [BBM^{+}21]. Part [BCHA23, GJH^{+}23, GJHL24, IAV24, JG23b, JG23c]. Partner [Fai20a]. Partnership [FMP21]. Past [CLM21, Hol20a, TEB^{+}22]. Patent [GMA^{+}24]. Patented [GMA^{+}24]. Path [LTB^{+}20]. Pattern [LLP^{+}20, LZXW23, ZPKS24]. Patterns [Amm23, WZ20]. Paul [Win22a]. Paying [MAM^{+}20]. Payment [GMZ21]. Pedagogy [TKHB23]. Peer [Ozk21e]. Peer-Review [Ozk21e]. People [HGS^{+}21, MAM^{+}20, Sel20]. Perceive [ÜiYÖTD23, WCG22]. Perceived [GGJ20]. Perceptible [NRPN21]. Perception [Ant20, SPAF21]. Perceptions [AVvBH22]. Performance [CPM^{+}20, CZ22, EUG^{+}21, KGSA20, ML21, MND^{+}21]. Personas [CABD20]. Perspective [BZS21, Eld24a, Nej21, Zac21]. Perspectives [CHL^{+}20, CCTH23, LHHK20, PZZ^{+}20]. Pervasive [Ano20j, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano22k, Ano22l, Ano23s, Ano23t, Ano23u, Ano20z]. Ph.D [Ano22-63]. Phases [KHK20]. Philosophy [AaPC23]. Phoenix [Con20]. Physical [SHG^{+}24]. Physics [ALB^{+}22]. Physics-Based [ALB^{+}22]. Pillars [CKB^{+}21]. Pipeline [Bor23a, ZPKS24].

Pipelines [DM20]. Pitfalls [FRP^{+}21, GLBS20]. Placebo [Abr21, BC21]. Platform [Doo23b, Fun23]. Platforms [LPT^{+}22]. Play [FVK21, MGD^{+}22]. Plea [Ozk21d]. Please [TJG21]. Plug [FVK21]. Plug-and-Play [FVK21]. Podcast [Ano22-62, Ano23-71, Ano23-72, Ano23-82, Ano23-83, Ano23-84, Ano24-33, Ano24-34]. Points [LP21]. Politics [BZA21]. Pooling [VJ23]. Popular [BPS20]. Portrays [BPS20]. Positionality [JM24c]. Postgres [Con21]. Potential [BOBC22]. Power [JM24c]. Powered [WYW20]. Practical [FWBW23, DCD^{+}23]. Practice [BK24, CLM21, DMBA20, ELFCFL20, EH23b, SPAF21, VKKA20, VORP24, lÜlzD24]. Practices [BMBDMT22, GLBS20, HFLR20, LPE^{+}20a, LPE^{+}20b, MS20, TZ22, VCMR22, WCG22]. Practitioner [Gre20]. Practitioners [Ant20, BVG^{+}23, CCSS21, DSE20, EL23, GSH20, LLA^{+}21, MTR^{+}21, ÜiYÖTD23]. Predictability [JG23c]. Predicting [Hol20a]. Prediction [KGSA20]. Prepare [SBVO22]. Preparing [Was23a]. Present [CLM21]. Pressure [KMFC21]. Prevention [CPM^{+}20]. Primary [LBF^{+}22]. Primer [ALB^{+}22, DSE20]. Principles [ED24, TK21]. Prioritization [MTR^{+}21]. Privacy [Ano23-82, BMBDMT22, SAHH23, Ano20n, Ano20k, Ano20l, Ano20m, Ano20-27, Ano21-31, Ano21-30, Ano22m, Ano22n, Ano22o, Ano23v, Ano23x, Ano23w, Ano24h, Ano24k]. Problems [CLSS22]. Procedural [NI21]. Process [BB20, Fun23, OSMB20, Ozk21e, RFP^{+}21]. Processes [Ozk23e]. Processing [JM21, RMLV^{+}22, SD21]. Processing-Intensive [RMLV^{+}22]. Procurement [RLTD23]. Product [VORP24]. Production [Spi24a]. Productivity [BOBZ21, BCHA23, EVGA24, GJH^{+}23, GJHL24, IAV24, JG23b, JG23c, JG23d]. Products [EAM24]. Professional [AGC^{+}24, Ano21-73, Ano20x, Ano20y, Ano21t, Ano21u, Ano21-41, Ano22i, Ano22j, Ano23n, Ano23o, Ano23-41, Ano23-32, Ano24o]. Professionalism [TWRG23]. Professionals [HLNP23, dSSCMS23, dSSSVM23]. Program [AGC^{+}24, CCPLS21, KWM^{+}21, LPRC21, PBB23, TVH21]. Programming [BGBM24, LaT20, LHHK20, Spi24b]. Programs [MS20, olTY23]. Project [BVG^{+}23, CABD20, CZ22, FMP21, KHK20, LMFQ24]. Project-Based [KHK20, LMFQ24]. Projects [BVG^{+}23, FRP^{+}21, GSH20]. Proposals [Ano22d, Ano24-35, Ano24i]. Prospects [LPT^{+}22]. Protect [PZC^{+}22]. Protecting [Ozk21e]. Protestware [CKT24]. Protocol [DVE^{+}23]. Provers [HM21]. Provider [KHK20]. Psychological [TmAM24]. Psychology [GGJ20, Kee22b]. Public [Fav23, Fre23, Hol23, LRBM23, MNDT22, NDY22, RLTD23, VJ23, Zac21]. Publications [Ano23-47]. Publishing [ZPKS24]. Pull [NRPN21]. Pulse [MBPB23]. Puzzle [rTRN^{+}22]. PVC [Ano22-49]. Python [Fun23].

Q&A [GFJL24]. Qualitative [GMdOB22]. Quality [ASS^{+}23a, BGBM24, CWGS24, CCDM22, ER23, GJHL24, GBDC22, GLBS20, HHJ^{+}21, Ozk21a, Rag21, SHG^{+}22, VORP24]. Quality-Aware [CCDM22]. Quantified [Tag24a]. Quantitative [GMdOB22]. Quantum [Ano20-59, Ano20-60, Ano23-73, HPEP21, ZMR21].

Race [NRPN21, Syk20]. Radio [BP22].

RADON [DCD^{+}23]. Rainbow

[Ano22-50, Ano23-71, Ano23-72, Ano23-81, Ano24-34, Ano22-62, Ano23-82, Ano23-83, Ano23-84, Ano24-33]. Ramp [GJH^{+}23].

Ramp-Up [GJH^{+}23]. Randy [Jun23]. Rapid [GRMR21]. Rational [BB20]. rationale [CCSS21]. Rationality [PREL21]. Re [AVvBH22, Gre20, SBVO22]. RE BBC [SL20]. Reaches [BP22]. Readiness [KGW23]. Realizing [BOBC22]. Really [Ozk20e, Ozk21a]. Rebooting [CJ23]. Rebuttal [Ano22-63]. Recommendations [FWBW23, MS20, dMOU^{+}23]. Records [Kee22a, Kee22b]. Recovery [ERL^{+}21]. Redesign [Kee21]. Redis [Man21]. Redoing [CG23]. Reducing [JS20, NH20]. Reduction [CWGS24]. Reference [AaPC23]. Reflections [BCR21, Ozk23d, Ozk23f, Pan22]. Regarding [BC21]. Related [ÜiYÖTD23]. Relationship [BOBZ21, RFP^{+}21]. Reliable [LL23]. Remote [Cla23, TmAM24]. Repair [BCE^{+}21, CCPLS21, KWM^{+}21, LPRC21, TVH21]. Repetition [CG23]. Replication [Wen23]. Report [DM20, Mah23, NI21, DCD^{+}23, VSS^{+}22, WAHJ24]. Reported [BOBZ21]. Reports [MBL^{+}20]. Reproducible [LZ22]. Requests [NRPN21]. Required [Bor23a]. Requirement [GLBS20]. Requirements [BZA21, BGH^{+}22, Bor23a, Bor23b, Bor24, CPM^{+}20, CP20, CAP21, DC20, DN20, DS24, DLJ^{+}21, DMBA20, ER21, Gil21, Gre20, Gre21, Gre22, GBDC22, HMU^{+}20, Hot23, LPT^{+}22, LLA^{+}21, MND^{+}21, Pan22, SL20, SAHH23, WDT21, ZB23, dSWLT23, dGEH^{+}21]. Research [ASB^{+}24, BATM24, CP20, CPSS21, CLM21, CKB^{+}21, GMPS24, SSC^{+}22, SAPH23, VORP24, lÜlzD24]. Researchers [GSH20]. Resilient [Mag24b]. Resist [FWBW23]. Resistance [CABD20]. Resource [Hoc23a, VJ23]. Respond [LHHK20]. Response [BC21]. Responsible [LZXW23, ZB23, LZXW23]. Responsible-AI-by-Design [LZXW23]. RESTful [Arc21]. Résumé [FWBW23]. Résumé-Driven [FWBW23]. REvelation [Hot23]. Review [Ozk21e, PS22]. Reviewing [HGS^{+}21]. Revolution [PS23]. Ri [Gre20]. Richer [Hen20b]. Right [Hol21b, JM24b]. Ripley [Shr21]. Risk [CAP21, GFVS20, KLRT23, LMPT22]. Risk-Based [GFVS20]. Risks [Ozk23a]. Rituals [Fai20b]. Robot [BCE^{+}21]. Robotic [ZLG^{+}20]. Robotics [Dau23]. Robust [GRMR21]. Role [SL20]. Roles [Was23a]. Room [SvC20]. Ross [Rag23]. Rough [CABD20]. Rules [SMS20]. Runaway [HvG22]. Runtime [JS20]. Ryan [Shr21].

Sadness [NCL20]. Safely [CABD20]. Safer [Mag24b]. Safety [AS21, CHAVMD21, TmAM24]. Safety-Critical [CHAVMD21]. Sampling [KGSA20]. Satisfying [MND^{+}21]. Say [dSSSVM23]. Scalability [BHHvH20]. Scalable [Gil21]. Scale [HFLR20, Mag23, SBVO22]. Scaling [LPBG^{+}23, TZ22]. Scenarios [GRMR21, Ozk20d]. Schema [MTR^{+}21]. Science [Ano21c, Ano21s, Ano21r, Ano21d, Ano22h, Ano23c, Ano23l, Ano23m, Ano24a, CMP^{+}21, Ozk20a, PBC^{+}20, SBHK21, TSJ21]. Science/Software [SBHK21]. Scientific [LSN21, WPP20]. Scott [Tag24b]. Scrum [MAWH22, Shr21]. ScrumScale [BHHvH20]. SE [CHM^{+}20, Men20, Men21, NCL20, TCHC23]. SE4AI [AGC^{+}24]. Search [Ano24-31, BATM24, FVK21, Ozk21f, TVH21]. Search-Based [FVK21, Ozk21f, TVH21]. Seas [CABD20]. Seasoned [HLNP23]. Secondary [LBF^{+}22]. Secret [BGH^{+}22]. Sector [Fav23, Fre23, Hol23, LRBM23, NDY22, VJ23]. Sectors [LBF^{+}22]. Security [Ano20n, Ano20k, Ano20l, Ano20m, Ano21-31, Ano21-30, Ano22m, Ano22n, Ano22o, Ano23v, Ano23x, Ano23w, Ano23-82, Ano24h, CPM^{+}20, ERK20, LTB^{+}20, SAHH23, dSWLT23, Ano20-27]. See [Hoc23b]. Seek [Ano23-47]. Seeks [Ano23g]. Selection [DLJ^{+}21]. Self [BOBZ21, LLP^{+}20, MJGP21]. Self-Adaptive [MJGP21]. Self-Reported [BOBZ21]. Self-Sovereign [LLP^{+}20]. Sell [Men23]. Semantic [vV20]. Sensors [SHG^{+}24]. Serebrenik [Men23]. Series [Win22b]. Server [DVE^{+}23]. Serverless [CPSS21, ESvE^{+}21, EUG^{+}21, LP21, LDM^{+}21, LSN21, LTHY21, DCD^{+}23, SGLSA^{+}21, TSW21]. Service [Ano20-57, Ano23-65, BATM24, FLB^{+}20, KHK20, LLP^{+}20, LTHY21, MNDT22, MS21, SL20, Ano23-66, Ano24-32, Ano24-36, Ano24-37]. Service-Level [SL20]. Service-Oriented [MS21]. Services [Ano22a]. Set [Fai23b, PSSH23]. Sets [MAM^{+}20]. Settings [MDB22]. Setup [ZPKS24]. Seven [Ant23]. Sharing [DLIR^{+}24, HFLR20]. Ship [CABD20]. Shockingly [Men21]. Shostack [Bey20]. Should [EB22, Her21, Ozk20f, RWS^{+}24, WPSK21]. Shoup [Jun23]. Shu [Gre20]. Shu-Ha-Ri [Gre20]. Silver [Ozk23f]. Simple [Hol21a, MSK^{+}21, Men21]. Situation [NH20]. Size [Gil21, lÜlzD24]. Sizing [Abr21, BC21]. Skills [Mag23]. Slow [GPS23]. Small [DM20, dNRR23]. Smart [Amm23, CC23, CS20, FLB^{+}20, JS20, LP20, PS23, WLL20, WZ20]. Smells [CPT^{+}20, SPAF21, dMOU^{+}23]. SMILE [AHW^{+}24]. Smith [Rag21]. SOBO [BGBM24]. Social [AGC^{+}24, KPM^{+}22, LTB^{+}20]. Society [Ano20-53, Ano20-57, Ano20-54, Ano20-55, Ano21-66, Ano22-37, Ano23-51, Ano23-66, Ano23-52, Ano24t, Ano24u, Ano24-32, Ano24v, KP20, KTH^{+}22, Ano20-32, Ano20-44, Ano20-48, Ano20-45, Ano20-49, Ano20-50, Ano20-46, Ano20-56, Ano20-51, Ano20-47, Ano20-52, Ano21-54, Ano21-67, Ano21-60, Ano21-55, Ano21-61, Ano21-68, Ano21-62, Ano21-69, Ano21-56, Ano21-63, Ano21-70, Ano21-59, Ano21-57, Ano21-64, Ano21-71, Ano21-58, Ano21-72, Ano21-65, Ano22-32, Ano22-38, Ano22-33, Ano22-31, Ano22-36, Ano22-39, Ano22-34, Ano22-40, Ano22-35, Ano22-41, Ano23-48, Ano23-58, Ano23-64, Ano23-49, Ano23-59, Ano23-60, Ano23-50, Ano23-56, Ano23-65, Ano23-53, Ano23-61, Ano23-44, Ano23-63, Ano23-55, Ano23-62, Ano23-57, Ano23-47, Ano23-46, Ano23-45, Ano23-54, Ano24-31, Ano24r, Ano24q, Ano24-27, Ano24w, Ano24s, Ano24-30, Ano24-28, Ano24x, Ano24p]. Society [Ano24z, Ano24-29, Ano24y, Ano20i, Ano21v, Ano21w, Ano21x, Ano21y, Ano23p, Ano23q, Ano23r]. Soft [SAHH23]. Software [AGC^{+}24, ASS^{+}23a, ASIM24, ASB^{+}24, AK20, ABD^{+}21, Amm23, Ano24-35, AS21, Ant23, ANR22, BKL^{+}21, BPS20, BGL^{+}21, BCF^{+}20, BCE^{+}21, BMBDMT22, BGH^{+}22, BVG^{+}23, BATM24, BP22, BK24, CC23, CG23, CWGS24, CJ23, CMP^{+}21, CABD20, CKB^{+}21, CMSG23, DLJ^{+}21, Dau23, EBW22, EL23, EF23, EVGA24, ED24, Fai21, Fai23b, FVK21, FMP21, FRP^{+}21, FSF^{+}23, GGT^{+}20, GSH20, GR21, GMA^{+}24, GSM22, GJHL24, GGJ20, GBDC22, GPS23, GRMR21, Hal21, HB22, HvG22, HLNP23, HDT^{+}23, Hol23, IAV24, JPvdHM23, JHK22, JJE^{+}20, JM21, JM24a, KGSA20, KP20, Kim20, KGW23, KPM^{+}22, KTH^{+}22, KMFC21, LZ22, LPE^{+}20a, LPE^{+}20b, LL23, LSN21, LRZB22, LRBM23, LHLLL24, LPBG^{+}23, LSPN21, LBF^{+}22, MCPdS21, Mag23, MAM^{+}20, Mah23, ME22, MHPV21, MDB22, NH20, Nej21, NVV^{+}24, OSMB20, Ozk20b, Ozk20a, Ozk20f, Ozk21a]. Software [Ozk21b, Ozk21e, Ozk21f, Ozk22a, Ozk22d, Ozk22f, Ozk23a, Ozk23b, Ozk23c, Ozk23e, PA21, PBB23, PS22, PAS^{+}22, PBC^{+}20, PLT23, PREL21, Rag21, Rag23, Rai20, RWS^{+}24, RJV^{+}22, dSSCMS23, dSSSVM23, SBL^{+}24, Sel20, SSC^{+}22, SZ22, SSQ^{+}20, SBHK21, SCTR23, SRR^{+}22, TZ22, TK21, TKHB23, TK24, TWRG23, TSJ21, TEB^{+}22, ÜiYÖTD23, VS20, VLEd21, VORP24, WYW20, WCG22, WCF^{+}23, Was23a, WFSH21, WPP20, ZHX22, dSTdAdS22, lÜlzD24, Ano20o, Ano24i]. Software-Development [GGJ20]. Software-Driven [NVV^{+}24]. Software-Intensive [CWGS24, RJV^{+}22]. Software-Repair [BCE^{+}21]. Solution [SYW22, vdBLvdHvG20]. Solutions [FRP^{+}21]. Solving [rTRN^{+}22]. Sound [BCF^{+}20]. Source [ASS^{+}23a, CZ22, Fav23, Fre23, FCTP21, Hol23, LMPT22, LRBM23, LBF^{+}22, RLTD23, TZ22, VJ23, WCG22, ZPKS24]. Source-Driven [Fav23]. Sovereign [LLP^{+}20]. SP [Ano22-50, Ano22-52, Ano22-51, Ano23-74]. Spaces [BATM24]. Spark [CZ22]. Spatiotemporal [IINN22]. Special [Ano22d, Ano24i, NCL20]. Specific [WZ20]. Specification [SL20]. Specify [Bor23b]. Spectrum [SCTR23]. Speed [Mag23]. SQL [Spi24c]. Staff [Ano20-67, Ano20-68, Ano20-69, Ano20-70, Ano20-71, Ano20-72, Ano21-82, Ano21-83, Ano21-84, Ano21-85, Ano21-86, Ano21-87, Ano22-64, Ano22-65, Ano22-66, Ano22-67, Ano22-68, Ano23-85, Ano23-86, Ano23-87, Ano23-88, Ano23-89, Ano23-90, Ano24-41, Ano24-42, Ano24-43]. Stakeholders [MS22]. Stand [SMS20]. Stand-Up [SMS20]. Standard [BC21]. Standards [GFVS20, LBF^{+}22]. Stars [DJF22]. Start [LPE^{+}20a, LPE^{+}20b, SMS20]. Start-Ups [LPE^{+}20a, LPE^{+}20b]. Starter [MS21]. Starting [FSF^{+}23]. Startup [Was23b]. State [VKKA20]. States [dNRR23, ZCG^{+}21]. Statistical [AHW^{+}24]. Stay [Ano20-73]. STEALTH [AM23]. Steering [HB22]. Steroids [AKK22]. Still [LRBM23]. Stop [TJG21]. Store [MGD^{+}22]. Stores [XJA21]. Story [Kee22a]. Strategies [BBM^{+}21, DLJ^{+}21, ET21, KLRT23, LBF^{+}22]. Strategy [LHHK20]. Strikes [MBMR21]. Student [KLRT23, LPBG^{+}23]. Students [dSSCMS23]. Studies [SPS23]. Study [ANR22, CG23, JJE^{+}20, NFH22, NCL20, RWS^{+}24, ÜiYÖTD23, Wen23, XJA21]. Subscribe [Ano20b, Ano20f, Ano20g, Ano20e, Ano20n, Ano20k, Ano20l, Ano20m, Ano20-27, Ano21n, Ano21-30, Ano22m, Ano22n, Ano22o, Ano22-52, Ano22-51, Ano23v, Ano23w, Ano23-74, Ano24h]. Subscription [Ano21r, Ano21-31, Ano23x]. Success [Amm23, BZA21, PZZ^{+}20]. Successful [EAM24, Eld24c, Fre23, FCTP21, GSH20]. Such [MSK^{+}21]. Supplier [Ant23, WPSK21]. Supply [LZ22]. Support [AS21]. Supporting [CPGR^{+}23, JHK22, PAS^{+}22]. Surfacing [MAWH22]. Surprise [NCL20]. Survey [EM22, NFH22, ÜiYÖTD23]. Survival [EAM24]. Sustainable [Ano20u, Ano20v, Ano20w, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano22w, Ano22x, Ano22s, Ano23-28, Ano23-29, Ano23-30]. Swing [RK20]. Symbolic [KNP^{+}24]. Symposium [Ano20-42]. Sysparency [GMPS24]. System [HP22, IINN22, MSW24, SL20, Hen20a]. Systematic [LL23, MBPB23, PS22]. Systems [AGC^{+}24, ASIM24, Ben20, CWGS24, CKWB22, CHAVMD21, CKM^{+}21, DLIR^{+}24, EKB21, EBW22, GMZ21, LHLLL24, LZXW23, Mag23, Mag24b, MAE24, MJGP21, Ozk20e, Ozk22f, PLT23, RJV^{+}22, RMLV^{+}22, SSvG22, SHG^{+}24, SPE20, VCMR22, WDT21].

Table [Ano20-74, Ano20-75, Ano20-76, Ano20-77, Ano20-78, Ano20-79, Ano21-88, Ano21-89, Ano21-90, Ano21-91, Ano21-92, Ano21-93, Ano22-69, Ano22-70, Ano22-71, Ano22-72, Ano22-73, Ano23-91, Ano23-92, Ano23-93, Ano23-94, Ano23-95, Ano23-96, Ano24-44, Ano24-45, Ano24-46]. Tabular [BATM24]. Tailor [GFVS20]. Taking [LTB^{+}20]. Tale [DJF22]. Taming [AK22, SHG^{+}22]. Task [MSK^{+}21]. Tasks [Ozk23a, Tag24a]. Taxonomy [GFVS20]. TD [MTR^{+}21, MBPB23]. TD-Experienced [MTR^{+}21]. Teach [MJRA23, NDY22]. Teaching [MAE24, PS22, PLT23]. Team [CPM^{+}20, JPvdHM23, MHPV21, rTRN^{+}22]. Teaming [Ozk20a]. Teams [GGJ20, SBL^{+}24, SSQ^{+}20, SCTR23]. Tech [Fai23a, KPM^{+}22, MAWH22, PSAH23, PS23, Was23b]. Technical [AGC^{+}24, ATA^{+}21, Bor23b, BZS21, CLSS22, CLM21, Dau23, Fai20d, FRP^{+}21, FB21, JG23a, KHF21, LDM^{+}21, MTR^{+}21, MBPB23, NI21, Ozk21d, RFP^{+}21, TSJ21]. Technique [SPS23]. Techniques [LL23]. Technologies [ELFCFL20, NDY22]. Technology [EKv20, ET21, EH23a, vV20]. Tell [BAEM23, TJG21]. Telling [TJG21]. Terminal [GMZ21]. Test [ERK20, ER21, GvV22, Hol20b]. Test-Driven [ER21]. Testing [Arc21, EBW22, Fai20c, GMZ21, GFVS20, KNP^{+}24, LP21, LSN21, dSSCMS23, STZC20, SAGS24, Win22a, ZLG^{+}20]. Tests [Fai20b, Spi24c, SAHH23]. Text [NCL20]. Their [ÜiYÖTD23]. Them [Bor23b, Kee22a]. Theorem [HM21]. Theory [Sel20, WAHJ24]. There [NRPN21]. Things [Her21, SHG^{+}22]. Think [GCF22, Mag24a]. Thinking [CEL20, GPS23, HMU^{+}20, MAM^{+}20, Pen20, RSN20]. Thought [Fai20a]. Threat [Bey20]. Three [KHK20]. Time [JM21, KMFC21, Win22b, ZCG^{+}21]. Time-Series [Win22b]. Times [MDB22]. Timestream [Win22b]. Together [Kee22a]. Tool [ATS20, JJE^{+}20, LP21, Ozk21d, Pan22, PSJPGdL21, WLO21, ZPKS24]. Tools [ATA^{+}21, JvdHPE22, LaT20, LPE^{+}20a, LPE^{+}20b, SAGS24, dGEH^{+}21]. Traceability [MSK^{+}21, WPSK21]. Tracing [BZA21, GCF22]. Track [SRP20]. Trademark [GMA^{+}24]. Tradeoffs [CWGS24]. Training [AGC^{+}24, GGJ20, LMFQ24, TKHB23, TK24]. Trans [Ano22p, Ano22q, Ano22r, Ano22s, PMGP21]. Trans-Inclusive [PMGP21]. Transaction [CCDM22]. Transactions [Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano23-31, Ano23y, Ano23z, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano24j, Ano24k, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano21-42]. Transformation [DSE20, WDT21]. Transforming [KHK20]. Transgender [dSSSVM23]. Transparency [Hol23, SGLSA^{+}21]. Transpilers [NHF24]. Transport [RLTD23]. Trenches [CP20, CPT^{+}20]. Trends [BGL^{+}21, EKv20, ET21, EH23a, GvV22, JHK22, PAS^{+}22, PLC20, ZWL20]. Triangle [EM22]. Triggers [Ant20]. Trust [HM21, Ozk22f]. Truth [TEB^{+}22]. tsusc [Ano22-53]. Tug [Man21]. Tuning [CZ22]. Turnover [SvC20]. Tutorial [ANR22]. Twin [CKWB22, LPT^{+}22, RJV^{+}22, dSWLT23]. Twin-Enabled [CKWB22]. Twins [ALB^{+}22, AK22, CSC^{+}22, EBC^{+}22, IINN22, MSW24, Ozk22b]. Two [Fai22, Her21, SPS23]. Type [ERL^{+}21]. TypeScript [Bla20]. Typology [MNDT22].

U.S. [GMA^{+}24, Hol23]. Ukraine [PSSH23]. UML [BB20]. Unanticipated [Ozk20d]. Uncertainty [FB21, SHG^{+}24]. Uncertainty-Aware [SHG^{+}24]. Unchain [PZZ^{+}20]. Undergraduate [SBHK21, olTY23]. Understanding [Ozk22f, WLL20, ZCG^{+}21]. Unfairness [JM23]. Unified [BB20, FLB^{+}20]. Union [WBR24]. Unit [Spi24c]. United [ZCG^{+}21].

University [Was23b]. Unlock [EVG22]. Unprofessional [BC21]. Unrequited [Kee22a]. Unscheduled [SSH23]. Unseen [Hot23]. Untestable [STZC20]. Unveiling [Hot23]. Upcoming [WBR24]. Ups [LPE^{+}20a, LPE^{+}20b]. Ur [Fai20d]. Ur-Technical [Fai20d]. Usability [Hen21, WCG22]. Usage [Hen22, NHF24]. Usage-Centric [Hen22]. Use [PSSH23]. Users [GCF22, MS22, WPC22]. Using [ATS20, CS22, GBDC22, LBF^{+}22, Mah23, Men23, dSWLT23, AVvBH22].

Validation [CCDM22]. Value [EVG22]. Valued [WCG22]. Values [WFSH21]. Vendor [LP21]. Vendors [Ozk21d]. Venues [SSC^{+}22]. Verification [AKG^{+}20, Nej21, PZC^{+}22, dSWLT23]. Versus [Doo20, GR21, Ozk22a]. Via [CWGS24, GMdOB22]. Video [BZS21]. View [LP21, Ozk20c]. Virtual [AKG^{+}20, SHG^{+}24]. Virtualization [NVV^{+}24]. Virtualize [GMZ21]. Virtuous [Fai23a]. Vision [BK24, Rai20, SSvG22]. Visualizing [CHAVMD21]. Viva [JAP23]. Voice [MTR^{+}21]. Volunteer [Ano20-57, Ano23-65, Ano23-66, Ano24-32, Ano24-36, Ano24-37].

Want [AMCM24]. War [PSSH23]. Wasteland [Poo20]. Watch [Ano20a]. Waterfall [Doo20, Poo20]. Watershed [Ozk21f]. Watts [Ano22-74]. Way [Mag24a, MJRA23]. Web [VSS^{+}22]. Week [Ano23-73, Ano20-59, Ano20-60]. Well [LLA^{+}21]. Well-Being [LLA^{+}21]. Where [DS24, Ozk21c, TSW21]. While [RK20]. White [Arc21]. White-Box [Arc21]. Will [WAHJ24]. Winning [EKv20]. Wireframes [Mah23]. Within [CJ23]. Women [HB22, MAWH22]. Words [HLNP23]. Work [JAP23, Mag24a, MDB22, Ozk21b, Shr21, SCTR23, SSH23, ÜiYÖTD23, WAHJ24]. Work-Related [ÜiYÖTD23]. Worker [WYW20]. Working [CHM^{+}20, Cla23, ZCG^{+}21]. Workplace [SCTR23, WCF^{+}23]. Workplaces [MSmM23]. World [Ano22a, JPvdHM23, WDT21]. Worldwide [GvV22]. Worry [EB22].

Y2K [ZMR21]. Year

[Con21, Eld24a, VORP24, Zac21]. Years [CLM21, LRBM23].

Zealand [MDB22]. Zones [JM21]. Zup [CPGR^{+}23].

References

Alonso:2023:EIT

[AaPC23] Juncal Alonso, Rados aw Piliszek, and Matija Cankar. Embracing IaC through the DevSecOps philosophy: Concepts, challenges, and a reference framework. IEEE Software, 40(1):56–62, January/February 2023. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Albusays:2021:DCS

[ABD^{+}21] Khaled Albusays, Pernille Bjorn, Laura Dabbish, Denae Ford, Emerson Murphy-Hill, Alexander Serebrenik, and Margaret-Anne Storey. The diversity crisis in software development. IEEE Software, 38(2):19–25, March/ April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Abran:2021:INF

[Abr21] Alain Abran. IEEE 2430 non-functional sizing measurements: a numerical placebo. IEEE Software, 38(3):113–120, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Abdellatif:2024:STP

[AGC^{+}24] Ahmad Abdellatif, Gita Ghiasi, Diego Elias Costa, Emad Shihab, and Tanja Tajmel. SE4AI: a training program considering technical, social, and professional aspects of AI-based software systems. IEEE Software, 41(2):44–51, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Aviv:2023:ICN

[AGS^{+}23] Itzhak Aviv, Ruti Gafni, Sofia Sherman, Berta Aviv, Asher Sterkin, and Etzik Bega. Infrastructure from code: The next generation of cloud lifecycle automation. IEEE Software, 40(1):42– 49, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Aslansefat:2024:EBBb

[AHW^{+}24] Koorosh Aslansefat, Mojgan Hashemian, Martin Walker, Mohammed Naveed Akram, Ioannis Sorokos, and Yiannis Papadopoulos. Explaining black boxes with a SMILE: Statistical model-agnostic interpretability with local explanations. IEEE Software, 41(1):87–97, January/ February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Adams:2020:DCS

[AK20] Bram Adams and Foutse Khomh. The diversity crisis of software engineering for artificial intelligence. IEEE Software, 37(5):104–108, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Atkinson:2022:TCD

[AK22] Colin Atkinson and Thomas Kühne. Taming the complexity of digital twins. IEEE Software, 39(2):27– 32, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Agren:2020:AVV

[AKG^{+}20] S. Magnus Agren, Eric Knauss, Paolo Giusto, Grant Soremekun, Rogardt Heldal, and Daniela Damian. The automotive virtual verification ecosystem: Impediments and enablers. IEEE Software, 37(5):67–76, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Apel:2022:FIS

[AKK22] Sven Apel, Christian Kästner, and Eunsuk Kang. Feature interactions on steroids: On the composition of ML models. IEEE Software, 39(3): 120–124, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Adams:2022:HDT

[ALB^{+}22] Matt Adams, Xiao Li, Lucas Boucinha, Sameer S. Kher, Prith Banerjee, and Jose-Luis Gonzalez. Hybrid digital twins: a primer on combining physics-based and data analytics approaches. IEEE Software, 39(2):47– 52, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Alvarez:2023:DLM

[AM23] Lauren Alvarez and Tim Menzies. Don’t lie to me: Avoiding malicious explanations with STEALTH. IEEE Software, 40(3):43–53, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Angelo:2024:WDD

[AMCM24] Sarah D Angelo, Ambar Murillo, Satish Chandra, and Andrew Macvean. What do developers want from AI? IEEE Software, 41 (3):11–15, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ammanath:2023:JSP

[Amm23] Brijesh Ammanath. Jon smart on patterns and antipatterns for enterprise software success. IEEE Software, 40(3):98–100, May/ June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:AW

[Ano20a] Anonymous. AI’s 10 to watch. IEEE Software, 37 (4):C4, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:CGA

[Ano20b]  Anonymous.  Computer

Graphics and Applications subscribe. IEEE Software, 37(3):29, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:IAHa

[Ano20c]  Anonymous.  IEEE An-

nals of the History of Computing. IEEE Software, 37(3):53, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:IAHb

[Ano20d] Anonymous. IEEE Annals of the History of Computing. IEEE Software, 37(6):98, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICGc

[Ano20e] Anonymous. IEEE Computer Graphics and Applications magazine subscribe. IEEE Software, 37(6):75, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICGa

[Ano20f] Anonymous. [IEEE] Computer Graphics and Applications subscribe. IEEE Software, 37(3):29, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICGb

[Ano20g] Anonymous. IEEE Computer Graphics and Applications subscribe. IEEE Software, 37(5):42, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICE

[Ano20h] Anonymous. IEEE Computing Edge. IEEE Software, 37(6):C4, November/ December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:IOA

[Ano20i] Anonymous. IEEE Open Access Journal of the Computer Society. IEEE Software, 37(6):104, November/ December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:IPCb

[Ano20j] Anonymous. IEEE Pervasive Computing call for articles. IEEE Software, 37 (4):74, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ISPb

[Ano20k] Anonymous. [IEEE] Security & Privacy subscribe. IEEE Software, 37(3):13, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ISPc

[Ano20l] Anonymous. IEEE Security & Privacy subscribe. IEEE Software, 37(5):103, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ISPd

[Ano20m] Anonymous. IEEE Security & Privacy subscribe. IEEE Software, 37(6):11, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ISPa

[Ano20n]  Anonymous.  IEEE Se-

curity and Privacy magazine subscribe. IEEE Software, 37(1):80, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ISM

[Ano20o] Anonymous. IEEE Software magazine call for articles. IEEE Software, 37 (2):88, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ITBa

[Ano20p] Anonymous. [IEEE] Transactions on Big Data. IEEE Software, 37(1):67, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ITBb

[Ano20q] Anonymous. [IEEE] Transactions on Big Data. IEEE Software, 37(2):67, March/ April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ITBc

[Ano20r] Anonymous. [IEEE] Transactions on Big Data. IEEE Software, 37(5):76, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ITCa

[Ano20s] Anonymous. [IEEE] Transactions on Computers call for papers. IEEE Software, 37(1):49, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ITCb

[Ano20t] Anonymous. [IEEE] Transactions on Computers call for papers. IEEE Software, 37(2):77, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ITSa

[Ano20u] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 37(1):24, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ITSb

[Ano20v] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 37 (2):24, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ITSc

[Ano20w] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 37 (5):48, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:IPCa

[Ano20x] Anonymous. IT Professional call for articles. IEEE Software, 37(3):92, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:IPCc

[Ano20y] Anonymous. IT Professional call for articles. IEEE Software, 37(6):25, November/ December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:PCC

[Ano20z] Anonymous. Pervasive Computing call for articles. IEEE Software, 37(1):6, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:SPS

[Ano20-27] Anonymous. Security & Privacy subscribe. IEEE Software, 37(3):13, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:TBDa

[Ano20-28] Anonymous. Transactions on Big Data. IEEE Software, 37(1):67, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:TBDb

[Ano20-29] Anonymous. Transactions on Big Data. IEEE Software, 37(2):67, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:TCCa

[Ano20-30] Anonymous. Transactions on Computers call for papers. IEEE Software, 37(1): 49, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:TCCb

[Ano20-31] Anonymous. Transactions on Computers call for papers. IEEE Software, 37 (2):77, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:CSI

[Ano20-32] Anonymous. Computer Society information. IEEE Software, 37(2):C3, March/ April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:C

[Ano20-33]  Anonymous.  Correction.

IEEE Software, 37(3):92, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:Ea

[Ano20-34] Anonymous. Errata. IEEE Software, 37(2):94, March/ April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:Eb

[Ano20-35] Anonymous. Errata. IEEE Software, 37(4):86, July/ August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:FCa

[Ano20-36] Anonymous. [Front cover].

IEEE Software, 37(1):C1, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:FCb

[Ano20-37] Anonymous. Front cover. IEEE Software, 37(2):C1, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:FCc

[Ano20-38] Anonymous. Front cover. IEEE Software, 37(3):C1, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:FCd

[Ano20-39] Anonymous. Front cover. IEEE Software, 37(4):C1, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:FCe

[Ano20-40] Anonymous. Front cover. IEEE Software, 37(5):C1, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:FCf

[Ano20-41] Anonymous. Front cover. IEEE Software, 37(6):C1, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:HS

[Ano20-42] Anonymous. HOST Symposium. IEEE Software, 37(2):97, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:I

[Ano20-43] Anonymous. ICCQ. IEEE Software, 37(5):12, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSa

[Ano20-44] Anonymous. IEEE Computer Society. IEEE Software, 37(1):97, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSc

[Ano20-45] Anonymous. IEEE Computer Society benefits. IEEE Software, 37(2):12, March/ April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSh

[Ano20-46] Anonymous. IEEE Computer Society call for papers. IEEE Software, 37 (4):42, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSl

[Ano20-47] Anonymous. IEEE Computer Society call for papers. IEEE Software, 37(6):64, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSb

[Ano20-48] Anonymous. IEEE Computer Society information. IEEE Software, 37(1):C3, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSe

[Ano20-49] Anonymous. IEEE Computer Society information. IEEE Software, 37(3):C3, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSg

[Ano20-50] Anonymous. IEEE Computer Society information. IEEE Software, 37(4):25, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSk

[Ano20-51] Anonymous. IEEE Computer Society information. IEEE Software, 37(5):C3, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSn

[Ano20-52] Anonymous. IEEE Computer Society information. IEEE Software, 37(6):C3, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSd

[Ano20-53] Anonymous. IEEE Computer Society Job Boards. IEEE Software, 37(2):C4, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSi

[Ano20-54] Anonymous. IEEE Computer Society Job Boards. IEEE Software, 37(5):6, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSm

[Ano20-55] Anonymous. IEEE Computer Society Jobs Board. IEEE Software, 37(6):109, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSj

[Ano20-56] Anonymous. IEEE Computer Society member benefits. IEEE Software, 37 (5):121, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:ICSf

[Ano20-57] Anonymous. IEEE Computer Society Volunteer Service Awards. IEEE Software, 37(3):C4, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:IJB

[Ano20-58]  Anonymous.  IEEE Job

Boards. IEEE Software, 37(1):C4, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:IQWa

[Ano20-59] Anonymous. IEEE Quantum Week. IEEE Software, 37(3):97, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:IQWb

[Ano20-60] Anonymous. IEEE Quantum Week. IEEE Software, 37(5):C4, September/ October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:Ma

[Ano20-61]  Anonymous.  Masthead.

IEEE Software, 37(1):C2, January/February 2020.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:Mb

[Ano20-62]  Anonymous.  Masthead.

IEEE Software, 37(2):C2, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:Mc

[Ano20-63]  Anonymous.  Masthead.

IEEE Software, 37(3):C2, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:Md

[Ano20-64]  Anonymous.  Masthead.

IEEE Software, 37(4):C2, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:Me

[Ano20-65]  Anonymous.  Masthead.

IEEE Software, 37(5):C2, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:Mf

[Ano20-66]  Anonymous.  Masthead.

IEEE Software, 37(6):C2, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:SLa

[Ano20-67]  Anonymous.  Staff list-

ing. IEEE Software, 37(1):4– 5, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:SLb

[Ano20-68]  Anonymous.  Staff list-

ing. IEEE Software, 37 (2):4–5, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:SLc

[Ano20-69] Anonymous. Staff listing. IEEE Software, 37(3):4–5, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:SLd

[Ano20-70]  Anonymous.  Staff list-

ing. IEEE Software, 37 (4):4–5, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:SLe

[Ano20-71]  Anonymous.  Staff list-

ing. IEEE Software, 37(5):4– 5, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:SLf

[Ano20-72] Anonymous. Staff listing.

IEEE Software, 37(6):4–5, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:SC

[Ano20-73]  Anonymous.  Stay con-

nected. IEEE Software, 37(1):86, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:TCa

[Ano20-74] Anonymous. Table of contents. IEEE Software, 37 (1):1–2, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:TCb

[Ano20-75] Anonymous. Table of contents. IEEE Software, 37 (2):1–2, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:TCd

[Ano20-76] Anonymous. Table of contents. IEEE Software, 37 (3):1–2, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:TCe

[Ano20-77] Anonymous. Table of contents. IEEE Software, 37 (4):1–2, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:TCf

[Ano20-78] Anonymous. Table of contents. IEEE Software, 37 (5):1–2, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2020:TCg

[Ano20-79] Anonymous. Table of contents. IEEE Software, 37(6): 1–2, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:A

[Ano21a] Anonymous. Agile 2021. IEEE Software, 38(2):113, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:AA

[Ano21b]  Anonymous.  Agile al-

liance. IEEE Software, 38 (4):6, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:CSE

[Ano21c]  Anonymous.  Computing

in Science and Engineering. IEEE Software, 38 (3):136, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSo

[Ano21d]  Anonymous.  Computing

in Science and Engineering. IEEE Software, 38(5):82, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:Cb

[Ano21e] Anonymous. ComputingEdge.

IEEE Software, 38(6):109, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IAHa

[Ano21f] Anonymous. IEEE Annals of the History of Computing. IEEE Software, 38 (1):122, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IAHb

[Ano21g]  Anonymous.  IEEE An-

nals of the History of Computing. IEEE Software, 38(2):6, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IAHc

[Ano21h] Anonymous. IEEE Annals of the History of Computing. IEEE Software, 38(3): 6, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IAHd

[Ano21i] Anonymous. IEEE Annals of the History of Computing. IEEE Software, 38 (4):21, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IAHe

[Ano21j] Anonymous. IEEE Annals of the History of Computing. IEEE Software, 38(5):31, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICGb

[Ano21k] Anonymous. IEEE Computer Graphics & Applications. IEEE Software, 38 (2):37, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICGc

[Ano21l] Anonymous. IEEE Computer Graphics & Applications. IEEE Software, 38 (4):27, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICGd

[Ano21m] Anonymous. IEEE Computer Graphics & Applications. IEEE Software, 38 (5):38, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICGa

[Ano21n] Anonymous. IEEE Computer Graphics and Applications subscribe. IEEE Software, 38(1):113, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICEa

[Ano21o] Anonymous. IEEE Computing Edge. IEEE Software, 38(1):C4, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICEb

[Ano21p] Anonymous. IEEE Computing Edge. IEEE Software, 38(3):131, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICEc

[Ano21q] Anonymous. IEEE Computing Edge. IEEE Software, 38 (4):114, July/August 2021.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSk

[Ano21r] Anonymous. IEEE Computing in Science & Engineering subscription. IEEE Software, 38(4):51, July/ August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSg

[Ano21s] Anonymous. IEEE Computing in Science and Engineering. IEEE Software, 38(3):34, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IIPa

[Ano21t]  Anonymous.  IEEE IT

Professional call for articles. IEEE Software, 38 (2):95, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IIPb

[Ano21u] Anonymous. IEEE IT Professional call for articles. IEEE Software, 38(3):71, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IOJa

[Ano21v] Anonymous. IEEE Open Journal of the Computer Society. IEEE Software, 38 (2):133, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IOJb

[Ano21w] Anonymous. IEEE Open Journal of the Computer Society. IEEE Software, 38 (4):125, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IOJc

[Ano21x] Anonymous. IEEE Open Journal of the Computer Society. IEEE Software, 38 (5):124, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IOJd

[Ano21y] Anonymous. IEEE Open Journal of the Computer Society. IEEE Software, 38 (6):16, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IPCa

[Ano21z] Anonymous. IEEE Pervasive Computing. IEEE Software, 38(1):53, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IPCc

[Ano21-27] Anonymous. IEEE Pervasive Computing call for articles. IEEE Software, 38 (2):106, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IPCd

[Ano21-28] Anonymous. IEEE Pervasive Computing call for articles. IEEE Software, 38 (5):123, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IPCe

[Ano21-29] Anonymous. IEEE Pervasive Computing call for articles. IEEE Software, 38 (6):41, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ISPb

[Ano21-30] Anonymous. IEEE Security & Privacy subscribe. IEEE Software, 38(2):112, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ISPa

[Ano21-31] Anonymous. IEEE Security and Privacy subscription. IEEE Software, 38(1):

15, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ITBa

[Ano21-32] Anonymous. [IEEE] Transactions on Big Data. IEEE Software, 38(1):24, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ITBb

[Ano21-33] Anonymous. [IEEE] Transactions on Big Data. IEEE Software, 38(2):118, March/ April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ITBc

[Ano21-34] Anonymous. IEEE Transactions on Big Data. IEEE Software, 38(3):141, May/ June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ITBd

[Ano21-35] Anonymous. IEEE Transactions on Big Data. IEEE Software, 38(5):22, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ITCa

[Ano21-36] Anonymous. IEEE Transactions on Computing. IEEE Software, 38(1):80, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ITSa

[Ano21-37] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 38(1):47, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ITSb

[Ano21-38] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 38 (2):132, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ITSc

[Ano21-39] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 38 (4):113, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ITSd

[Ano21-40] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 38 (6):97, November/December 2021. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IPCb

[Ano21-41] Anonymous. IT Professional call for articles. IEEE Software, 38(1):74, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ITCb

[Ano21-42] Anonymous. Transactions on Computers call for papers. IEEE Software, 38 (6):64, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:CMA

[Ano21-43] Anonymous. Call for 2021 major awards nominations. IEEE Software, 38(3):125, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:CEA

[Ano21-44] Anonymous. Call for education awards nominations. IEEE Software, 38(3):25, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:Ca

[Ano21-45]  Anonymous.  Correction.

IEEE Software, 38(6):102, November/December 2021.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:E

[Ano21-46]  Anonymous.  Erratum.

IEEE Software, 38(2):127, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:FCa

[Ano21-47] Anonymous. Front cover. IEEE Software, 38(1):C1, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:FCb

[Ano21-48] Anonymous. Front cover. IEEE Software, 38(2):C1, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:FCc

[Ano21-49] Anonymous. Front cover. IEEE Software, 38(3):C1, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:FCd

[Ano21-50] Anonymous. [Front cover]. IEEE Software, 38(4):C1, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:FCe

[Ano21-51] Anonymous. [Front cover].

IEEE Software, 38(5):C1, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:FCf

[Ano21-52] Anonymous. [Front cover].

IEEE Software, 38(6):C1, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:I

[Ano21-53] Anonymous. ICCQ. IEEE Software, 38(5):15, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSa

[Ano21-54] Anonymous. IEEE Computer Society call for papers. IEEE Software, 38(1): 12, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSd

[Ano21-55] Anonymous. IEEE Computer Society call for papers. IEEE Software, 38 (2):44, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSj

[Ano21-56] Anonymous. IEEE Computer Society call for papers. IEEE Software, 38 (4):15, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSp

[Ano21-57] Anonymous. IEEE Computer Society call for papers. IEEE Software, 38(5):100, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSt

[Ano21-58] Anonymous. IEEE Computer Society call for papers. IEEE Software, 38(6):15, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSn

[Ano21-59] Anonymous. IEEE Computer Society election. IEEE Software, 38(5):23, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSc

[Ano21-60] Anonymous. IEEE Computer Society information. IEEE Software, 38(1):C3, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSe

[Ano21-61] Anonymous. IEEE Computer Society information. IEEE Software, 38(2):C3, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSh

[Ano21-62] Anonymous. IEEE Computer Society information. IEEE Software, 38(3):C3, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSl

[Ano21-63] Anonymous. IEEE Computer Society information. IEEE Software, 38(4):C3, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSr

[Ano21-64] Anonymous. IEEE Computer Society information. IEEE Software, 38(5):C3, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSv

[Ano21-65] Anonymous. IEEE Computer Society information. IEEE Software, 38(6):C3, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSq

[Ano21-66] Anonymous. IEEE Computer Society Job Boards. IEEE Software, 38(5):117, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSb

[Ano21-67] Anonymous. IEEE Computer Society member benefits. IEEE Software, 38(1): 87, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSf

[Ano21-68] Anonymous. IEEE Computer Society member benefits. IEEE Software, 38 (2):C4, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSi

[Ano21-69] Anonymous. IEEE Computer Society member benefits. IEEE Software, 38 (3):C4, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSm

[Ano21-70] Anonymous. IEEE Computer Society member benefits. IEEE Software, 38 (4):C4, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSs

[Ano21-71] Anonymous. IEEE Computer Society member benefits. IEEE Software, 38 (5):C4, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:ICSu

[Ano21-72] Anonymous. IEEE Computer Society member benefits. IEEE Software, 38(6): 103, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IIPc

[Ano21-73]  Anonymous.  IEEE IT

Professional call for articles. IEEE Software, 38 (4):93, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IJBb

[Ano21-74]  Anonymous.  IEEE Job

Boards. IEEE Software, 38 (6):C4, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:IJBa

[Ano21-75] Anonymous. IEEE Jobs Board. IEEE Software, 38 (4):121, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:Ma

[Ano21-76]  Anonymous.  Masthead.

IEEE Software, 38(1):C2, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:Mb

[Ano21-77]  Anonymous.  Masthead.

IEEE Software, 38(2):C2, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:Mc

[Ano21-78]  Anonymous.  Masthead.

IEEE Software, 38(3):C2, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:Md

[Ano21-79]  Anonymous.  Masthead.

IEEE Software, 38(4):C2, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:Me

[Ano21-80]  Anonymous.  Masthead.

IEEE Software, 38(5):C2, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:Mf

[Ano21-81]  Anonymous.  Masthead.

IEEE Software, 38(6):C2, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:SLa

[Ano21-82]  Anonymous.  Staff list-

ing. IEEE Software, 38(1):4– 5, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:SLb

[Ano21-83]  Anonymous.  Staff list-

ing. IEEE Software, 38 (2):4–5, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:SLc

[Ano21-84] Anonymous. Staff listing.

IEEE Software, 38(3):4–5, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:SLd

[Ano21-85]  Anonymous.  Staff list-

ing. IEEE Software, 38 (4):4–5, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:SLe

[Ano21-86]  Anonymous.  Staff list-

ing. IEEE Software, 38(5):4– 5, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:SLf

[Ano21-87] Anonymous. Staff listing. IEEE Software, 38(6):4–5, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:TCa

[Ano21-88] Anonymous. Table of contents. IEEE Software, 38 (1):1–2, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:TCb

[Ano21-89] Anonymous. Table of contents. IEEE Software, 38 (2):1–2, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:TCc

[Ano21-90] Anonymous. Table of contents. IEEE Software, 38 (3):1–2, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:TCd

[Ano21-91] Anonymous. Table of contents. IEEE Software, 38 (4):1–2, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:TCe

[Ano21-92] Anonymous. Table of contents. IEEE Software, 38 (5):1–2, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2021:TCf

[Ano21-93] Anonymous. Table of contents. IEEE Software, 38(6): 1–2, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IWC

[Ano22a] Anonymous. 2022 IEEE World Congress on Services. IEEE Software, 39 (2):6, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:AA

[Ano22b]  Anonymous.  Agile al-

liance. IEEE Software, 39 (4):7, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:C

[Ano22c] Anonymous. ComputingEdge. IEEE Software, 39(1):30, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICC

[Ano22d] Anonymous. IEEE Computer call for special issue proposals. IEEE Software, 39(2):111, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICEa

[Ano22e] Anonymous. IEEE Computing Edge. IEEE Software, 39(2):100, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICEc

[Ano22f] Anonymous. IEEE Computing Edge. IEEE Software, 39(6):C2, November/ December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICEb

[Ano22g] Anonymous. IEEE Computing Edge filler. IEEE Software, 39(3):21, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSk

[Ano22h] Anonymous. IEEE Computing in Science and Engineering. IEEE Software, 39(6): 108, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IIPa

[Ano22i]  Anonymous.  IEEE IT

Professional call for articles. IEEE Software, 39(2): 19, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IIPb

[Ano22j] Anonymous. IEEE IT Professional call for articles. IEEE Software, 39(6):104, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IPCa

[Ano22k] Anonymous. IEEE Pervasive Computing call for articles. IEEE Software, 39 (2):38, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IPCb

[Ano22l] Anonymous. IEEE Pervasive Computing filler. IEEE Software, 39(4):83, July/ August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ISPa

[Ano22m] Anonymous. IEEE Security and Privacy magazine subscribe. IEEE Software, 39(1):54, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ISPb

[Ano22n]  Anonymous.  IEEE Se-

curity and Privacy subscribe. IEEE Software, 39 (2):46, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ISPc

[Ano22o] Anonymous. IEEE Security and Privacy subscribe. IEEE Software, 39(6):57, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ITBb

[Ano22p] Anonymous. IEEE Trans on Big Data filler. IEEE Software, 39(3):42, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ITBc

[Ano22q] Anonymous. IEEE Trans on Big Data filler. IEEE Software, 39(4):47, July/ August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ITCb

[Ano22r] Anonymous. IEEE Trans on Computers generic cfp filler. IEEE Software, 39(3):20, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ITSc

[Ano22s] Anonymous. IEEE Trans On Sustainable Computing filler. IEEE Software, 39(3):11, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ITBa

[Ano22t] Anonymous. IEEE Transactions on Big Data. IEEE Software, 39(1):69, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ITCa

[Ano22u] Anonymous. IEEE Transactions on Computers call for papers. IEEE Software, 39(1):92, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ITCc

[Ano22v] Anonymous. IEEE Transactions on Computers call for papers. IEEE Software, 39(6):50, November/ December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ITSa

[Ano22w] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 39(1):104, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ITSb

[Ano22x] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 39 (2):79, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:CE

[Ano22y]  Anonymous.  Computing

edge. IEEE Software, 39 (4):C2, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:FCa

[Ano22z] Anonymous. [Front cover]. IEEE Software, 39(1):C1, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:FCb

[Ano22-27] Anonymous. Front cover.

IEEE Software, 39(2):C1, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:FCc

[Ano22-28] Anonymous. Front cover.

IEEE Software, 39(3):C1, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:FCd

[Ano22-29] Anonymous. Front cover.

IEEE Software, 39(4):C1, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:FCe

[Ano22-30] Anonymous. Front cover.

IEEE Software, 39(6):C1, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSe

[Ano22-31] Anonymous. IEEE Computer Society. IEEE Software, 39(2):C3, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSa

[Ano22-32] Anonymous. IEEE Computer Society call for papers. IEEE Software, 39(1): 28, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSd

[Ano22-33] Anonymous. IEEE Computer Society call for papers. IEEE Software, 39 (2):15, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSh

[Ano22-34] Anonymous. IEEE Computer Society call for papers. IEEE Software, 39 (4):17, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSj

[Ano22-35] Anonymous. IEEE Computer Society call for papers. IEEE Software, 39(6):27, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSf

[Ano22-36] Anonymous. IEEE Computer Society CFP filler. IEEE Software, 39(3):6, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSb

[Ano22-37] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. IEEE Software, 39(1):111, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSc

[Ano22-38] Anonymous. IEEE Computer Society information. IEEE Software, 39(1):C3, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSg

[Ano22-39] Anonymous. IEEE Computer Society information. IEEE Software, 39(3):C3, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSi

[Ano22-40] Anonymous. IEEE Computer Society information. IEEE Software, 39(4):96, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICSl

[Ano22-41] Anonymous. IEEE Computer Society information. IEEE Software, 39(6):C3, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ICJ

[Ano22-42] Anonymous. IEEE CS Jobs Board. IEEE Software, 39 (4):C4, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IGF

[Ano22-43] Anonymous. IEEE general filler. IEEE Software, 39(3):57, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IJBb

[Ano22-44] Anonymous. IEEE Jobs Board filler. IEEE Software, 39(3):C4, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IJBa

[Ano22-45] Anonymous. IEEE Jobs Boards. IEEE Software, 39(1):C4, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IJ

[Ano22-46] Anonymous. IEEE Jobsboard. IEEE Software, 39 (2):C4, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IOJa

[Ano22-47] Anonymous. IEEE open journal computer filler. IEEE Software, 39(3):125, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IOJb

[Ano22-48] Anonymous. IEEE open journal CS. IEEE Software, 39(4):C3, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IPG

[Ano22-49] Anonymous. IEEE PVC general filler. IEEE Software, 39(3):115, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ISR

[Ano22-50]  Anonymous.  IEEE SP

over the rainbow filler. IEEE Software, 39(3):119, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ISSb

[Ano22-51] Anonymous. IEEE SP subscribe. IEEE Software, 39 (4):62, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ISSa

[Ano22-52] Anonymous. IEEE SP subscribe filler. IEEE Software, 39(3):105, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:ITF

[Ano22-53] Anonymous. IEEE tsusc filler. IEEE Software, 39 (4):33, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:IC

[Ano22-54] Anonymous. Ivica Crnković: 1955–2022. IEEE Software, 39(3):7–8, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:JB

[Ano22-55] Anonymous. Job boards. IEEE Software, 39(6):C4, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:L

[Ano22-56] Anonymous. Letters. IEEE Software, 39(3):9–11, May/ June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:Ma

[Ano22-57]  Anonymous.  Masthead.

IEEE Software, 39(1):C2, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:Mb

[Ano22-58]  Anonymous.  Masthead.

IEEE Software, 39(2):C2, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:Mc

[Ano22-59]  Anonymous.  Masthead.

IEEE Software, 39(3):C2, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:Md

[Ano22-60]  Anonymous.  Masthead.

IEEE Software, 39(4):1, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:Me

[Ano22-61]  Anonymous.  Masthead.

IEEE Software, 39(6):1, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:RP

[Ano22-62] Anonymous. Over the Rainbow podcast. IEEE Software, 39(6):98, November/ December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:RLA

[Ano22-63] Anonymous. Rebuttal letter by Alain Abran (Ph.D.). IEEE Software, 39(1):29–30, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). See [BC21].

Anonymous:2022:SLa

[Ano22-64] Anonymous. Staff listing. IEEE Software, 39(1):4–5, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:SLb

[Ano22-65]  Anonymous.  Staff list-

ing. IEEE Software, 39(2): 4–5, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:SLc

[Ano22-66] Anonymous. Staff listing. IEEE Software, 39(3):4–5, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:SLd

[Ano22-67]  Anonymous.  Staff list-

ing. IEEE Software, 39 (4):5–6, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:SLe

[Ano22-68] Anonymous. Staff listing.

IEEE Software, 39(6):5–6, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:TCa

[Ano22-69] Anonymous. Table of contents. IEEE Software, 39(1):1–2, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:TCb

[Ano22-70] Anonymous. Table of contents. IEEE Software, 39(2): 1–2, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:TCc

[Ano22-71] Anonymous. Table of contents. IEEE Software, 39 (3):1–2, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:TCd

[Ano22-72] Anonymous. Table of contents. IEEE Software, 39 (4):2–3, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:TCe

[Ano22-73] Anonymous. Table of contents. IEEE Software, 39(6): 2–3, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2022:WH

[Ano22-74] Anonymous. Watts Humphrey. IEEE Software, 39(4):26, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICA

[Ano23a] Anonymous. 2023 IEEE Conference on Artificial Intelligence (CAI). IEEE Software, 40(3):C2, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:CE

[Ano23b]  Anonymous.  Computing

Edge. IEEE Software, 40 (4):C4, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:CSE

[Ano23c]  Anonymous.  Computing

in Science & Engineering. IEEE Software, 40(3):88, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:C

[Ano23d] Anonymous. ComputingEdge.

IEEE Software, 40(6):C4, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IAHa

[Ano23e] Anonymous. IEEE Annals of the History of Computing. IEEE Software, 40(5):57, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IAHb

[Ano23f] Anonymous. IEEE Annals of the History of Computing. IEEE Software, 40(6):108, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSp

[Ano23g] Anonymous. IEEE Computer seeks 2025–2026 Editor in Chief. IEEE Software, 40(5):125, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICEa

[Ano23h] Anonymous. IEEE Computing Edge. IEEE Software, 40(1):C2, January/ February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICEb

[Ano23i] Anonymous. IEEE Computing Edge. IEEE Software, 40(2):C4, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICEc

[Ano23j] Anonymous. IEEE Computing Edge. IEEE Software, 40(3):C4, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICEd

[Ano23k] Anonymous. IEEE Computing Edge. IEEE Software, 40(5):C4, September/ October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSb

[Ano23l] Anonymous. IEEE Computing in Science and Engineering. IEEE Software, 40(1): 85, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSv

[Ano23m] Anonymous. IEEE Computing in Science and Engineering. IEEE Software, 40(6): 125, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IIPa

[Ano23n]  Anonymous.  IEEE IT

Professional call for articles. IEEE Software, 40(1): 76, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IIPb

[Ano23o] Anonymous. IEEE IT Professional call for articles. IEEE Software, 40(2):102, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IOJa

[Ano23p] Anonymous. IEEE Open Journal of the Computer Society. IEEE Software, 40(1): 41, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IOJb

[Ano23q] Anonymous. IEEE Open Journal of the Computer Society. IEEE Software, 40 (4):100, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IOJc

[Ano23r] Anonymous. IEEE Open Journal of the Computer Society. IEEE Software, 40(6): 148, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IPCa

[Ano23s] Anonymous. IEEE Pervasive Computing call for articles. IEEE Software, 40(1): 40, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IPCb

[Ano23t] Anonymous. IEEE Pervasive Computing call for articles. IEEE Software, 40 (2):110, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IPCc

[Ano23u] Anonymous. IEEE Pervasive Computing: Call for articles. IEEE Software, 40(4):87, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ISPa

[Ano23v] Anonymous. IEEE Security and Privacy subscribe. IEEE Software, 40(1):28, January/ February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ISPc

[Ano23w] Anonymous. IEEE Security and Privacy subscribe. IEEE Software, 40(6):54, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ISPb

[Ano23x] Anonymous. IEEE Security and Privacy subscription. IEEE Software, 40 (2):89, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ITBa

[Ano23y] Anonymous. IEEE Transactions on Big Data. IEEE Software, 40(1):22, January/ February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ITBb

[Ano23z] Anonymous. IEEE Transactions on Big Data. IEEE Software, 40(2):79, March/ April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ITC

[Ano23-27] Anonymous. IEEE Transactions on Computers call for papers. IEEE Software, 40(2):59, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ITSa

[Ano23-28] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 40(1):17, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ITSb

[Ano23-29] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 40 (2):49, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ITSc

[Ano23-30] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Software, 40 (6):138, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ITA

[Ano23-31] Anonymous. IEEE Transactions on the Annals of Computing. IEEE Software, 40 (2):116, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IPCd

[Ano23-32] Anonymous. IT Professional: Call for articles. IEEE Software, 40(4):99, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:FCa

[Ano23-33] Anonymous. Front cover.

IEEE Software, 40(1):C1, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:FCb

[Ano23-34] Anonymous. Front cover.

IEEE Software, 40(2):C1, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:FCc

[Ano23-35] Anonymous. Front cover.

IEEE Software, 40(3):C1, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:FCd

[Ano23-36] Anonymous. Front cover.

IEEE Software, 40(4):C1, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:FCe

[Ano23-37] Anonymous. Front cover. IEEE Software, 40(5):C1, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:FCf

[Ano23-38] Anonymous. Front cover. IEEE Software, 40(6):C1, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:HAa

[Ano23-39]  Anonymous.  Humphrey

Award. IEEE Software, 40(3):33, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:HAb

[Ano23-40]  Anonymous.  Humphrey

Award. IEEE Software, 40 (4):38, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IIPc

[Ano23-41] Anonymous. IEEE IT Professional call for articles. IEEE Software, 40(5):84, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICC

[Ano23-42]  Anonymous.  IEEE Ca-

reer Center. IEEE Software, 40(5):129, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICGb

[Ano23-43] Anonymous. IEEE computer graphics & applications. IEEE Software, 40(6): 116, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSn

[Ano23-44] Anonymous. IEEE Computer Society. IEEE Software, 40(5):40, September/ October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSw

[Ano23-45] Anonymous. IEEE Computer Society. IEEE Software, 40(6):143, November/ December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSu

[Ano23-46] Anonymous. IEEE Computer Society — call for papers. IEEE Software, 40 (6):39, November/December 2023. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSt

[Ano23-47] Anonymous. IEEE Computer Society — publications seek 2025 Editors in Chief. IEEE Software, 40 (6):35, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSa

[Ano23-48] Anonymous. IEEE Computer Society call for papers. IEEE Software, 40(1): 8, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSe

[Ano23-49] Anonymous. IEEE Computer Society call for papers. IEEE Software, 40 (2):7, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSh

[Ano23-50] Anonymous. IEEE Computer Society: Call for papers. IEEE Software, 40(4):9, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSi

[Ano23-51] Anonymous. IEEE Computer Society Career Center. IEEE Software, 40 (4):10, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSz

[Ano23-52] Anonymous. IEEE Computer Society Career Center. IEEE Software, 40(6):C3, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSl

[Ano23-53] Anonymous. IEEE Computer Society D&I fund. IEEE Software, 40(4):117, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSy

[Ano23-54] Anonymous. IEEE Computer Society diversity and inclusion. IEEE Software, 40(6):C2, November/ December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSq

[Ano23-55] Anonymous. IEEE Computer Society drive diversity and inclusion in computing. IEEE Software, 40(5):C2, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSj

[Ano23-56] Anonymous. IEEE Computer Society has you covered! IEEE Software, 40 (4):18, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSs

[Ano23-57] Anonymous. IEEE Computer Society has you covered! IEEE Software, 40 (6):26, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSc

[Ano23-58] Anonymous. IEEE Computer Society information. IEEE Software, 40(1):C3, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSf

[Ano23-59] Anonymous. IEEE Computer Society information. IEEE Software, 40(2):C3, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSg

[Ano23-60] Anonymous. IEEE Computer Society information. IEEE Software, 40(3):C3, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSm

[Ano23-61] Anonymous. IEEE Computer Society information. IEEE Software, 40(4):C3, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSr

[Ano23-62] Anonymous. IEEE Computer Society information. IEEE Software, 40(5):C3, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSo

[Ano23-63] Anonymous. IEEE Computer Society member benefits. IEEE Software, 40 (5):116, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSd

[Ano23-64] Anonymous. IEEE Computer Society member benifits. IEEE Software, 40(2):C2, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSk

[Ano23-65] Anonymous. IEEE Computer Society volunteer service awards. IEEE Software, 40(4):45, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICSx

[Ano23-66] Anonymous. IEEE Computer Society Volunteer Service awards. IEEE Software, 40(6):147, November/ December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ICGa

[Ano23-67]  Anonymous.  IEEE CS

generic. IEEE Software, 40(3):71, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IC

[Ano23-68] Anonymous. IEEE CSCFP. IEEE Software, 40(3):53, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IJBa

[Ano23-69]  Anonymous.  IEEE Job

Boards. IEEE Software, 40(1):C4, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IJBb

[Ano23-70] Anonymous. IEEE Jobs Board. IEEE Software, 40 (2):117, March/April 2023.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IRPa

[Ano23-71] Anonymous. IEEE over the rainbow podcast. IEEE Software, 40(2):98, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IRPb

[Ano23-72] Anonymous. IEEE over the rainbow podcast. IEEE Software, 40(3):83, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:IQW

[Ano23-73] Anonymous. IEEE quantum week. IEEE Software, 40(4):C2, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:ISS

[Ano23-74] Anonymous. IEEE SP subscribe. IEEE Software, 40(3):94, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:Ma

[Ano23-75]  Anonymous.  Masthead.

IEEE Software, 40(1):3, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:Mb

[Ano23-76]  Anonymous.  Masthead.

IEEE Software, 40(2):3, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:Mc

[Ano23-77]  Anonymous.  Masthead.

IEEE Software, 40(3):3, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:Md

[Ano23-78]  Anonymous.  Masthead.

IEEE Software, 40(4):3, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:Me

[Ano23-79]  Anonymous.  Masthead.

IEEE Software, 40(5):3, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:Mf

[Ano23-80]  Anonymous.  Masthead.

IEEE Software, 40(6):3, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:R

[Ano23-81] Anonymous. Over the rainbow. IEEE Software, 40 (4):24, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:RCS

[Ano23-82] Anonymous. Over the Rainbow: 21st century security & privacy podcast. IEEE Software, 40(6):34, November/ December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:RPa

[Ano23-83] Anonymous. Over the Rainbow podcast. IEEE Software, 40(1):49, January/ February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:RPb

[Ano23-84] Anonymous. Over the Rainbow podcast. IEEE Software, 40(5):49, September/ October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:SLa

[Ano23-85]  Anonymous.  Staff list-

ing. IEEE Software, 40(1):5– 6, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:SLb

[Ano23-86]  Anonymous.  Staff list-

ing. IEEE Software, 40 (2):5–6, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:SLc

[Ano23-87] Anonymous. Staff listing. IEEE Software, 40(3):5–6, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:SLd

[Ano23-88]  Anonymous.  Staff list-

ing. IEEE Software, 40 (4):5–6, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:SLe

[Ano23-89]  Anonymous.  Staff list-

ing. IEEE Software, 40(5):5– 6, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:SLf

[Ano23-90] Anonymous. Staff listing. IEEE Software, 40(6):5–6, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:TCa

[Ano23-91] Anonymous. Table of contents. IEEE Software, 40 (1):1–2, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:TCb

[Ano23-92] Anonymous. Table of contents. IEEE Software, 40 (2):1–2, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:TCc

[Ano23-93] Anonymous. Table of contents. IEEE Software, 40 (3):1–2, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:TCd

[Ano23-94] Anonymous. Table of contents. IEEE Software, 40 (4):1–2, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:TCe

[Ano23-95] Anonymous. Table of contents. IEEE Software, 40 (5):1–2, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2023:TCf

[Ano23-96] Anonymous. Table of contents. IEEE Software, 40(6): 1–2, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:CSE

[Ano24a]  Anonymous.  Computing

in Science & Engineering. IEEE Software, 41(3):79, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:IAH

[Ano24b]  Anonymous.  IEEE An-

nals of the History of Computing. IEEE Software, 41(3):66, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICGb

[Ano24c] Anonymous. IEEE Computer Graphics and Applications. IEEE Software, 41(3):75, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICGa

[Ano24d] Anonymous. IEEE Computers Graphics & Applications. IEEE Software, 41 (2):112, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICEb

[Ano24e] Anonymous. IEEE Computing Edge. IEEE Software, 41(1):C4, January/ February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICEc

[Ano24f] Anonymous. IEEE Computing Edge. IEEE Software, 41(2):C4, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICEd

[Ano24g] Anonymous. IEEE Computing Edge. IEEE Software, 41(3):C4, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ISPb

[Ano24h] Anonymous. IEEE Security & Privacy subscribe. IEEE Software, 41(1):54, January/ February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ISCb

[Ano24i] Anonymous. IEEE Software — call for special issue proposals. IEEE Software, 41(3):19, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ITB

[Ano24j] Anonymous. IEEE Transactions on Big Data. IEEE Software, 41(3):89, May/ June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ITP

[Ano24k] Anonymous. IEEE Transactions on Privacy. IEEE Software, 41(3):48, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:FCb

[Ano24l] Anonymous. Front cover. IEEE Software, 41(1):C1, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:FCc

[Ano24m] Anonymous. Front cover. IEEE Software, 41(2):C1, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:FCd

[Ano24n] Anonymous. Front cover. IEEE Software, 41(3):C1, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:IIPb

[Ano24o]  Anonymous.  IEEE IT

Professional call for articles. IEEE Software, 41(1): 44, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSr

[Ano24p] Anonymous. IEEE Computer Society — call for papers. IEEE Software, 41(3):6, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSf

[Ano24q] Anonymous. IEEE Computer Society benefits. IEEE Software, 41(1):137, January/February 2024. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSd

[Ano24r] Anonymous. IEEE Computer Society call for papers. IEEE Software, 41(1): 49, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSm

[Ano24s] Anonymous. IEEE Computer Society call for papers. IEEE Software, 41 (2):16, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSl

[Ano24t] Anonymous. IEEE Computer Society Career Center. IEEE Software, 41(1): C3, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSq

[Ano24u] Anonymous. IEEE Computer Society Career Center. IEEE Software, 41 (2):C3, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSw

[Ano24v] Anonymous. IEEE Computer Society Career Center. IEEE Software, 41(3):C3, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSj

[Ano24w] Anonymous. IEEE Computer Society diversity & inclusion. IEEE Software, 41(1):C2, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSp

[Ano24x] Anonymous. IEEE Computer Society diversity and inclusion. IEEE Software, 41(2):C2, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSv

[Ano24y] Anonymous. IEEE Computer Society diversity and inclusion fund. IEEE Software, 41(3):C2, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSs

[Ano24z] Anonymous. IEEE Computer Society has you covered! IEEE Software, 41(3):25, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSh

[Ano24-27] Anonymous. IEEE Computer Society information. IEEE Software, 41(1):145, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSo

[Ano24-28] Anonymous. IEEE Computer Society information. IEEE Software, 41(2):103, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSt

[Ano24-29] Anonymous. IEEE Computer Society information. IEEE Software, 41(3):90, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSn

[Ano24-30] Anonymous. IEEE Computer Society member benefits. IEEE Software, 41 (2):76, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSb

[Ano24-31] Anonymous. IEEE Computer Society search for an Editor In Chief. IEEE Software, 41(1):8, January/ February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ICSu

[Ano24-32] Anonymous. IEEE Computer Society Volunteer Service awards. IEEE Software, 41(3):95, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:IRPb

[Ano24-33] Anonymous. IEEE Over the Rainbow podcast. IEEE Software, 41(1):30, January/ February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:IRPc

[Ano24-34] Anonymous. IEEE over the rainbow podcast. IEEE Software, 41(2):97, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:ISCa

[Ano24-35] Anonymous. IEEE software call for proposals. IEEE Software, 41(2):98, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:IVSb

[Ano24-36] Anonymous. IEEE Volunteer Service awards. IEEE Software, 41(1):132, January/February 2024. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:IVSc

[Ano24-37] Anonymous. IEEE Volunteer Service awards. IEEE Software, 41(2):108, March/ April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:Mb

[Ano24-38]  Anonymous.  Masthead.

IEEE Software, 41(1):3, January/February 2024.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:Mc

[Ano24-39]  Anonymous.  Masthead.

IEEE Software, 41(2):3, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:Md

[Ano24-40]  Anonymous.  Masthead.

IEEE Software, 41(3):3, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:SLb

[Ano24-41]  Anonymous.  Staff list-

ing. IEEE Software, 41(1):5– 6, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:SLc

[Ano24-42]  Anonymous.  Staff list-

ing. IEEE Software, 41 (2):5, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:SLd

[Ano24-43]  Anonymous.  Staff list-

ing. IEEE Software, 41(3): 5, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:TCb

[Ano24-44] Anonymous. Table of contents. IEEE Software, 41 (1):1–2, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:TCc

[Ano24-45] Anonymous. Table of contents. IEEE Software, 41 (2):1–2, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2024:TCd

[Ano24-46] Anonymous. Table of contents. IEEE Software, 41 (3):1–2, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Arya:2022:DCS

[ANR22] Deeksha M. Arya, Mathieu Nassif, and Martin P. Robillard. A data-centric study of software tutorial design. IEEE Software, 39(3): 106–115, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Antinyan:2020:EEA

[Ant20] Vard Antinyan. Evaluating essential and accidental code complexity triggers by practitioners perception. IEEE Software, 37(6):86–93, November/December 2020.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Antinyan:2023:SLL

[Ant23] Vard Antinyan. Seven lessons learned from automotive software supplier collaborations. IEEE Software, 40(1):77–85, January/ February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Arcuri:2021:ABW

[Arc21] Andrea Arcuri. Automated black- and white-box testing of RESTful APIs with Evo-Master. IEEE Software, 38 (3):72–78, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Antinyan:2021:SSA

[AS21] Vard Antinyan and Henrik Sandgren. Software safety analysis to support ISO 26262-6 compliance in agile development. IEEE Software, 38(3):52–60, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Abrahao:2024:RHEb

[ASB^{+}24] Silvia Abrahão, Miroslaw Staron, Maria Teresa Baldassarre, Jennifer Horkoff, Birgit Penzenstadler, Paul Ralph, and Alexander Serebrenik. Research highlights in evidence-based software engineering. IEEE Software, 41(1):133–136, January/February 2024. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Abrahao:2024:MAC

[ASIM24] Silvia Abrahão, Miroslaw Staron, Emilio Insfran, and Henry Muccini. Modeling and architecting of complex software systems. IEEE Software, 41(3):76–79, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Abrahao:2023:OSS

[ASS^{+}23a] Silvia Abrahão, Miroslaw Staron, Alexander Serebrenik, Birgit Penzenstadler, and Rafael Capilla. Open source software: Communities and quality. IEEE Software, 40(4):96–99, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Abrahao:2023:FDE

[ASS^{+}23b] Silvia Abrahão, Miroslaw Staron, Alexander Serebrenik, Birgit Penzenstadler, Rafael Prikladnicki, and Henry Muccini. Focusing on developers in the era of AI and ML. IEEE Software, 40(6):126–129, November/ December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Avgeriou:2021:OCT

[ATA^{+}21] Paris C. Avgeriou, Davide Taibi, Apostolos Ampatzoglou, Francesca Arcelli Fontana, Terese Besker, Alexander Chatzigeorgiou, Valentina Lenarduzzi, Antonio Martini, Athanasia Moschou, Ilaria Pigazzini, Nyyti Saarimaki, Darius Daniel Sas, Saulo Soares de Toledo, and Angeliki Agathi Tsintzira. An overview and comparison of technical debt measurement tools. IEEE Software, 38 (3):61–71, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Araujo:2020:TGH

[ATS20] Andre Araujo, Valeria Times, and Marcus Silva. A tool for generating health applications using archetypes. IEEE Software, 37(1):60– 67, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Alorwu:2022:RUC

[AVvBH22] Andy Alorwu, Aku Visuri, Niels van Berkel, and Simo Hosio. (Re)Using crowdsourced health data: Perceptions of data contributors. IEEE Software, 39 (1):36–42, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Borg:2023:HWM

[BAEM23] Markus Borg, Emil Aasa, Khashayar Etemadi, and Martin Monperrus. Human, what must I tell you? IEEE Software, 40(3):9–14, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Berenguer:2024:RDI

[BATM24] Alberto Berenguer, Olimpia Alcaraz, David Tomás, and Jose-Norberto Mazón. From research on data-intensive software to innovation in data spaces: a search service for tabular data. IEEE Software, 41(3):59–66, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bernstein:2020:URU

[BB20] Dave Bernstein and Grady Booch. The UML and the rational unified process. IEEE Software, 37(6):12, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bernardino:2021:IPG

[BBM^{+}21] Matheus Tavares Bernardino, Giuliano Belinassi, Paulo Meirelles, Eduardo Martins Guerra, and Alfredo Goldman. Improving parallelism in Git and GNU Compiler Collection: Strategies, difficulties, and lessons learned. IEEE Software, 38(5):92–100, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ben-Cnaan:2021:UAR

[BC21] Talmon Ben-Cnaan. Unprofessional article regarding non-functional sizing: The IEEE Standard 2430 Chair response to IEEE 2430 Non-Functional Sizing Measurements: a numerical placebo. IEEE Software, 38(5):16, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic). See rebuttal [Ano22-63].

Baudry:2021:SRR

[BCE^{+}21] Benoit Baudry, Zimin Chen, Khashayar Etemadi, Han Fu, Davide Ginelli, Steve Kommrusch, Matias Martinez, Martin Monperrus, Javier Ron Arteaga, He Ye, and Zhongxing Yu. A software-repair robot based on continual learning. IEEE Software, 38(4):28–35, July/ August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Barton:2020:SSD

[BCF^{+}20] Laura Barton, Gulipek Candan, Thomas Fritz, Thomas Zimmermann, and Gail C. Murphy. The sound of software development: Music listening among software engineers. IEEE Software, 37(2):78–85, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Brown:2023:DPH

[BCHA23] Adam Brown, Alison Chang, Ben Holtz, and Sarah D Angelo. Developer productivity for humans, part 6: Measuring flow, focus, and friction for developers. IEEE Software, 40(6):16–21, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bucchiarone:2021:WFM

[BCL^{+}21] Antonio Bucchiarone, Federico Ciccozzi, Leen Lambers, Alfonso Pierantonio, Matthias Tichy, Massimo Tisi, Andreas Wortmann, and Vadim Zaytsev. What is the future of modeling? IEEE Software, 38(2): 119–127, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Boehme:2021:FCR

[BCR21] Marcel Boehme, Cristian Cadar, and Abhik ROY-CHOUDHURY. Fuzzing: Challenges and reflections. IEEE Software, 38(3):79–86, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Benton:2020:MLS

[Ben20] William C. Benton. Machine learning systems and intelligent applications. IEEE Software, 37(4):43–49, July/ August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Beyer:2020:AST

[Bey20]  Justin Beyer.  Adam

Shostack on threat modeling. IEEE Software, 37(6):110–112, November/ December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bobadilla:2024:SFB

[BGBM24] Sofia Bobadilla, Richard Glassey, Alexandre Bergel, and Martin Monperrus. SOBO: a feedback bot to nudge code quality in programming courses. IEEE Software, 41(2):68–76, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bencomo:2022:SBA

[BGH^{+}22] Nelly Bencomo, Jin L. C.

Guo, Rachel Harrison, Hans-Martin Heyn, and Tim Menzies. The secret to better AI and better software (is requirements engineering). IEEE Software, 39 (1):105–110, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Baresi:2021:TCS

[BGL^{+}21] Luciano Baresi, William G.

Griswold, Grace A. Lewis, Marco Autili, Ivano Malavolta, and Christine Julien. Trends and challenges for software engineering in the mobile domain. IEEE Software, 38(1):88–96, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Beetz:2022:GED

[BH22] Florian Beetz and Simon Harrer. GitOps: The evolution of DevOps? IEEE Software, 39(4):70–75, July/ August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Brataas:2020:ASE

[BHHvH20] Gunnar Brataas, Geir Kjetil Hanssen, Nikolas Herbst, and Andre van Hoorn. Agile scalability engineering:

The ScrumScale method. IEEE Software, 37(5):77–84, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bull:2024:GAI

[BK24] Christopher Bull and Ahmed Kharrufa. Generative artificial intelligence assistants in software development education: a vision for integrating generative artificial intelligence into educational practice, not instinctively defending against it. IEEE Software, 41(2): 52–59, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bader:2021:ASE

[BKL^{+}21] Johannes Bader, Sonia Seohyun Kim, Frank Sifei Luan, Satish Chandra, and Erik Meijer. AI in software engineering at Facebook. IEEE Software, 38(4):52–61, July/ August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Black:2020:BCT

[Bla20] Nate Black. Boris Cherny on TypeScript. IEEE Software, 37(2):98–100, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Belles-Munoz:2022:NPP

[BMBDMT22] Marta Bellés-Muñoz, Jordi Baylina, Vanesa Daza, and José L. Muñoz-Tapia. New privacy practices for blockchain software. IEEE Software, 39(3):43–49, May/ June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bosch:2022:AER

[BOBC22] Jan Bosch, Helena Holmström Olsson, Björn Brinne, and Ivica Crnkovic. AI engineering: Realizing the potential of AI. IEEE Software, 39(6):23–27, November/December 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Beller:2021:MGR

[BOBZ21] Moritz Beller, Vince Orgovan, Spencer Buja, and Thomas Zimmermann. Mind the gap: On the relationship between automatically measured and self-reported productivity. IEEE Software, 38(5):24–31, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Borg:2023:PIR

[Bor23a]  Markus Borg.  Pipeline

infrastructure required to meet the requirements on AI. IEEE Software, 40(1): 18–22, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Borg:2023:RTD

[Bor23b] Markus Borg. Requirements on technical debt: Dare to specify them! IEEE Software, 40(2):8–12, March/ April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Borg:2024:REL

[Bor24] Markus Borg. Requirements engineering and large language models: Insights from a panel. IEEE Software, 41 (2):6–10, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bosch:2024:FAH

[Bos24] Jan Bosch. The future is already here. IEEE Software, 41(3):26–32, May/ June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Blumen:2022:SER

[BP22] Robert Blumen and Bill Pollak. Software engineering radio reaches milestone 500th episode. IEEE Software, 39(4):94–96, July/ August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Baltes:2020:NHP

[BPS20] Sebastian Baltes, George Park, and Alexander Serebrenik. Is 40 the new 60? How popular media portrays the employability of older software developers. IEEE Software, 37(6):26–31, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Beraza:2023:IPM

[BVG^{+}23] Iñigo Luja Beraza, Juan Manuel Vara, David Granada, Cristian Gómez, and Francisco Javier Pérez-Blanco. On the impact of project management certification for software projects and practitioners. IEEE Software, 40(5):100–107, September/ October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Bano:2021:RPI

[BZA21] Muneera Bano, Didar Zowghi, and Chetan Arora. Requirements, politics, or individualism: What drives the success of COVID-19 contacttracing apps? IEEE Software, 38(1):7–12, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Borowa:2021:LTD

[BZS21] Klara Borowa, Andrzej Zalewski, and Adam Saczko. Living with technical debt: a perspective from the video game industry. IEEE Software, 38(6):65–70, November/December 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cheikh-Ammar:2020:NRS

[CABD20] Mustapha Cheikh-Ammar, Simon Bourdeau, and Jessica Darveau. Navigating the rough seas of software project resistance: Four project manager personas to help bring the ship safely to harbor. IEEE Software, 37(6):38–43, November/December 2020. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2021:ERM

[CAP21] Jeffrey C. Carver, Silvia Abrahão, and Birgit Penzenstadler. Extracting requirements and modeling information and controlling risk. IEEE Software, 38(3): 121–124, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cabot:2023:LCS

[CC23] Jordi Cabot and Robert Clarisó. Low code for smart software development. IEEE Software, 40(1):89– 93, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Comuzzi:2022:TQA

[CCDM22] Marco Comuzzi, Cinzia Cappiello, Florian Daniel, and Giovanni Meroni. Toward quality-aware transaction validation in blockchains. IEEE Software, 39(4):54–62, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2021:APR

[CCPLS21] Jeffrey Carver, Ricardo Colomo-Palacios, Xabier Larrucea, and Miroslaw Staron. Automatic program repair. IEEE Software, 38(4):122–124, July/ August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2021:AAO

[CCSS21] Jeffrey C. Carver, Rafael Capilla, Alexander Serebrenik, and Miroslaw Staron. Automated authoring, onboarding developers, and extracting decision rationale [practitioners digest]. IEEE Software, 38(5):110–112, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cito:2023:EPE

[CCTH23] Jürgen Cito, Satish Chandra, Chakkrit Tantithamthavorn, and Hadi Hemmati. Expert perspectives on explainability. IEEE Software, 40(3):84–88, May/ June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Combelles:2020:DT

[CEL20] Annie Combelles, Christof Ebert, and Percival Lucena. Design thinking. IEEE Software, 37(2):21–24, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Camara:2023:LRE

[CG23] Javier Cámara and David Garlan. Learning by redoing: an experimental study on the impact of repetition of formative assessments in a formal methods course for software engineers. IEEE Software, 40(6):95–101, November/ December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cleland-Huang:2021:VCA

[CHAVMD21] Jane Cleland-Huang, Ankit Agrawal, Michael Vierhauser, and Christoph Mayr-Dorn. Visualizing change in agile safety-critical systems. IEEE Software, 38 (3):43–51, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carleton:2020:EPA

[CHL^{+}20] Anita D. Carleton, Erin Harper, Michael R. Lyu, Sigrid Eldh, Tao Xie, and Tim Menzies. Expert perspectives on AI. IEEE Software, 37(4):87–94, July/ August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carleton:2020:AEW

[CHM^{+}20] Anita D. Carleton, Erin Harper, Tim Menzies, Tao Xie, Sigrid Eldh, and Michael R. Lyu. The AI effect: Working at the intersection of AI and SE. IEEE Software, 37(4):26–35, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Capiluppi:2023:BRA

[CJ23] Andrea Capiluppi and Letizia Jaccheri. Booting and rebooting academia industry collaborations within software engineering courses. IEEE Software, 40(6):63–71, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cohen:2021:FPR

[CKB^{+}21] Jeremy Cohen, Daniel S. Katz, Michelle Barker, Neil Chue Hong, Robert Haines, and Caroline Jay. The four pillars of research software engineering. IEEE Software, 38(1):97–105, January/February 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Combemale:2021:HGM

[CKM^{+}21] Benoit Combemale, Jorg Kienzle, Gunter Mussbacher, Hyacinth Ali, Daniel Amyot, Mojtaba Bagherzadeh, Edouard Batot, Nelly Bencomo, Benjamin Benni, Jean-Michel Bruel, Jordi Cabot, Betty H. C. Cheng, Philippe Collet, Gregor Engels, Robert Heinrich, Jean-Marc Jezequel, Anne Koziolek, Sebastien Mosser, Ralf Reussner, Houari Sahraoui, Rijul Saini, June Sallou, Serge Stinckwich, Eugene Syriani, and Manuel Wimmer. A hitchhiker’s guide to model-driven engineering for data-centric systems. IEEE Software, 38(4): 71–84, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cheong:2024:ECT

[CKT24] Marc Cheong, Raula Gaikovina Kula, and Christoph Treude. Ethical considerations toward protestware. IEEE Software, 41(3):67–75, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Clark:2022:EDT

[CKWB22] Tony Clark, Vinay Kulkarni, Jon Whittle, and Ruth Breu. Engineering digital twin-enabled systems. IEEE Software, 39(2):16– 19, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Clarke:2023:RWG

[Cla23] Paul Clarke. The remote working genie is out of the office bottle. IEEE Software, 40(4):88–95, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ciolkowski:2021:YTD

[CLM21] Marcus Ciolkowski, Valentina Lenarduzzi, and Antonio Martini. 10 years of technical debt research and practice: Past, present, and future. IEEE Software, 38(6): 24–29, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2022:TDP

[CLSS22] Jeffrey C. Carver, Xabier Larrucea, Alexander Serebrenik, and Miroslaw Staron. Technical debt problems and concerns. IEEE Software, 39(3):116–119, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2021:BSD

[CMP^{+}21] Jeffrey C. Carver, Henry Muccini, Birgit Penzenstadler, Rafael Prikladnicki, Alexander Serebrenik, and Thomas Zimmermann. Behavioral science and diversity in software engineering. IEEE Software, 38(2): 107–112, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Conboy:2023:FHS

[CMSG23] Kieran Conboy, Nils Brede Moe, Viktoria Stray, and Jan Henrik Gundelsby. The future of hybrid software development: Challenging current assumptions. IEEE Software, 40(2):26–33, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Conrad:2020:CMP

[Con20] Adam Conrad. Chris Mc-Cord on Phoenix’s LiveView functionality. IEEE Software, 37(3):98–100, May/ June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Conrad:2021:DYP

[Con21] Adam Conrad. Database of the year: Postgres. IEEE Software, 38(5):130–132, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2020:RER

[CP20] Jeffrey C. Carver and Birgit Penzenstadler. Requirements engineering research: News from the trenches. IEEE Software, 37(2):86–88, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Cardoso-Pereira:2023:SCD

[CPGR^{+}23] Isadora Cardoso-Pereira, Geraldo Gomes, Danilo Monteiro Ribeiro, Alberto de Souza, Danilo Lucena, and Gustavo Pinto. Supporting the careers of developers with disabilities: Lessons from zup innovation. IEEE Software, 40(5):58–65, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2020:CHJ

[CPM^{+}20] Jeffrey C. Carver, Birgit Penzenstadler, Leandro L. Minku, Ricardo Colomo-Palacios, and Xabier Larrucea. Conference highlights: JIT fault prevention, motivated modeling, security in requirements, and improving team performance. IEEE Software, 37(4):83–86, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2021:RIS

[CPSS21] Jeffrey C. Carver, Birgit Penzenstadler, Joel Scheuner, and Miroslaw Staron. (Research) insights for serverless application engineering. IEEE Software, 38(1):123–125, January/February 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Catolino:2020:GDC

[CPT^{+}20] Gemma Catolino, Fabio Palomba, Damian A. Tamburri, Alexander Serebrenik, and Filomena Ferrucci. Gender diversity and community smells: Insights from the trenches. IEEE Software, 37(1):10–16, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2020:BSC

[CS20] Jeffrey C. Carver and Miroslaw Staron. Blockchain and smart contract engineering. IEEE Software, 37(5): 94–96, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2022:CUO

[CS22] Jeffrey C. Carver and Miroslaw Staron. Collaboration using open data. IEEE Software, 39(1):102– 104, February 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Carver:2022:DT

[CSC^{+}22] Jeffrey Carver, Miroslaw Staron, Rafael Capilla, Henry Muccini, and Lorin Hochstein. Digital twins. IEEE Software, 39(2):97– 99, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Camara:2024:FWMb

[CWGS24] Javier Cámara, Rebekka Wohlrab, David Garlan, and Bradley Schmerl. Focusing on what matters: Explaining quality tradeoffs in software-intensive systems via dimensionality reduction. IEEE Software, 41(1): 64–73, January/February 2024. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Chen:2022:OSP

[CZ22] Donghua Chen and Runtong Zhang. An open source project for tuning and analyzing MapReduce performance in Hadoop and Spark. IEEE Software, 39 (1):61–69, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Daun:2023:SEE

[Dau23] Marian Daun. Software engineering education for technical engineering degrees: a comparison with the needs of robotics software engineering education. IEEE Software, 40(6):40–46, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dalpiaz:2020:GRE

[DC20] Fabiano Dalpiaz and Kendra M. L. Cooper. Games for requirements engineers: Analysis and directions. IEEE Software, 37(1):50–59, January/February 2020. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Palma:2023:GSR

[DCD^{+}23] Stefano Dalla Palma, Gemma Catolino, Dario Di Nucci, Damian Andrew Tamburri, and Willem-Jan van den Heuvel. Go serverless with RADON! A practical DevOps experience report. IEEE Software, 40(2): 80–89, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

deGea:2021:RET

[dGEH^{+}21] Juan M. Carrillo de Gea, Christof Ebert, Mohamed Hosni, Aurora Vizcaíno, Joaquín Nicolás, and José L. Fernández-Alemán. Requirements engineering tools: an evaluation. IEEE Software, 38(3):17–24, May/ June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dey:2022:KGS

[DJF22] Tapajit Dey, Willem Jiang, and Brian Fitzgerald. Knights and gold stars: a tale of InnerSource incentivization. IEEE Software, 39(6):88–98, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Daubaris:2024:EOSb

[DLIR^{+}24] Paulius Daubaris, Simo Linkola, Juan F. Inglés-Romero, Javier Berrocal, Cristina Vicente-Chicote, Juan Manuel Murillo, Tommi Mikkonen, and Niko Mäkitalo. Explainability with observation sharing in long collaboration chains of automated systems of systems. IEEE Software, 41(1):74– 86, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Damian:2021:CSM

[DLJ^{+}21] Daniela Damian, Johan Linåker, David Johnson, Tony Clear, and Kelly Blincoe. Challenges and strategies for managing requirements selection in software ecosystems. IEEE Software, 38(6):76–87, November/December 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Debroy:2020:OCC

[DM20] Vidroha Debroy and Senecca Miller. Overcoming challenges with continuous integration and deployment pipelines: an experience report from a small company. IEEE Software, 37 (3):21–29, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Duboc:2020:CRE

[DMBA20] Leticia Duboc, Curtis Mc-Cord, Christoph Becker, and Syed Ishtiaque Ahmed. Critical requirements engineering in practice. IEEE Software, 37(1):17–24, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

deMello:2023:RDI

[dMOU^{+}23] Rafael de Mello, Roberto Oliveira, Anderson Uchôa, Willian Oizumi, Alessandro Garcia, Baldoino Fonseca, and Fernanda de Mello. Recommendations for developers identifying code smells. IEEE Software, 40(2):90–98, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dalpiaz:2020:RED

[DN20] Fabiano Dalpiaz and Nan Niu. Requirements engineering in the days of artificial intelligence. IEEE Software, 37(4):7–10, July/ August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

deNobrega:2023:ICS

[dNRR23] Kristel M. de Nobrega, Anne-Françoise Rutkowski, and Piet Ribbers. The impact of cyberattacks on small states. IEEE Software, 40(4):101–105, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Doolittle:2020:JMW

[Doo20]  Jeff Doolittle.  Jeremy

Miller on waterfall versus agile. IEEE Software, 37(4): 107–C3, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Doolittle:2022:LHI

[Doo22] Jeff Doolittle. Luke Hoban on infrastructure as code. IEEE Software, 39(2):112– 114, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Doolittle:2023:AG

[Doo23a] Jeff Doolittle. APIs with GraphQL. IEEE Software, 40(2):118–120, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Doolittle:2023:LGP

[Doo23b] Jeff Doolittle. Luca Galante on platform engineering. IEEE Software, 40(6):144– 146, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dalpiaz:2024:WRA

[DS24] Fabiano Dalpiaz and Jan-Philipp Steghöfer. Where requirements and agility meet: No-man’s-land or a land of opportunity? IEEE Software, 41(3):7–10, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Doukidis:2020:DTP

[DSE20] Georgios Doukidis, Diomidis Spinellis, and Christof Ebert. Digital transformation — a primer for practitioners. IEEE Software, 37(5):13–21, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Santos:2023:MFA

[dSSCMS23] Ronnie de Souza Santos, Luiz Fernando Capretz, Cleyton Magalhães, and Rodrigo Souza. Myths and facts about a career in software testing: a comparison between students beliefs and professionals experience. IEEE Software, 40(5): 76–84, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Santos:2023:WDT

[dSSSVM23] Ronnie de Souza Santos, Brody Stuart-Verner, and Cleyton Magalhães. What do transgender software professionals say about a career in the software industry? IEEE Software, 40(5):34–40, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

daSilva:2022:TAS

[dSTdAdS22] Marcio Alexandre Pereira da Silva, Valéria Cesário Times, André Magno Costa de Araújo, and Paulo Caetano da Silva. Toward an adaptive software architecture for archetypebased health-care applications. IEEE Software, 39(2): 89–96, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

daSilva:2023:UCD

[dSWLT23] Ana Cristina Franco da Silva, Stefan Wagner, Eddie Lazebnik, and Eyal Traitel. Using a cyber digital twin for continuous automotive security requirements verification. IEEE Software, 40(1):69– 76, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Darbord:2023:MCP

[DVE^{+}23] Gabriel Darbord, Benoît Verhaeghe, Anne Etien, Nicolas Anquetil, Anas Shatnawi, Abderrahmane Seriai, and Mustapha Derras. Migrating the communication protocol of client server applications. IEEE Software, 40(4):11–18, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2024:IIS

[EAM24] Christof Ebert, Silvia Abrahão, and V. S. Mani. From idea to impact: Survival guide for successful products. IEEE Software, 41 (3):20–25, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ernst:2022:ADD

[EB22] Neil A. Ernst and Gabriele Bavota. AI-driven development is here: Should you worry? IEEE Software, 39(2):106–110, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2023:AIC

[EB23] Christof Ebert and Maximilian Beck. Artificial intelligence for cybersecurity. IEEE Software, 40(6):27–34, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Eramo:2022:CDT

[EBC^{+}22] Romina Eramo, Francis Bordeleau, Benoit Combemale, Mark van den Brand, Manuel Wimmer, and Andreas Wortmann. Conceptualizing digital twins. IEEE Software, 39(2):39– 46, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2022:TSS

[EBW22] Christof Ebert, Divith Bajaj, and Michael Weyrich. Testing software systems. IEEE Software, 39(4):8–17, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2024:SP

[ED24] Christof Ebert and Alan M. Davis. Software principles. IEEE Software, 41(2): 21–27, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2023:HBS

[EF23] Christof Ebert and Joachim Fetzer. How to become a software company. IEEE Software, 40(2):19–25, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2023:TTC

[EH23a] Christof Ebert and Ulrich Hemel. Technology trends 2023: The competence challenge. IEEE Software, 40 (3):20–28, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2023:DP

[EH23b] Christof Ebert and Lorin Hochstein. DevOps in practice. IEEE Software, 40(1):

29–36, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2021:ASE

[EKB21] Christof Ebert and Frank Kirschke-Biller. Agile systems engineering. IEEE Software, 38(4):7–15, July/ August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2020:TTW

[EKv20] Christof Ebert, Anh Kim, and Jeroen van Genuchten. Technology trends: Winning with ACES. IEEE Software, 37(3):6–13, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2023:GAS

[EL23] Christof Ebert and Panos Louridas. Generative AI for software practitioners. IEEE Software, 40(4):30–38, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Eldh:2024:YIPb

[Eld24a] Sigrid Eldh. A 40-year impact perspective: Meet your new Editor in Chief. IEEE Software, 41(1):4– 7, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Eldh:2024:GAC

[Eld24b] Sigrid Eldh. Generative AI is changing how and what we learn. IEEE Software, 41 (2):4–5, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Eldh:2024:MYI

[Eld24c] Sigrid Eldh. Making your ideas successful. IEEE Software, 41(3):4–6, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2020:BTP

[ELFCFL20] Christof Ebert, Panos Louridas, Tiago M. Fernández-Caramés, and Paula Fraga-Lamas. Blockchain technologies in practice. IEEE Software, 37(4):17–25, July/ August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2022:ISM

[EM22] Christof Ebert and Jan Maximilian Montenbruck. Industry survey: The magic triangle of the new normal. IEEE Software, 39 (3):12–20, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2021:TDR

[ER21] Christof Ebert and Ruschil Ray. Test-driven requirements engineering. IEEE Software, 38(1):16–24, January/February 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2023:CQA

[ER23] Christof Ebert and Valentina Roldan. Code quality analysis. IEEE Software, 40(5): 20–28, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2020:ST

[ERK20] Christof Ebert, Youssef Rekik, and Rahul Karade. Security test. IEEE Software, 37(2):13–20, March/ April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Elzanaty:2021:ARM

[ERL^{+}21] Farida Elzanaty, Christophe Rezk, Sander Lijbrink, Willem van Bergen, Mark Cote, and Shane McIntosh. Automatic recovery of missing issue type labels. IEEE Software, 38 (3):35–42, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Eismann:2021:SAW

[ESvE^{+}21] Simon Eismann, Joel Scheuner, Erwin van Eyk, Maximilian Schwinger, Johannes Grohmann, Nikolas Herbst, Cristina L. Abad, and Alexandru Iosup. Serverless applications: Why, when, and how? IEEE Software, 38(1):32–39, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2021:TTS

[ET21] Christof Ebert and Bertrand Tavernier. Technology trends: Strategies for the new normal. IEEE Software, 38(2):7–14, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Elordi:2021:BDN

[EUG^{+}21] Unai Elordi, Luis Unzueta, Jon Goenetxea, Sergio Sanchez-Carballido, Ignacio Arganda-Carreras, and Oihana Otaegui. Benchmarking deep neural network inference performance on serverless environments with MLPerf. IEEE Software, 38(1):81–87, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2022:UBV

[EVG22] Christof Ebert, Aurora Vizcaino, and Ruben Grande. Unlock the business value of gamification. IEEE Software, 39(6):15–22, November/December 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2024:MISb

[EVGA24] Christof Ebert, Aurora Vizcaino, Félix O. García, and Julio Suarez Albanchez. Measure and improve software productivity. IEEE Software, 41(1):35–44, January/February 2024. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ebert:2020:IG

[EVM20] Christof Ebert, Aurora Vizcaino, and Antonio Manjavacas. IT governance. IEEE Software, 37(6):13–20, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2020:CYP

[Fai20a] George Fairbanks. Code is your partner in thought. IEEE Software, 37(5):109– 112, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2020:RIT

[Fai20b] George Fairbanks. The rituals of iterations and tests. IEEE Software, 37(6):105– 108, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2020:TNU

[Fai20c] George Fairbanks. Testing numbs us to our loss of intellectual control. IEEE Software, 37(3):93–96, May/ June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2020:UTD

[Fai20d]  George Fairbanks.  Ur-

technical debt. IEEE Software, 37(4):95–98, July/ August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2021:WIG

[Fai21] George Fairbanks. Why is it getting harder to apply software architecture? IEEE Software, 38(4):126– 129, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2022:TKI

[Fai22] George Fairbanks. Two kinds of iteration. IEEE Software, 39(1):114–117, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2023:FTD

[Fai23a]  George Fairbanks.  Fix

tech debt with virtuous cycles. IEEE Software, 40(2): 111–116, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fairbanks:2023:SAS

[Fai23b] George Fairbanks. Software architecture is a set of abstractions. IEEE Software, 40(4):110–113, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Favario:2023:TFO

[Fav23] Leonardo Favario. Toward a free and open sourcedriven public sector: an Italian journey. IEEE Software, 40(4):55–61, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fungprasertkul:2021:MTD

[FB21] Suwichak Fungprasertkul and Rami Bahsoon. Managing technical debt under uncertainty. IEEE Software, 38(6):30–34, November/December 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Furtado:2021:HSO

[FCTP21] Leonardo B. Furtado, Bruno Cartaxo, Christoph Treude, and Gustavo Pinto. How successful are open source contributions from countries with different levels of human development? IEEE Software, 38(2):58–63, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fabijan:2020:OCE

[FDHB20] Aleksander Fabijan, Pavel Dmitriev, Helena Holmstrom Olsson, and Jan Bosch. The online controlled experiment lifecycle. IEEE Software, 37(2): 60–67, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Falazi:2020:UIS

[FLB^{+}20] Ghareeb Falazi, Andrea Lamparelli, Uwe Breitenbuecher, Florian Daniel, and Frank Leymann. Unified integration of smart contracts through service orientation. IEEE Software, 37(5):60–66, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Flores:2021:ASP

[FMP21] Gabriela Castro Flores, Ana M. Moreno, and Lawrence Peters. Agile and software project management antipatterns: Clarifying the partnership. IEEE Software, 38(5):39–47, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Frey:2023:HWL

[Fre23] Rasmus Frey. How we lead successful open-source collaborations in the Danish public sector. IEEE Software, 40(4):19–24, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Freire:2021:PST

[FRP^{+}21] Sávio Freire, Nicolli Rios, Boris Pérez, Camilo Castellanos, Darío Correal, Robert Ramač, Vladimir Mandić, Nebojša Taušan, Alexia Pacheco, Gustavo López, Manoel Mendonça, Clemente Izurieta, Davide Falessi, Carolyn Seaman, and Rodrigo Spínola. Pitfalls and solutions for technical debt management in agile software projects. IEEE Software, 38(6):42–49, November/December 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Furda:2023:JSS

[FSF^{+}23] Andrei Furda, Alessandro Soro, Erwin Fielt, Nicola Bartlett, Colin Fidge, Jesse Steineck, and Lionel van den Berg. Jump-starting software engineering careers. IEEE Software, 40(5):108– 115, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Funk:2023:CLC

[Fun23] Dan Funk. Creating a lowcode business process execution platform with Python, BPMN, and DMN. IEEE Software, 40(1):9–17, January/February 2023. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ferreira:2021:NIP

[FVK21] Thiago Ferreira, Silvia Regina Vergilio, and Marouane Kessentini. Nautilus: an interactive plug-and-play search-based software engineering framework. IEEE Software, 38(5):73–82, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fritzsch:2023:RHP

[FWBW23] Jonas Fritzsch, Marvin Wyrich, Justus Bogner, and Stefan Wagner. Resist the hype!: Practical recommendations to cope with résumé-driven development. IEEE Software, 40(5):41–49, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Griva:2022:SRQ

[GBDC22] Anastasia Griva, Stephen Byrne, Denis Dennehy, and Kieran Conboy. Software requirements quality: Using analytics to challenge assumptions at Intel. IEEE Software, 39(2):80–88, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Garousi:2022:WUT

[GCF22] Vahid Garousi, David Cutting, and Michael Felderer. What users think of COVID-19 contact-tracing apps: an analysis of eight European apps. IEEE Software, 39 (3):22–30, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gorton:2024:OQb

[GFJL24] Ian Gorton, Liz Fong-Jones, and Alf Larsson. Observability Q&A. IEEE Software, 41(1):50–54, January/ February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Grossmann:2020:TAT

[GFVS20] Juergen Grossmann, Michael Felderer, Johannes Viehmann, and Ina Schieferdecker. A taxonomy to assess and tailor risk-based testing in recent testing standards. IEEE Software, 37(1):40– 49, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gren:2020:GDP

[GGJ20] Lucas Gren, Alfredo Goldman, and Christian Jacobsson. Group-development psychology training: The perceived effects on agile software-development teams. IEEE Software, 37 (3):63–69, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Garousi:2020:CGB

[GGT^{+}20] Vahid Garousi, Gorkem Giray, Eray Tuzun, Cagatay Catal, and Michael Felderer. Closing the gap between software engineering education and industrial needs. IEEE Software, 37(2):68–77, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gilb:2021:SRO

[Gil21] Tom Gilb. Scalable requirements: One size can fit all. IEEE Software, 38(4): 16–21, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Green:2023:DPH

[GJH^{+}23] Collin Green, Ciera Jaspan, Maggie Hodges, Lanting He, Demei Shen, and Nan Zhang. Developer productivity for humans, part 5: Onboarding and ramp-up. IEEE Software, 40(5):13–19, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Green:2024:DPHb

[GJHL24] Collin Green, Ciera Jaspan, Maggie Hodges, and Jessica Lin. Developer productivity for humans, part 7: Software quality. IEEE Software, 41(1):25–30, January/ February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Guizani:2020:GIQ

[GLBS20] Mariam Guizani, Lara Letaw, Margaret Burnett, and Anita Sarma. Gender inclusivity as a quality requirement: Practices and pitfalls. IEEE Software, 37(6):7–11, November/December 2020. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Georgiou:2024:WBPb

[GMA^{+}24] Konstantinos Georgiou, Nikolaos Mittas, Apostolos Ampatzoglou, Alexander N. Chatzigeorgiou, and Lefteris Angelis. What is being patented in software engineering?: Empirical evidence from the U.S. Patent and Trademark Office. IEEE Software, 41(1): 123–131, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gallina:2022:MDC

[GMdOB22] Barbara Gallina, Leonardo Montecchi, André Luiz de Oliveira, and Lucas Bressan. Multiconcern, dependability-centered assurance via a qualitative and quantitative coanalysis. IEEE Software, 39(4): 39–47, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Geist:2024:IIR

[GMPS24] Verena Geist, Michael Moser, Josef Pichler, and Florian Schnitzhofer. Innovating industry with research: eknows and Sysparency. IEEE Software, 41 (3):41–48, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gloor:2021:APT

[GMZ21] Martin Gloor, Matija Mazalin, and Adrian Zimmermann. Automated payment terminal testing: How to achieve continuous integration for systems that are almost impossible to virtualize. IEEE Software, 38(6): 17–23, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Guizzardi:2023:TFS

[GPS23] Giancarlo Guizzardi, Oscar Pastor, and Veda C. Storey. Thinking fast and slow in software engineering. IEEE Software, 40(6):139– 142, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Garousi:2021:GLV

[GR21] Vahid Garousi and Austen Rainer. Gray literature versus academic literature in software engineering: a call for epistemological analysis. IEEE Software, 38(5):65–72, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gregory:2020:SRR

[Gre20] Sarah Gregory. Shu-Ha-Ri for RE? An agile approach to requirements engineering practitioner maturity.

IEEE Software, 37(1):68– 71, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gregory:2021:RNN

[Gre21] Sarah Gregory. Requirements for the new normal: Requirements engineering in a pandemic. IEEE Software, 38(2):15–18, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Gregory:2022:WDF

[Gre22] Sarah Gregory. What does the future hold for requirements engineers? IEEE Software, 39(4):18–21, July/ August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Guntha:2021:RYR

[GRMR21] Ramesh Guntha, Sethuraman N. Rao, Henry Muccini, and Maneesha Vinodini Ramesh. Rapid yet robust continuous delivery of software for disaster management scenarios. IEEE Software, 38(4):104–113, July/ August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Garousi:2020:SEP

[GSH20] Vahid Garousi, David C. Shepherd, and Kadir Herkiloglu. Successful engagement of practitioners and software engineering researchers: Evidence from 26 international industry academia collaborative projects. IEEE Software, 37(6): 65–75, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Granlund:2022:MSN

[GSM22] Tuomas Granlund, Vlad Stirbu, and Tommi Mikkonen. Medical software needs calm compliance. IEEE Software, 39(1):19–28, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Garousi:2022:TMM

[GvV22] Vahid Garousi and Erik van Veenendaal. Test maturity model integration: Trends of worldwide test maturity and certifications. IEEE Software, 39(2):71–79, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Halloran:2021:FSD

[Hal21] Timothy J. Halloran. The fog of software design. IEEE Software, 38(3):132–135, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Happe:2022:FSW

[HB22] Lucia Happe and Barbora Buhnova. Frustrations steering women away from software engineering. IEEE Software, 39(4):63–69, July/ August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hoda:2023:AAH

[HDT^{+}23] Rashina Hoda, Hoa Dam, Chakkrit Tantithamthavorn, Patanamon Thongtanunam, and Margaret-Anne Storey. Augmented agile: Humancentered AI-assisted software management. IEEE Software, 40(4):106–109, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Henry:2020:BHD

[Hen20a] Gavin Henry. Bert Hubert on the Domain Name System. IEEE Software, 37(5):122–124, September/ October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Henry:2020:JRO

[Hen20b] Gavin Henry. Justin Richer on OAuth. IEEE Software, 37(1):98–100, January/February 2020. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Henry:2021:MLM

[Hen21] Gavin Henry. Matt Lacey on mobile app usability. IEEE Software, 38(2):134– 136, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Henry:2022:IUC

[Hen22] Gavin Henry. The importance of a usage-centric approach. IEEE Software, 39(6):118–120, November/ December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Henry:2024:DCGb

[Hen24] Gavin Henry. Dave Cross on GitHub actions. IEEE Software, 41(1):146–148, January/February 2024. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hermans:2021:TTE

[Her21] Felienne Hermans. Two things every Java developer should know. IEEE Software, 38(6):110–112, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hemon:2020:IPK

[HFLR20] Aymeric Hemon, Brian Fitzgerald, Barbara Lyonnet, and Frantz Rowe. Innovative practices for knowledge sharing in large-scale DevOps. IEEE Software, 37 (3):30–37, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Huebner:2021:WPF

[HGS^{+}21] Johannes Huebner, Andrea Girardello, Olivia Sliz, Elgar Fleisch, and Alexander Ilic. What people focus on when reviewing your app — an analysis across app categories. IEEE Software, 38 (3):96–105, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hampson:2021:OCD

[HHJ^{+}21] Gary Hampson, Neil Hargreaves, Helmut Jakubowicz, Gareth Williams, and Les Hatton. Open collaboration, data quality, and COVID-19. IEEE Software, 38(3): 137–141, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hochstein:2023:DYS

[HLNP23] Lorin Hochstein, Filippo Lanubile, Laura Nolan, and Rafael Prikladnicki. Developing your software engineering career: Words of advice from seasoned professionals. IEEE Software, 40(5):29–33, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Howar:2021:CWT

[HM21] Falk Howar and Malte Mues. Can we trust theorem provers for industrial AI? IEEE Software, 38(6):104–108, November/ December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hehn:2020:IDT

[HMU^{+}20] Jennifer Hehn, Daniel Mendez, Falk Uebernickel, Walter Brenner, and Manfred Broy. On integrating design thinking for human-centered requirements engineering. IEEE Software, 37(2):25–31, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hochstein:2023:RE

[Hoc23a] Lorin Hochstein. Resource exhaustion. IEEE Software, 40(5):126–128, September/ October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hochstein:2023:WDW

[Hoc23b]  Lorin Hochstein.  Why

don’t we see even more failures? IEEE Software, 40(4): 114–116, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2020:PP

[Hol20a] Gerard J. Holzmann. Predicting the past. IEEE Software, 37(2):10–12, March/ April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2020:TF

[Hol20b] Gerard J. Holzmann. Test fatigue. IEEE Software, 37(4):11–16, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2021:KIS

[Hol21a] Gerard J. Holzmann. Keeping it simple: Agile analysis. IEEE Software, 38 (3):13–16, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holzmann:2021:RC

[Hol21b] Gerard J. Holzmann. Right code. IEEE Software, 38(1): 13–15, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Holbrook:2023:CAT

[Hol23] Luke D. Holbrook. Clarity about transparency: Software transparency and open source software licenses in the U.S. public sector. IEEE Software, 40(4):71–76, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hotomski:2023:MRU

[Hot23] Sofija Hotomski. My REvelation: Unveiling an unseen career in requirements. IEEE Software, 40(5):9–12, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hugues:2022:CSD

[HP22] Jérôme Hugues and Sam Procter. Contracts in system development: From multiconcern analysis to assurance with the architecture analysis and design language. IEEE Software, 39(4):34–38, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hevia:2021:QC

[HPEP21] Jose Luis Hevia, Guido Peterssen, Christof Ebert, and Mario Piattini. Quantum computing. IEEE Software, 38(5):7–15, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Hatton:2022:KCM

[HvG22] Les Hatton and Michiel van Genuchten. Keyless or clueless? Managing runaway software ambition. IEEE Software, 39(4):87–89, July/ August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Inman:2024:DPH

[IAV24] Sarah Inman, Sarah D Angelo, and Bogdan Vasilescu. Developer productivity for humans, part 8: Creativity in software engineering. IEEE Software, 41(2): 11–16, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Isomura:2022:ADS

[IINN22] Atsushi Isomura, Yasuhiro Iida, Ichibe Naito, and Takao Nakamura. Axispot: a distributed spatiotemporal data management system for digital twins of moving objects. IEEE Software, 39(2): 33–38, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

John:2023:EEE

[JAP23] Blooma John, Zeena Alsamarra’i, and Niki Panteli. Enhancing employee experience in the era of hybrid work: The case of Microsoft Viva. IEEE Software, 40(2): 70–79, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jaspan:2023:DMM

[JG23a] Ciera Jaspan and Collin Green. Defining, measuring, and managing technical debt. IEEE Software, 40 (3):15–19, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jaspan:2023:DPHa

[JG23b] Ciera Jaspan and Collin Green. Developer productivity for humans, part 2: Hybrid productivity. IEEE Software, 40(2):13–18, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jaspan:2023:DPHb

[JG23c] Ciera Jaspan and Collin Green. Developer productivity for humans, part 4: Build latency, predictability, and developer productivity. IEEE Software, 40(4): 25–29, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jaspan:2023:HCA

[JG23d] Ciera Jaspan and Collin Green. A human-centered approach to developer productivity. IEEE Software, 40(1):23–28, January/ February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jaskolka:2022:SDT

[JHK22] Jason Jaskolka, Brahim Hamid, and Sahar Kokaly. Software design trends supporting multiconcern assurance. IEEE Software, 39(4): 22–26, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jaspan:2020:ESS

[JJE^{+}20] Ciera Jaspan, Matt Jorde, Carolyn Egelman, Collin Green, Ben Holtz, Edward Smith, Maggie Hodges, Andrea Knight, Liz Kammer, Jill Dicker, Caitlin Sadowski, James Lin, Lan Cheng, Mark Canning, and Emerson Murphy-Hill. Enabling the study of software development behavior with cross-tool logs. IEEE Software, 37(6):44–51, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jayakumar:2021:BMC

[JM21] Sanjeev Jayakumar and Gaurav Maini. Building medical claims-processing software from 12 time zones away. IEEE Software, 38(5):125–129, September/ October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Johnson:2023:UEW

[JM23] Brittany Johnson and Tim Menzies. Unfairness is everywhere, so what to do? An interview with Jeanna Matthews. IEEE Software, 40(6):135–138, November/ December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Johnson:2024:EWSb

[JM24a] Brittany Johnson and Tim Menzies. Ethics: Why software engineers can’t afford to look away. IEEE Software, 41(1):142–144, January/February 2024. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Johnson:2024:FWR

[JM24b] Brittany Johnson and Tim Menzies. Fighting for what’s right: an interview with Marc Canellas. IEEE Software, 41(2): 104–107, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Johnson:2024:PPW

[JM24c] Brittany Johnson and Tim Menzies. The power of positionality why accessibility? An interview with Kevin Moran and Arun Krishnavajjala. IEEE Software, 41 (3):91–94, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jackson:2023:TCH

[JPvdHM23] Victoria Jackson, Rafael Prikladnicki, André van der Hoek, and Lisa Marshall. Team creativity in a hybrid software development world: Eight approaches. IEEE Software, 40(2):60–69, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jumnongsaksub:2020:RSC

[JS20]  Siwapol  Jumnongsaksub

and Kunwadee Sripanidkulchai. Reducing smart contract runtime errors on Ethereum. IEEE Software, 37(5):55–59, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jung:2023:RSE

[Jun23] Jeremy Jung. Randy Shoup on evolving architecture and organization at eBay. IEEE Software, 40(1):98–100, January/February 2023. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Jackson:2022:CTD

[JvdHPE22] Victoria Jackson, André van der Hoek, Rafael Prikladnicki, and Christof Ebert. Collaboration tools for developers. IEEE Software, 39 (2):7–15, April 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Keeling:2021:HR

[Kee21] Michael Keeling. Headwinds to redesign. IEEE Software, 38(2):128–132, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Keeling:2022:LUS

[Kee22a] Michael Keeling. Love unrequited: The story of architecture, agile, and how architecture decision records brought them together. IEEE Software, 39(4):90–93, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Keeling:2022:PAD

[Kee22b] Michael Keeling. The psychology of architecture decision records. IEEE Software, 39(6):114–117, November/ December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kaltenecker:2020:ISM

[KGSA20] Christian Kaltenecker, Alexander Grebhahn, Norbert Siegmund, and Sven Apel. The interplay of sampling and machine learning for software performance prediction. IEEE Software, 37(4): 58–66, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Klotins:2023:CSE

[KGW23] Eriks Klotins, Tony Gorschek, and Magnus Wilson. Continuous software engineering: Introducing an industry readiness model. IEEE Software, 40(4):77–87, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Keeling:2021:GCY

[KHF21] Michael Keeling, Tim Halloran, and George Fairbanks. Garbage collect your technical debt. IEEE Software, 38(5):113–116, September/ October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kuula:2020:TPT

[KHK20] Seppo Kuula, Harri Haapasalo, and Juha-Matti Kosonen. Three phases of transforming a projectbased IT company into a lean and design-led digital service provider. IEEE Software, 37(2):41–48, March/ April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kim:2020:SED

[Kim20] Miryung Kim. Software engineering for data analytics. IEEE Software, 37(4): 36–42, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kirk:2023:ESR

[KLRT23] Diana Kirk, Andrew Luxton-Reilly, and Ewan Tempero. Evaluating student risk group mitigation strategies. IEEE Software, 40(6): 79–86, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kuutila:2021:WDW

[KMFC21] Miikka Kuutila, Mika V.

Mantyla, Umar Farooq, and Maelick Claes. What do we know about time pressure in software development? IEEE Software, 38(5):32–38, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kadron:2024:FSEb

[KNP^{+}24] Ismet Burak Kadron, Yannic Noller, Rohan Padhye, Tevfik Bultan, Corina S. Păsăreanu, and Koushik Sen. Fuzzing, symbolic execution, and expert guidance for better testing. IEEE Software, 41(1):98–104, January/February 2024. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kazman:2020:SES

[KP20] Rick Kazman and Liliana Pasquale. Software engineering in society. IEEE Software, 37(1):7–9, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Knutas:2022:CCS

[KPM^{+}22] Antti Knutas, Victoria Palacin, Giovanni Maccani, Pablo Aragón, Annika Wolff, and Lukas Mocek. Civic code for social change: Lessons in civic tech grassroots for software engineers. IEEE Software, 39(6):64–72, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kula:2022:CIS

[KTH^{+}22] Raula Gaikovina Kula, Christoph Treude, Hideaki Hata, Sebastian Baltes, Igor Steinmacher, Marco Aurélio Gerosa, and Winifred Kula Amini. Challenges for inclusion in software engineering: The case of the emerging Papua New Guinean society. IEEE Software, 39 (3):67–76, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Kirbas:2021:IAP

[KWM^{+}21] Serkan Kirbas, Etienne Windels, Olayori McBello, Kevin Kells, Matthew Pagano, Rafal Szalanski, Vesna Nowack, Emily Rowan Winter, Steve Counsell, David Bowes, Tracy Hall, Saemundur Haraldsson, and John Woodward. On the introduction of automatic program repair in Bloomberg. IEEE Software, 38(4):43–51, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

LaToza:2020:INL

[LaT20] Thomas D. LaToza. Information needs: Lessons for programming tools. IEEE Software, 37(6):52–57, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

LaToza:2023:CDC

[LaT23] Thomas D. LaToza. Connecting design to code. IEEE Software, 40(1):94– 97, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lundell:2022:ESU

[LBF^{+}22] Björn Lundell, Simon Butler, Thomas Fischer, Jonas Gamalielsson, Christoffer Brax, Jonas Feist, Tomas Gustavsson, Andrew Katz, Bengt Kvarnström, Erik Lönroth, and Anders Mattsson. Effective strategies for using open source software and open standards in organizational contexts: Experiences from the primary and secondary software sectors. IEEE Software, 39 (1):84–92, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lenarduzzi:2021:TTD

[LDM^{+}21] Valentina Lenarduzzi, Jeremy Daly, Antonio Martini, Sebastiano Panichella, and Damian Andrew Tamburri. Toward a technical debt conceptualization for serverless computing. IEEE Software, 38(1):40–47, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lindman:2020:EPA

[LHHK20] Juho Lindman, Jennifer Horkoff, Imed Hammouda, and Eric Knauss. Emerging perspectives of application programming interface strategy: a framework to respond to business concerns. IEEE Software, 37(2): 52–59, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Liu:2024:OESb

[LHLLL24] Yan Liu, Abdelwahab Hamou-Lhadj, Jiye Li, and Qinghua Lu. Observability and explainability for software systems decision making. IEEE Software, 41(1):45– 49, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lee:2023:TRS

[LL23] Gichan Lee and Scott Uk-Jin Lee. Toward reliable software analytics: Systematic integration of explanations from different model-agnostic techniques. IEEE Software, 40(3):34–42, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Levy:2021:REW

[LLA^{+}21] Meira Levy, Lin Liu, Daniel Amyot, Eric Yu, Muneef Alshammari, Malak Baslyman, Elizabeth Bjarnason, Christopher Bull, Carlos Henrique C. Duarte, Eduard C. Groen, Sami Jantunen, Sylwia Kopczyńska, Lysanne Lessard, Ita Richardson, Jens Weber, and Ziqi Wei. Requirements engineering for well-being, aging, and health: an overview for practitioners. IEEE Software, 38(3):7–12, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Liu:2020:DPS

[LLP^{+}20] Yue Liu, Qinghua Lu, Hye-Young Paik, Xiwei Xu, Shiping Chen, and Liming Zhu. Design pattern as a service for blockchainbased self-sovereign identity. IEEE Software, 37(5):30–36, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lanubile:2024:TFM

[LMFQ24] Filippo Lanubile, Silverio Martínez-Fernández, and Luigi Quaranta. Training future machine learning engineers: a projectbased course on MLOps. IEEE Software, 41(2):60–67, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Li:2022:OAR

[LMPT22] Xiaozhou Li, Sergio Moreschini, Fabiano Pecorelli, and Davide Taibi. OSSARA: Abandonment risk assessment for embedded open source components. IEEE Software, 39(4):48–53, July/ August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Larrucea:2020:BSC

[LP20] Xabier Larrucea and Cesare Pautasso. Blockchain and smart contract engineering. IEEE Software, 37(5):23–29, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lenarduzzi:2021:STT

[LP21] Valentina Lenarduzzi and Annibale Panichella. Serverless testing: Tool vendors’ and experts’ points of view. IEEE Software, 38(1):54– 60, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lopez-Pernas:2023:SSF

[LPBG^{+}23] Sonsoles López-Pernas, Enrique Barra, Aldo Gordillo, Álvaro Alonso, and Juan Quemada. Scaling student feedback in software engineering massive open online courses. IEEE Software, 40(5):50–57, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

LapasiniLeal:2020:PTS

[LPE^{+}20a] G. C. Lapasini Leal, R. Prikladnicki, C. Ebert, R. Balancieri, and L. Bento Pompermaier. Practices and tools for software start-ups. IEEE Software, 37(1):72– 77, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Leal:2020:PTS

[LPE^{+}20b] Gislaine Camila Lapasini Leal, Rafael Prikladnicki, Christof Ebert, Renato Balancieri, and Leandro Bento Pompermaier. Practices and tools for software start-ups. IEEE Software, 37(1):72– 77, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

LeGoues:2021:APR

[LPRC21] Claire Le Goues, Michael Pradel, Abhik Roychoudhury, and Satish Chandra. Automatic program repair. IEEE Software, 38(4): 22–27, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lehner:2022:DTP

[LPT^{+}22] Daniel Lehner, Jérôme Pfeiffer, Erik-Felix Tinsel, Matthias Milan Strljic, Sabine Sint, Michael Vierhauser, Andreas Wortmann, and Manuel Wimmer. Digital twin platforms: Requirements, capabilities, and future prospects. IEEE Software, 39(2):53–61, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Linaaker:2023:OSS

[LRBM23] Johan Linåker, Gregorio Robles, Deborah Bryant, and Sachiko Muto. Open source software in the public sector: 25 years and still in its infancy. IEEE Software, 40(4):39–44, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Linaaker:2022:CAO

[LRZB22] Johan Linåker, Per Runeson, Anneke Zuiderwijk, and Amanda Brock. Collaborative aspects of open data in software engineering. IEEE Software, 39 (1):31–35, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lin:2021:SST

[LSN21] Xuanyi Lin, Michelle Simon, and Nan Niu. Scientific software testing goes serverless: Creating and invoking metamorphic functions. IEEE Software, 38(1):61– 67, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lopez:2021:BCE

[LSPN21] Tamara Lopez, Helen Sharp, Marian Petre, and Bashar Nuseibeh. Bumps in the code: Error handling during software development. IEEE Software, 38(3):26–34, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lopez:2020:TMP

[LTB^{+}20] Tamara Lopez, Thein T. Tun, Arosha K. Bandara, Mark Levine, Bashar Nuseibeh, and Helen Sharp. Taking the middle path: Learning about security through online social interaction. IEEE Software, 37(1):25–30, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Long:2021:LDS

[LTHY21] Ju Long, Hung-Ying Tai, Shen-Ta Hsieh, and Michael Juntao Yuan. A lightweight design for serverless function as a service. IEEE Software, 38(1):75–80, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

lu:2024:SSM

[lÜlzD24] Tuna Hacalo lu, Hüseyin Ünlü, Ali Y ld z, and Onur Demirörs. Software size measurement: Bridging research and practice. IEEE Software, 41(3):49–58, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lamb:2022:RBI

[LZ22] Chris Lamb and Stefano Zacchiroli. Reproducible builds: Increasing the integrity of software supply chains. IEEE Software, 39(2):62–70, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Lu:2023:RAD

[LZXW23] Qinghua Lu, Liming Zhu, Xiwei Xu, and Jon Whittle. Responsible-AI-by-Design: a pattern collection for designing responsible artificial intelligence systems. IEEE Software, 40 (3):63–71, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mashkoor:2024:TEA

[MAE24] Atif Mashkoor, Wesley K. G.

Assunção, and Alexander Egyed. Teaching engineering of AI-intensive systems. IEEE Software, 41(2):30–35, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maguire:2023:CSS

[Mag23] Laura Maguire. Cognitive skills in software engineering: Operating complex, adaptive systems at speed and scale. IEEE Software, 40(6):130–134, November/ December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maguire:2024:ADWb

[Mag24a] Laura Maguire. Automation doesn’t work the way we think it does. IEEE Software, 41(1):138–141, January/February 2024. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maguire:2024:JCS

[Mag24b] Laura Maguire. Joint cognitive systems: Ideas with impact for designing safer, more resilient automated systems. IEEE Software, 41 (3):96–99, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mahmoud:2023:ERU

[Mah23] Anas Mahmoud. An experience report on using wireframes in a capstone software engineering class.

IEEE Software, 40(6):109– 116, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mahe:2020:MSF

[MAM^{+}20] Nolwen Mahe, Bram Adams, Josianne Marsan, Mathieu Templier, and Sylvie Bissonnette. Migrating a software factory to design thinking: Paying attention to people and mindsets. IEEE Software, 37(2): 32–40, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Manchale:2021:TGR

[Man21]  Akshay Manchale.  Tug

Grall on Redis. IEEE Software, 38(4):130–132, July/ August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Marsden:2022:SCS

[MAWH22] Nicola Marsden, Michael Ahmadi, Volker Wulf, and Karen Holtzblatt. Surfacing challenges in scrum for women in tech. IEEE Software, 39(6):80–87, November/December 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mellegard:2020:CBB

[MBL^{+}20] Niklas Mellegard, Hakan Burden, Daniel Levin, Kenneth Lind, and Ana Magazinius. Contrasting big bang with continuous integration through defect reports. IEEE Software, 37 (3):14–20, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mendonca:2021:MSB

[MBMR21] Nabor C. Mendonça, Craig Box, Costin Manolache, and Louis Ryan. The monolith strikes back: Why Istio migrated from microservices to a monolithic architecture. IEEE Software, 38(5):17–22, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Martini:2023:TPA

[MBPB23] Antonio Martini, Terese Besker, Torsten Posch, and Jan Bosch. TD pulse: Assessing the systematic management of technical debt. IEEE Software, 40(3):54–62, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Machado:2021:GES

[MCPdS21] Leticia S. Machado, Clara Caldeira, Marcelo Gattermann Perin, and Cleidson R. B. de Souza. Gendered experiences of software engineers during the COVID-19 crisis. IEEE Software, 38(2): 38–44, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Masood:2022:HNZ

[MDB22] Zainab Masood, Daniela Damian, and Kelly Blincoe. How New Zealand software companies are adapting work settings with changing times. IEEE Software, 39 (3):77–84, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mani:2022:MS

[ME22] V. S. Mani and Christof Ebert. Medical software. IEEE Software, 39(1):11–18, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Menzies:2020:FLS

[Men20] Tim Menzies. The five laws of SE for AI. IEEE Software, 37(1):81–85, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Menzies:2021:SSK

[Men21] Tim Menzies. Shockingly simple: “KEYS” for better AI for SE. IEEE Software, 38(2):114–118, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Menzies:2023:HSE

[Men23] Tim Menzies. How to sell ethics (using AI): an interview with Alexander Serebrenik. IEEE Software, 40 (3):95–97, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Malavolta:2022:EMA

[MGD^{+}22] Ivano Malavolta, Taher A. Ghaleb, Istvan David, Jasper van Rooijen, and Marielle Stoelinga. Engineering mobile apps for disaster management: The case of COVID-19 apps in the Google Play Store. IEEE Software, 39(3):31–42, May/ June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Menzies:2023:BDF

[MH23] Tim Menzies and Chris Hazard. The best data are fake data?: an interview with Chris Hazard. IEEE Software, 40(5):121–124, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Marcos:2021:AET

[MHPV21] Esperanza Marcos, Roberto Hens, Teresa Puebla, and Juan Manuel Vara. Applying emotional team coaching to software development. IEEE Software, 38(4): 85–93, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mendonca:2021:DSA

[MJGP21] Nabor C. Mendonca, Pooyan Jamshidi, David Garlan, and Claus Pahl. Developing self-adaptive microservice systems: Challenges and directions. IEEE Software, 38(2):70–79, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Menzies:2023:EME

[MJRA23] Tim Menzies, Brittany Johnson, David L. Roberts, and Lauren Alvarez. The engineering mindset is an ethical mindset (we just don’t teach it that way yet). IEEE Software, 40(2): 103–110, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Marnewick:2021:DOP

[ML21] Carl Marnewick and Josef Langerman. DevOps and organizational performance: The fallacy of chasing maturity. IEEE Software, 38(5):

48–55, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mertz:2021:SIP

[MND^{+}21] Jhonny Mertz, Ingrid Nunes, Luca Della Toffola, Marija Selakovic, and Michael Pradel. Satisfying increasing performance requirements with caching at the application level. IEEE Software, 38 (3):87–95, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Makasi:2022:TCP

[MNDT22] Tendai Makasi, Alireza Nili, Kevin C. Desouza, and Mary Tate. A typology of chatbots in public service delivery. IEEE Software, 39 (3):58–66, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Malladi:2020:BBP

[MS20] Suresh S. Malladi and Hemang C. Subramanian. Bug bounty programs for cybersecurity: Practices, issues, and recommendations. IEEE Software, 37(1):31– 39, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Megargel:2021:DBA

[MS21] Alan Megargel and Venky Shankarararman. Digital Banking Accelerator: a service-oriented architecture starter kit for banks. IEEE Software, 38(3):106– 112, May/June 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Muller:2022:SGB

[MS22] Ingo Müller and Bran Selic’.

Stakeholders: Going beyond just end users. IEEE Software, 39(1):112–113, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Maro:2021:MTI

[MSK^{+}21] Salome Maro, Jan-Philipp Steghofer, Eric Knauss, Jennifer Horkoff, Rashida Kasauli, Rebekka Wohlrab, Jesper Lysemose Korsgaard, Florian Wartenberg, Niels Jørgen Strøm, and Ruben Alexandersson. Managing traceability information models: Not such a simple task after all? IEEE Software, 38(5):101–109, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Moe:2023:AWW

[MSmM23] Nils Brede Moe, Viktoria Stray, Darja mite, and Marius Mikalsen. Attractive workplaces: What are engineers looking for? IEEE Software, 40(5):85–93, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Michael:2024:ECSb

[MSW24] Judith Michael, Maike Schwammberger, and Andreas Wortmann. Explaining cyberphysical system behavior with digital twins. IEEE Software, 41(1):55– 63, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Mandic:2021:TNP

[MTR^{+}21] Vladimir Mandić, Nebojša Taušan, Robert Ramač, Sávio Freire, Nicolli Rios, Boris Pérez, Camilo Castellanos, Darío Correal, Alexia Pacheco, Gustavo Lopez, Clemente Izurieta, Davide Falessi, Carolyn Seaman, and Rodrigo Spínola. Technical and nontechnical prioritization schema for technical debt: Voice of TDexperienced practitioners. IEEE Software, 38(6):50–58, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Novielli:2020:LJA

[NCL20] Nicole Novielli, Fabio Calefato, and Filippo Lanubile. Love, joy, anger, sadness, fear, and surprise: SE needs special kinds of AI: a case study on text mining and SE. IEEE Software, 37 (3):86–91, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nili:2022:WCP

[NDY22] Alireza Nili, Kevin C. Desouza, and Tan Yigitcanlar. What can the public sector teach us about deploying artificial intelligence technologies? IEEE Software, 39(6): 58–63, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nejati:2021:NGS

[Nej21] Shiva Nejati. Next-generation software verification: an AI perspective. IEEE Software, 38(3):126–130, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nascimento:2022:JAD

[NFH22] Romulo Nascimento, Eduardo Figueiredo, and Andre Hora. JavaScript API deprecation landscape: a survey and mining study. IEEE Software, 39(3):96– 105, May/June 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nagaria:2020:RSD

[NH20] Bhaveet Nagaria and Tracy Hall. Reducing software developer human errors by improving situation awareness. IEEE Software, 37(6):32–37, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nicolini:2024:UNJb

[NHF24] Thiago Nicolini, Andre Hora, and Eduardo Figueiredo. On the usage of new JavaScript features through transpilers: The Babel case. IEEE Software, 41(1): 105–112, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nelson:2021:GAT

[NI21] Grant Nelson and Clemente Izurieta. A gap in the analysis of technical debt in procedural languages: an experiential report on Go. IEEE Software, 38(6):71–75, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nadri:2021:INP

[NRPN21] Reza Nadri, Gema Rodriguez-Perez, and Meiyappan Nagappan. Insights into nonmerged pull requests in GitHub: Is there evidence of bias based on perceptible race? IEEE Software, 38(2): 51–57, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nogales:2024:SDC

[NVV^{+}24] Borja Nogales, Ivan Vidal, Francisco Valera, Victor Sanchez-Aguero, and Diego R. Lopez. Softwaredriven connectivity orchestration for multidomain network functions virtualization ecosystems. IEEE Software, 41(2):88–97, March/ April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

oglu:2023:EEA

[olTY23] Selma Nazl og lu, Cigdem Turhan, and Ali Yazici. Education with experience: Assessment of a co-op model in undergraduate engineering programs in computing. IEEE Software, 40(6):87–94, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ochodek:2020:LDI

[OSMB20] Miroslaw Ochodek, Miroslaw Staron, Wilhelm Meding, and Jan Bosch. LegacyPro: a DNA-inspired method for identifying process legacies in software development organizations. IEEE Software, 37(6):76–85, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2020:BSS

[Ozk20a] Ipek Ozkaya. The behavioral science of software engineering and human machine teaming. IEEE Software, 37(6):3–6, November/ December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2020:BBS

[Ozk20b]  Ipek Ozkaya.  Building

blocks of software design. IEEE Software, 37(2):3–5, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2020:DV

[Ozk20c] Ipek Ozkaya. The deployment view. IEEE Software, 37(3):3–5, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2020:US

[Ozk20d]  Ipek Ozkaya.  Unantici-

pated scenarios. IEEE Software, 37(5):3–5, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2020:WRD

[Ozk20e] Ipek Ozkaya. What is really different in engineering AI-enabled systems? IEEE Software, 37(4):3–6, July/ August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2020:WSS

[Ozk20f] Ipek Ozkaya. What should a software engineer know? IEEE Software, 37(1):3– 6, January/February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2021:CWR

[Ozk21a] Ipek Ozkaya. Can we really achieve software quality? IEEE Software, 38(3):3–6, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2021:FSE

[Ozk21b] Ipek Ozkaya. The future of software engineering work. IEEE Software, 38(5):3–6, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2021:MWG

[Ozk21c] Ipek Ozkaya. Mom, where are the girls? IEEE Software, 38(2):3–6, March/ April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2021:PTV

[Ozk21d] Ipek Ozkaya. A plea to tool vendors: Do not mislead how technical debt is managed. IEEE Software, 38(6):3–6, November/ December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2021:PHL

[Ozk21e] Ipek Ozkaya. Protecting the health and longevity of the peer-review process in the software engineering community. IEEE Software, 38(1):3–6, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2021:WMS

[Ozk21f] Ipek Ozkaya. A watershed moment for searchbased software engineering. IEEE Software, 38(4):3–6, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2022:AEV

[Ozk22a] Ipek Ozkaya. An AI engineer versus a software engineer. IEEE Software, 39(6):4–7, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2022:ACD

[Ozk22b]  Ipek Ozkaya.  Architec-

tural concerns of digital twins. IEEE Software, 39 (2):3–6, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2022:BB

[Ozk22c] Ipek Ozkaya. Barry Boehm:

1935–2022. IEEE Software, 39(6):13–14, November/December 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2022:CGD

[Ozk22d] Ipek Ozkaya. Crossing the great divide of software engineering. IEEE Software, 39 (4):4–7, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2022:DN

[Ozk22e] Ipek Ozkaya. The developer nation. IEEE Software, 39(1):3–6, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2022:UBT

[Ozk22f] Ipek Ozkaya. Understanding and building trust in software systems. IEEE Software, 39(3):3–6, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2023:ALL

[Ozk23a] Ipek Ozkaya. Application of large language models to software engineering tasks: Opportunities, risks, and implications. IEEE Software, 40(3):4–8, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2023:CAK

[Ozk23b] Ipek Ozkaya. Can architecture knowledge guide software development with generative AI? IEEE Software, 40(5):4–8, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2023:ICS

[Ozk23c] Ipek Ozkaya. Infrastructure as code and software architecture conformance checking. IEEE Software, 40(1):4– 8, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2023:LBL

[Ozk23d] Ipek Ozkaya. Looking back, looking forward: Reflections of an Editor in Chief. IEEE Software, 40(6):4–10, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2023:NFS

[Ozk23e] Ipek Ozkaya. The next frontier in software development: AI-augmented software development processes. IEEE Software, 40(4):4–9, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ozkaya:2023:NSB

[Ozk23f] Ipek Ozkaya. No silver bullets, only disciplined reflections to guide course corrections [from the Editor]. IEEE Software, 40 (2):4–7, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pahl:2021:CDD

[PA21] Claus Pahl and Shelernaz Azimi. Constructing dependable data-driven software with machine learning. IEEE Software, 38(6):88–97, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Panis:2022:RCN

[Pan22] Michael Panis. Reflections on choosing a new requirements management tool. IEEE Software, 39(1):7–10, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Penzenstadler:2022:SDT

[PAS^{+}22] Birgit Penzenstadler, Silvia Abrahão, Miroslaw Staron, Jeffrey C. Carver, and Lorin Hochstein. Software design trends supporting multiconcern assurance. IEEE Software, 39(4):84–86, July/ August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Paulisch:2023:LLL

[PBB23] Frances Paulisch, Matthias Backert, and Thomas Blum. Lessons learned from a learning program for software architects. IEEE Software, 40(6):55–62, November/December 2023. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Petre:2020:BSS

[PBC^{+}20] Marian Petre, Jim Buckley, Luke Church, Margaret-Anne Storey, and Thomas Zimmermann. Behavioral science of software engineering. IEEE Software, 37(6): 21–25, November/December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Penzenstadler:2020:WDD

[Pen20] Birgit Penzenstadler. When does design help thinking, and when does design thinking help? IEEE Software, 37 (2):6–9, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Prikladnicki:2020:TAB

[PLC20] Rafael Prikladnicki, Casper Lassenius, and Jeffrey C. Carver. Trends in agile: Business agility. IEEE Software, 37(1):78–80, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pfeiffer:2023:LLT

[PLT23] Rolf-Helge Pfeiffer, Mircea Lungu, and Paolo Tell. Live is life: Teaching software engineering on live systems. IEEE Software, 40(6):117– 125, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Prado:2021:HTI

[PMGP21] Rafa Prado, Wendy Mendes, Kiev S. Gama, and Gustavo Pinto. How transinclusive are hackathons? IEEE Software, 38(2):26–31, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Poort:2020:BWW

[Poo20] Eltjo Poort. Between the waterfall wasteland and the agile outback. IEEE Software, 37(1):92–97, January/ February 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pretorius:2021:CIR

[PREL21] Carianne Pretorius, Maryam Razavian, Katrin Eling, and Fred Langerak. Combined intuition and rationality increases software feature novelty for female software designers. IEEE Software, 38(2):64–69, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

PazurAnicic:2022:TMS

[PS22] Katarina Pažur Aničić and Zlatko Stapić. Teaching methods in software engineering: a systematic review. IEEE Software, 39(6): 73–79, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pirozhnikov:2023:MST

[PS23] Alexey Pirozhnikov and Han Schaminée. Maritime’s smart tech revolution. IEEE Software, 40 (3):89–94, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Penzenstadler:2023:ATE

[PSAH23] Birgit Penzenstadler, Miroslaw Staron, Silvia Abrahão, and Lorin Hochstein. AI, tech, energy, and collaboration. IEEE Software, 40(3):80–83, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Perez-Soler:2021:CCD

[PSJPGdL21] Sara Perez-Soler, Sandra Juarez-Puerta, Esther Guerra, and Juan de Lara. Choosing a chatbot development tool. IEEE Software, 38(4): 94–103, July/August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Penzenstadler:2023:WUC

[PSSH23] Birgit Penzenstadler, Alexander Serebrenik, Miroslaw Staron, and Lorin Hochstein. From the war in Ukraine to cannabis use: Exploring a diverse set of papers. IEEE Software, 40(2): 99–102, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pasquale:2022:DAC

[PZC^{+}22] Liliana Pasquale, Paola Zippo, Cliona Curley, Brian O Neill, and Marina Mongiello. Digital age of consent and age verification: Can they protect children? IEEE Software, 39(3):50–57, May/ June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Pautasso:2020:UBH

[PZZ^{+}20] Cesare Pautasso, Olaf Zimmermann, Liming Zhu, Thomas Bocek, and Xabier Larrucea. Unchain or block the hype? : Decision drivers, success factors, and perspectives for blockchain adoption. IEEE Software, 37(5):97–103, September/ October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Quattrocchi:2023:IC

[QT23] Giovanni Quattrocchi and Damian A. Tamburri. Infrastructure as code. IEEE Software, 40(1):37–40, January/February 2023. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rudmark:2022:FLO

[RA22] Daniel Rudmark and Magnus Andersson. Feedback loops in open data ecosystems. IEEE Software, 39 (1):43–47, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Raghavan:2021:JSS

[Rag21] Priyanka Raghavan. James Smith on software bugs and quality. IEEE Software, 38(3):142–144, May/June 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Raghavan:2023:RJA

[Rag23] Priyanka Raghavan. Ross John Anderson on software obsolescence. IEEE Software, 40(5):130–132, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rainer:2020:AJV

[Rai20] Austen Rainer. Anaya’s journey: a vision for a future software academy. IEEE Software, 37(2):95–96, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rios:2021:RBT

[RFP^{+}21] Nicolli Rios, Savio Freire, Boris Perez, Camilo Castellanos, Dario Correal, Manoel Mendonca, Davide Falessi, Clemente Izurieta, Carolyn B. Seaman, and Rodrigo Oliveira Spinola. On the relationship between technical debt management and process models. IEEE Software, 38(5):56–64, September/October 2021.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rivera:2022:TAS

[RJV^{+}22] Luis F. Rivera, Miguel Jiménez, Norha M. Villegas, Gabriel Tamura, and Hausi A. Müller. Toward autonomic, software-intensive digital twin systems. IEEE Software, 39(2):20–26, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Robillard:2020:LLW

[RK20] Martin P. Robillard and Kaylee Kutschera. Lessons learned while migrating from Swing to JavaFX. IEEE Software, 37(3):78–85, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rudmark:2023:BPH

[RLTD23] Daniel Rudmark, Juho Lindman, Andreas Tryti, and Brede Dammen. Beyond procurement: How Entur navigated the open source journey to advance public transport. IEEE Software, 40(4):62–70, July/ August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Rodriguez-Mancini:2022:TFD

[RMLV^{+}22] Diego Rodríguez-Mancini, Zheng Li, Camilo Valenzuela, Guillermo Cabrera-Vives, and Francisco Förster. Toward fractal development of data processing-intensive artificial intelligence systems. IEEE Software, 39(6): 28–34, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Ristov:2021:DDA

[RPWF21] Sasko Ristov, Stefan Pedratscher, Jakob Wallnoefer, and Thomas Fahringer. DAF: Dependency-aware FaaSifier for Node.js monolithic applications. IEEE Software, 38(1):48–53, January/February 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

RozantedePaula:2020:DTI

[RSN20] Talita Rozante de Paula, Tatiane Santana Amancio, and Jose Adan Nonato Flores. Design thinking in industry. IEEE Software, 37(2): 49–51, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

rincThurnay:2022:SOL

[rTRN^{+}22] L rinc Thurnay, Bernhard Riedl, Anna-Sophie Novak, Verena Schmid, and Thomas J. Lampoltshammer. Solving an open legal data puzzle with an interdisciplinary team. IEEE Software, 39(1):55–60, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Randall:2024:WAE

[RWS^{+}24] Natasha Randall, Dennis Wäckerle, Nils Stein, Dennis Goßler, and Stefan Bente. What an AI-embracing software engineering curriculum should look like: an empirical study. IEEE Software, 41(2):36–43, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Staron:2024:TDL

[SAGS24] Miroslaw Staron, Silvia Abrahão, Gregory Gay, and Alexander Serebrenik. Testing, debugging, and log analysis with modern AI tools. IEEE Software, 41(2): 99–102, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Staron:2023:PSS

[SAHH23] Miroslaw Staron, Silvia Abrahão, Chetan Honnenahalli, and Jennifer Horkoff. Privacy, security, soft requirements, and flaky tests. IEEE Software, 40(5):117– 120, September/October 2023. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Staron:2023:RRI

[SAPH23] Miroslaw Staron, Silvia Abrahão, Birgit Penzenstadler, and Lorin Hochstein. Recent research into infrastructure as code. IEEE Software, 40(1):86–88, January/ February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Simmonds:2021:IAA

[SBHK21] Jocelyn Simmonds, Maria Cecilia Bastarrica, and Nancy Hitschfeld-Kahler. Impact of affirmative action on female computer science/software engineering undergraduate enrollment. IEEE Software, 38(2):32–37, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Seetharam:2024:EDK

[SBL^{+}24] Balachandran Seetharam, Druthil Belur, Xabier Larrucea, Giuliano Fragola, Ricardo Colomo-Palacios, Rue Bertsch, and Michiel van Genuchten. Embedding domain knowledge in new software teams. IEEE Software, 41(3):83–89, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Scharinger:2022:CRH

[SBVO22] Boris Scharinger, Markus Borg, Andreas Vogelsang, and Thomas Olsson. Can RE help better prepare industrial AI for commercial scale? IEEE Software, 39(6):8–12, November/December 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Smite:2023:FWC

[SCTR23] Darja Smite, Emily Laue Christensen, Paolo Tell, and Daniel Russo. The future workplace: Characterizing the spectrum of hybrid work arrangements for software teams. IEEE Software, 40(2):34–41, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sawant:2021:NHB

[SD21] Anand Ashok Sawant and Premkumar Devanbu. Naturally!: How breakthroughs in natural language processing can dramatically help developers. IEEE Software, 38(5):118–123, September/ October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Selic:2020:CCW

[Sel20] Bran Selic. Controlling the controllers: What software people can learn from control theory. IEEE Software, 37(6):99–103, November/ December 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sampe:2021:TMA

[SGLSA^{+}21] Josep Sampe, Pedro Garcia-Lopez, Marc Sanchez-Artigas, Gil Vernik, Pol Roca-Llaberia, and Aitor Arjona. Toward multicloud access transparency in serverless computing. IEEE Software, 38(1):68–74, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sen:2022:TDQ

[SHG^{+}22] Sagar Sen, Erik Johannes Husom, Arda Goknil, Simeon Tverdal, Phu Nguyen, and Iker Mancisidor. Taming data quality in AIenabled industrial Internet of Things. IEEE Software, 39(6):35–42, November/December 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sen:2024:UAV

[SHG^{+}24] Sagar Sen, Erik Johannes Husom, Arda Goknil, Simeon Tverdal, and Phu Nguyen. Uncertainty-aware virtual sensors for cyber-physical systems. IEEE Software, 41(2):77–87, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shringi:2021:RRM

[Shr21] Kanchan Shringi. Ryan Ripley on making Scrum work. IEEE Software, 38(1): 126–128, January/February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shringi:2023:MFB

[Shr23]  Kanchan Shringi.  Matt

Frisbie on browser extensions. IEEE Software, 40(4): 118–120, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Singh:2020:RRE

[SL20] Irish Singh and Seok-Won Lee. RE BBC: Requirements engineering in a blockchain-based cloud system: Its role in servicelevel agreement specification. IEEE Software, 37(5): 7–12, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Stray:2020:DSM

[SMS20] Viktoria Stray, Nils Brede Moe, and Dag I. K. Sjoberg. Daily stand-up meetings: Start breaking the rules. IEEE Software, 37(3):70–77, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sas:2021:PAS

[SPAF21] Darius Sas, Ilaria Pigazzini, Paris Avgeriou, and Francesca Arcelli Fontana. The perception of architectural smells in industrial practice. IEEE Software, 38(6):35–41, November/December 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spalazzese:2020:IIM

[SPE20] Romina Spalazzese, Patrizio Pelliccione, and Ulrik Eklund. INTERO: an interoperability model for large systems. IEEE Software, 37 (3):38–45, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2023:CAC

[Spi23] Diomidis Spinellis. Commands as AI conversations. IEEE Software, 40(6):22–26, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2024:AMO

[Spi24a]  Diomidis Spinellis.  Au-

tomating a massive open online course’s production. IEEE Software, 41(2):17–20, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2024:PPG

[Spi24b] Diomidis Spinellis. Pair programming with generative AI. IEEE Software, 41 (3):16–18, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Spinellis:2024:UTSb

[Spi24c] Diomidis Spinellis. Unit tests for SQL. IEEE Software, 41(1):31–34, January/ February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

SacchieSouza:2023:ICF

[SPS23] Isac Sacchi e Souza, Daniel Pinheiro Franco, and João Pedro São Gregorio Silva. Infrastructure as code as a foundational technique for increasing the DevOps maturity level: Two case studies. IEEE Software, 40(1): 63–68, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Subramanian:2022:AFC

[SRNK22] Vikram N. Subramanian, Ifraz Rehman, Meiyappan Nagappan, and Raula Gaikovina Kula. Analyzing first contributions on GitHub: What do newcomers do?

IEEE Software, 39(1):93– 101, February 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sedano:2020:DTD

[SRP20] Todd Sedano, Paul Ralph, and Cecile Peraire. Dualtrack development. IEEE Software, 37(6):58–64, November/December 2020. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Stirbu:2022:TMS

[SRR^{+}22] Vlad Stirbu, Mikko Raatikainen, Joel Röntynen, Vlas Sokolov, Timo Lehtonen, and Tommi Mikkonen. Toward multiconcern software development with everything as code. IEEE Software, 39(4): 27–33, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Serebrenik:2022:AER

[SSC^{+}22] Alexander Serebrenik, Miroslaw Staron, Jordi Cabot, Birgit Penzenstadler, Lorin Hochstein, and Jeffrey C. Carver. AI engineering research in software engineering venues. IEEE Software, 39(6):105–108, November/ December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sporsem:2023:UMH

[SSH23] Tor Sporsem, Audun Fauchald Strand, and Geir Kjetil Hanssen. Unscheduled meetings in hybrid work. IEEE Software, 40(2):42–49, March/April 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Siebra:2020:CFE

[SSQ^{+}20] Clauirton Siebra, Leonardo Sodre, Jonysberg Quintino, Fabio Q. B. da Silva, and Andre L. M. Santos. Collaborative feedback and its effects on software teams. IEEE Software, 37(5):85–93, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sagodi:2022:EAE

[SSvG22] André Sagodi, Johannes Schniertshauer, and Benjamin van Giffen. Engineering AI-enabled computer vision systems: Lessons from manufacturing. IEEE Software, 39(6):51–57, November/December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Segura:2020:MTT

[STZC20] Sergio Segura, Dave Towey, Zhi Quan Zhou, and Tsong Yueh Chen. Metamorphic testing: Testing the untestable.

IEEE Software, 37(3):46–53, May/June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Smite:2020:OER

[SvC20] Darja Smite, Rini van Solingen, and Panagiota Chatzipetrou. The offshoring elephant in the room: Turnover. IEEE Software, 37(3):54–62, May/ June 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sokolowski:2023:DIC

[SWS23] Daniel Sokolowski, Pascal Weisenburger, and Guido Salvaneschi. Decentralizing infrastructure as code. IEEE Software, 40(1):50– 55, January/February 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sykacek:2020:BAR

[Syk20]  Peter Sykacek.  Biased

against race or age? IEEE Software, 37(5):22, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shi:2022:IIS

[SYW22] Lin Shi, Ye Yang, and Qing Wang. iSPY: Issue solution pairing from developer live chats. IEEE Software, 39(6):109–113, November/ December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Shaw:2022:CSE

[SZ22] Mary Shaw and Liming Zhu. Can software engineering harness the benefits of advanced AI? IEEE Software, 39(6):99–104, November/ December 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Taggart:2024:JCM

[Tag24a] Sam Taggart. Jason C. McDonald on quantified tasks. IEEE Software, 41(3): 100–102, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Taggart:2024:MSF

[Tag24b] Sam Taggart. M. Scott Ford on managing dependency freshness. IEEE Software, 41(2):109–112, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tantithamthavorn:2023:EAS

[TCHC23] Chakkrit Tantithamthavorn, Jürgen Cito, Hadi Hemmati, and Satish Chandra. Explainable AI for SE: Challenges and future directions. IEEE Software, 40 (3):29–33, May/June 2023.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tuzun:2022:GTD

[TEB^{+}22] Eray Tüzün, Hakan Erdogmus, Maria Teresa Baldassarre, Michael Felderer, Robert Feldt, and Burak Turhan. Ground-truth deficiencies in software engineering: When codifying the past can be counterproductive. IEEE Software, 39 (3):85–95, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tantithamthavorn:2021:AAS

[TJG21] Chakkrit Tantithamthavorn, Jirayus Jiarpakdee, and John Grundy. Actionable analytics: Stop telling me what it is; please tell me what to do. IEEE Software, 38(4):115–120, July/ August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tang:2021:DMP

[TK21] Antony Tang and Rick Kazman. Decision-making principles for better software design decisions. IEEE Software, 38(6):98–102, November/December 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tenbergen:2024:FSE

[TK24] Bastian Tenbergen and Stephan Krusche. The future of software engineering education and training in the age of AI. IEEE Software, 41(2):28–29, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tenbergen:2023:SEE

[TKHB23] Bastian Tenbergen, Stephan Krusche, Robert Hanna, and Muneera Bano. Software engineering education and training: Industry demands, curriculum deficits, and pedagogy direction. IEEE Software, 40(6):36–39, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tkalich:2024:WHPb

[TmAM24] Anastasiia Tkalich, Darja mite, Nina Haugland Andersen, and Nils Brede Moe. What happens to psychological safety when going remote? IEEE Software, 41(1):113–122, January/February 2024. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Trendowicz:2021:DDT

[TSJ21] Adam Trendowicz, Julien Siebert, and Andreas Jedlitschka. Data-driven technical debt management: Software engineering or data science challenge? IEEE Software, 38(6):59–64, November/December 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Taibi:2021:SCW

[TSW21] Davide Taibi, Josef Spillner, and Konrad Wawruch. Serverless computing-where are we now, and where are we heading? IEEE Software, 38(1):25–31, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Trujillo:2021:NAS

[TVH21] Leonardo Trujillo, Omar M.

Villanueva, and Daniel Eduardo Hernandez. A novel approach for search-based program repair. IEEE Software, 38(4):36–42, July/ August 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Terragni:2023:FPS

[TWRG23] Valerio Terragni, Catherine Watson, Nicholas Rowe, and Nasser Giacaman. Fostering professionalism in software engineering: an early-exposure approach. IEEE Software, 40(6):47–54, November/December 2023.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Tan:2022:SOS

[TZ22] Xin Tan and Minghui Zhou. Scaling open source software communities: Challenges and practices of decentralization. IEEE Software, 39 (1):70–75, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Unlu:2023:HSP

[ÜiYÖTD23] Hüseyin Ünlü, Ozan Ra it Yürüm, Özden Özcan-Top, and Onur Demirörs. How software practitioners perceive work-related barriers and benefits based on their educational backgrounds: Insights from a survey study. IEEE Software, 40(5):66–75, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Vaidhyanathan:2022:ALA

[VCMR22]  Karthik  Vaidhyanathan,

Anish Chandran, Henry Muccini, and Regi Roy. Agile4MLS leveraging agile practices for developing machine learning-enabled systems: an industrial experience. IEEE Software, 39(6): 43–50, November/December 2022. CODEN IESOEG.

ISSN 0740-7459 (print), 1937-4194 (electronic).

Velasco:2020:NDL

[VCPC20] Carlos Velasco, Ricardo Colomo-Palacios, and Ramon Cano. Neural distributed ledger. IEEE Software, 37(5):43–48, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

vandeBeek:2020:BDC

[vdBLvdHvG20] Gerbrand van de Beek, Ronald Liem, Henk Jan van der Horst, and Michiel van Genuchten. Building and deploying a COVID-19 monitoring solution in March. IEEE Software, 37(5):113–120, September/ October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Viseur:2023:COO

[VJ23] Robert Viseur and Nicolas Jullien. Communes-Plone: an original open source model of resource pooling in the public sector. IEEE Software, 40(4): 46–54, July/August 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Vakkuri:2020:CSI

[VKKA20] Ville Vakkuri, Kai-Kristian Kemell, Joni Kultanen, and Pekka Abrahamsson. The current state of industrial practice in artificial intelligence ethics. IEEE Software, 37(4):50–57, July/ August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Verdecchia:2021:GIG

[VLEd21] Roberto Verdecchia, Patricia Lago, Christof Ebert, and Carol de Vries. Green IT and green software. IEEE Software, 38(6):7–15, November/December 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Verdugo:2024:CRP

[VORP24] Javier Verdugo, Jesús Oviedo, Moisés Rodríguez, and Mario Piattini. Connecting research and practice for software product quality evaluation and certification: a software laboratory’s 25-year journey. IEEE Software, 41(3):33–40, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Venkataswamy:2020:ISE

[VS20] Varalakshmi Venkataswamy and Balachandran Seetharam. The impact of software on eyecare in India. IEEE Software, 37(4):99–106, July/ August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Verhaeghe:2022:GAE

[VSS^{+}22] Benoît Verhaeghe, Anas Shatnawi, Abderrahmane Seriai, Anne Etien, Nicolas Anquetil, Mustapha Derras, and Stéphane Ducasse. From GWT to angular: an experiment report on migrating a legacy Web application. IEEE Software, 39(4):76–83, July/August 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

vanLuijt:2020:BSK

[vV20] Bob van Luijt and Micha Verhagen. Bringing semantic knowledge graph technology to your data. IEEE Software, 37(2):89–94, March/April 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wilson:2024:ERI

[WAHJ24] Greg Wilson, Jorge Aranda, Michael Hoye, and Brittany Johnson. Experience report: It will never work in theory. IEEE Software, 41 (3):80–82, May/June 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wasserman:2023:PSE

[Was23a] Anthony I. Wasserman. Preparing software engineers for management roles. IEEE Software, 40(5):94–99, September/October 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wasserman:2023:ETS

[Was23b] Anthony I. Tony Wasserman. Emulating a tech startup in a university: Everything but the code. IEEE Software, 40(6):72–78, November/December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wagner:2024:NUEb

[WBR24] Matthias Wagner, Markus Borg, and Per Runeson. Navigating the upcoming European Union AI Act. IEEE Software, 41(1):19– 24, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wang:2023:CDH

[WCF^{+}23] Zhendong Wang, Yi-Hung Chou, Kayla Fathi, Tobias Schimmer, Peter Colligan, David Redmiles, and Rafael Prikladnicki. Co-designing for a hybrid workplace experience in software development. IEEE Software, 40(2): 50–59, March/April 2023.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wang:2022:HDO

[WCG22] Wenting Wang, Jinghui Cheng, and Jin L. C. Guo. How do open source software contributors perceive and address usability?: Valued factors, practices, and challenges. IEEE Software, 39(1):76–83, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Weyer:2021:CWA

[WDT21] Thorsten Weyer, Marian Daun, and Bastian Tenbergen. The changing world and the adapting machine: How digital transformation changes requirements engineering in the embedded and cyberphysical systems industry. IEEE Software, 38(5):83–91, September/October 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wendt:2023:REO

[Wen23] Kevin D. Wendt. A replication of and extension to the original code mangler study. IEEE Software, 40(6):102–108, November/ December 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Whittle:2021:CHV

[WFSH21] Jon Whittle, Maria Angela Ferrario, Will Simm, and Waqar Hussain. A case for human values in software engineering. IEEE Software, 38(1):106–113, January/February 2021. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Winston:2022:PBF

[Win22a]  Philip Winston.  Paul

Butcher on fuzz testing. IEEE Software, 39(1):118– 120, February 2022. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Winston:2022:TSD

[Win22b]  Philip Winston.  Time-

series databases and Amazon Timestream. IEEE Software, 39(3):126–128, May/June 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wang:2020:NUF

[WLL20] Dabao Wang, Kui Liu, and Li Li. On the need of understanding the failures of smart contracts. IEEE Software, 37(5):49–54, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Williams:2021:DNT

[WLO21] Kevin Williams, Yan Li, and Lorne Olfman. DRT: a novel tool for data archiving. IEEE Software, 38(2): 88–95, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Worthington:2022:WEN

[WPC22] Simon Worthington, Harshada Patel, and Michal Charemza. What else is new? Open data users need to know what’s changed. IEEE Software, 39(1):48–54, February 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wiese:2020:NPD

[WPP20] Igor Wiese, Ivanilton Polato, and Gustavo Pinto. Naming the pain in developing scientific software. IEEE Software, 37(4):75–82, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wohlrab:2021:WHY

[WPSK21] Rebekka Wohlrab, Patrizio Pelliccione, Ali Shahrokni, and Eric Knauss. Why and how your traceability should evolve: Insights from an automotive supplier. IEEE Software, 38(4): 62–70, July/August 2021.

CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wang:2020:AIP

[WYW20] Junjie Wang, Ye Yang, and Qing Wang. Artificial intelligence-powered worker engagement in software crowdsourcing. IEEE Software, 37(6):94–98, November/December 2020. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wohrer:2020:DSL

[WZ20] Maximilian Wohrer and Uwe Zdun. From domain-specific language to code: Smart contracts and the application of design patterns. IEEE Software, 37(5):37–42, September/October 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Wessel:2023:GDB

[WZGS23] Mairieli Wessel, Andy Zaidman, Marco A. Gerosa, and Igor Steinmacher. Guidelines for developing bots for GitHub. IEEE Software, 40 (3):72–79, May/June 2023. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Xiu:2021:ESM

[XJA21] Minke Xiu, Zhen Ming Jack Jiang, and Bram Adams. An exploratory study of machine learning model stores. IEEE Software, 38(1):114–122, January/ February 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zacchiroli:2021:GDP

[Zac21] Stefano Zacchiroli. Gender differences in public code contributions: a 50-year perspective. IEEE Software, 38(2):45–50, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zowghi:2023:WMR

[ZB23] Didar Zowghi and Muneera Bano. What’s missing in requirements engineering for responsible AI? IEEE Software, 40(6):11–15, November/December 2023. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zhang:2021:UWT

[ZCG^{+}21] Jiayun Zhang, Yang Chen, Qingyuan Gong, Xin Wang, Aaron Yi Ding, Yu Xiao, and Pan Hui. Understanding the working time of developers in IT companies in China and the United States. IEEE Software, 38(2):96–106, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zhou:2022:OAH

[ZHX22] Minghui Zhou, Xinwei Hu, and Wei Xiong. openEuler: Advancing a hardware and software application ecosystem. IEEE Software, 39 (2):101–105, April 2022. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zhang:2020:DLB

[ZLG^{+}20] Tao Zhang, Ying Liu, Jerry Gao, Li Peng Gao, and Jing Cheng. Deep learningbased mobile application isomorphic GUI identification for automated robotic testing. IEEE Software, 37(4): 67–74, July/August 2020. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zhang:2021:QAY

[ZMR21] Lei Zhang, Andriy Miranskyy, and Walid Rjaibi. Quantum advantage and the Y2K bug: a comparison. IEEE Software, 38(2): 80–87, March/April 2021. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zimmermann:2024:CIDb

[ZPKS24] Olaf Zimmermann, Cesare Pautasso, Stefan Kapferer, and Mirko Stocker. Continuous integration and delivery in open source development and pattern publishing: Lessons learned with tool setup and pipeline evolution. IEEE Software, 41 (1):9–18, January/February 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Zdun:2020:ETC

[ZWL20] Uwe Zdun, Erik Wittern, and Philipp Leitner. Emerging trends, challenges, and experiences in DevOps and microservice APIs. IEEE Software, 37(1):87–91, January/February 2020. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).